### v1

# One missing piece in Vision and Language: A Survey on Comics Understanding

Abstract
Vision-language models have recently evolved into versatile systems capable of high performance across a range of tasks, such as document understanding, visual question answering, and grounding, often in zero-shot settings. Comics Understanding, a complex and multifaceted field, stands to greatly benefit from these advances. Comics, as a medium, combine rich visual and textual narratives, challenging AI models with tasks that span image classification, object detection, instance segmentation, and deeper narrative comprehension through sequential panels. However, the unique structure of comics—characterized by creative variations in style, reading order, and non-linear storytelling—presents a set of challenges distinct from those in other visual-language domains. In this survey, we present a comprehensive review of Comics Understanding  from both dataset and task perspectives. Our contributions are fivefold: (1) We analyze the structure of the comics medium, detailing its distinctive compositional elements; (2) We survey the widely used datasets and tasks in comics research, emphasizing their role in advancing the field; (3) We introduce the Layer of Comics Understanding  (LoCU) framework, a novel taxonomy that redefines vision-language tasks within comics and lays the foundation for future work; (4) We provide a detailed review and categorization of existing methods following the LoCU framework; (5) Finally, we highlight current research challenges and propose directions for future exploration, particularly in the context of vision-language models applied to comics. This survey is the first to propose a task-oriented framework for comics intelligence and aims to guide future research by addressing critical gaps in data availability and task definition. A project associated with this survey is available at https://github.com/emanuelevivoli/awesome-comics-understanding.
Index Terms:
Survey, Comics, Manga, Layers of Understanding, Comics Understanding, Vision-Language tasks
1 Introduction
Comics serve as a sophisticated medium that combines visual and textual elements to tell stories [1]. Applying recent machine learning tools to understand comics is complicated. This complexity comes from its unique domain that, driven by the author’s creativity, exhibits unique variations in both style and content. For example, modeling the drawing style is complicated, the reading order changes from one book to another, and, unlike natural images, comics frequently depict transformations that defy physical laws, adding another layer of complexity.
Motivated by these challenges, advanced artificial intelligence (AI) approaches have been increasingly applied to Comics Understanding. Given that AI thrives on tackling complex and diverse tasks, many researchers are now focusing on problems like object detection [2], semantic segmentation [3], optical character recognition (OCR) [4], recurrence of characters and objects in varying contexts [5], pose estimation [6], depth estimation [7], and integration of visual and textual narratives [8]. Comics also include complex narrative elements such as satire and irony and spatial and temporal dynamics within the story that have not been deeply investigated. Many of these tasks, individually or in combination, represent the cutting edge of advanced AI, particularly in the area of Multimodal Vision-Language understanding [9, 10, 11, 12, 13].
Figure 1: Number of publications on Comics and Manga (from Google Scholar). The publications have been filtered by keywords (manga, comics, graphic novels), topics and journals (computer science).
Historically, in machine learning (and deep learning), innovation has been driven by the available data, both in terms of tasks and more capable models. As new types of annotated data and more complex tasks emerged, the need for innovative approaches grew, particularly in domains that combine multiple modalities. The domain of multimodal learning, especially the convergence of vision and language, has witnessed significant advancements. These advancements include the shift away from independent uni-modal representations [14, 15] toward the alignment and integration of different modalities [16, 17, 18, 19, 20, 21] and the creation of multimodal embedding space combining aligned uni-modal inputs [22, 23], with early and late fusion. Trained on extensive multimodal datasets, these models have shown remarkable abilities in tasks like multimodal retrieval [14, 22, 23], visual question answering, image captioning [17], reasoning [20, 21], and retrieval-augmented multimodal generation [24] as well as image generation [25, 26].
The comics domain is uniquely well-suited to driving advancements in these types of multimodal reasoning models. Research in comics has extensively explored a range of questions, from the human ability to derive meaning from sequential images [27] to machine interpretation of comic strips, particularly through closure tasks [8].
With the continuous advancement in AI models and the intensive interest in harvesting vast knowledge from Comics, a lot of work is being published constantly. This is evidenced by a great number of recent papers as shown in Fig. 1. These contributions shifted the landscape around Comics Understanding, leaving previous surveys outdated [6]. Nowadays, the Comics Understanding  research community lacks a comprehensive survey that catalogs these papers and the various challenges in the Vision-Language domain.
To address this need, we introduce in this paper the novel Layer of Comics Understanding (LoCU) framework. This taxonomy, illustrated in Table I, delineates tasks by their input/output modalities and the spatio-temporal dimensions necessary for reasoning over comic data, inviting a structured and progressive approach to comics analysis. Motivated by the uniquely broad task distribution of the comics medium, this survey aggregates and examines all extant research on machine learning in comics, aimed at highlighting critical unresolved issues in the field. It is the first to review the breadth of comics research in a hierarchical structure, and to our knowledge, the first to propose a big-picture taxonomy of comics intelligence from a task-oriented perspective.
Figure 2: Visualization of the Layers of Comics Understanding tasks. From panel level to multipage, from unimodal to multimodal, from simplest to more complex.
Within this framework, the survey enumerates the state of the art in Comics Understanding tasks, including image classification, object detection, semantic segmentation, style transfer, multimodal retrieval, synthesis, and generation. Our comprehensive analysis uncovers fundamental gaps—specifically in data availability and task definition.
In summary, the main contributions of this work are threefold. First, it presents a systematic review of Comics Understanding tasks. This is the first survey of comics tasks for the Vision-Language domain, which provides a big picture of this promising research field with a comprehensive summary and categorization of existing studies. Second, it studies the up-to-date progress of Comics Understanding, including a comprehensive benchmarking and discussion of existing work over multiple public datasets. Third, it shares several research challenges and potential research directions that could be pursued in VLMs for Comics Understanding.
The rest of this survey is organized as follows. Section 2 introduces the Comics medium as a composition of Visual-Language tasks, and several related surveys. Section 3 describes the Foundations of Comics, including their structure, similarities, and differences among various types. Section 4 introduces existing tasks and the commonly used datasets in Comics Understanding, also focusing on evaluation procedures. In Section 5 we propose the Layer of Comics Understanding taxonomy (LoCU), and for each layer a collection of tasks grouped by difficulty (see Annex LABEL:annex:values). In Section 6 we start with Layer 1: Tagging and Augmentation, Section 7 with Layer 2: Grounding, Analysis, and Segmentation, in Section 8 with Layer 3: Retrieval and Modification, Section 9 with Layer 4: Understanding, and in Section 10 with Layer 5: Generation and Synthesis. Finally, we conclude with Section 11.
2 Background
This section first presents the development of the comics research field based on the learning paradigm and how it evolves towards more complex tasks. Then, we introduce the main limitations encountered by the research community, as well as discuss several related surveys to highlight the scope and contributions of this work.
2.1 The Comics Research Epochs
The development of the comics field can be broadly divided into three phases, including (1) Traditional Machine Learning, (2) Deep Learning, and (3) Modern Foundational Models. In what follows, we introduce, compare, and analyze the three paradigms and how they shaped Comics advancements in detail.
2.1.1 Traditional Machine Learning
Early machine learning approaches heavily depended on feature engineering, utilizing hand-crafted features [28, 29, 30], which were capable of, e.g., classify the style of a manga using SVMs [31], retrieve a manga character from a sketch with edge orientation histograms [32] or segment comic balloons with energy functions and active contour models [33]. However, these paradigms require domain experts to craft effective features for specific tasks, which do not cope well with complex tasks and have poor generalizability. Moreover, the feature engineering obstacles the extension of the Comics Understanding to more tasks.
2.1.2 Deep Learning approaches
With the development of deep learning, researchers have achieved major improvements by using end-to-end trainable deep neural networks (DNNs). These models eliminate the need for complex feature engineering and shift the focus to designing network architectures that learn features automatically. Applying these paradigms to comics came as a natural step. Models became better at detection [34, 35] and segmentation [36, 37, 38], and more complex tasks started being investigated and solved (e.g. visual and textual closure [8]). However, the turn from traditional machine learning toward deep learning raises a new grand challenge: the laborious collection of large-scale, task-specific, and crowd-labeled data in DNN training. Moreover, while deep learning methods excel in specific tasks, they are often difficult to adapt to multi-task. In the domain of comics, where data is scarce, protected by copyright, and challenging to annotate, these constraints have significantly hindered the development of more generalized, multi-task models for Comics Understanding.
2.1.3 Modern Foundational models
The era of foundational models marks a significant evolution in the application of AI to comics, with models that can learn from vast amounts of unstructured data to perform a variety of complex tasks [14, 39, 40]. These models relax previous limitations by “solving” simple yet important tasks as a zero-shot counter effect, e.g. panel, character, and text-box detection [41]. By leveraging these foundational models, more advanced tasks can be proposed and takled [42, 43], opening the doors to yet unexplored more complex ones.
TABLE I: Overview of Vision-Language Tasks of the Layers of Comics Understanding. The ranking is based on input and output modalities and dimensions, as illustrated in Appendix B. Taxonomy.
Layer Category Task Input Output
0 (Sec.5) View (Sec. 5.1) Basic Image Viewing (BIV) Text Command Image Display
Image Classification (I-CLS) Image Tag
Emotion Classification Comic Panels/Images Emotion Labels
Action Detection Multiple Panels Tag
Tagging (Sec. 6.1) Page Stream Segmentation (PSS) Images Tags sequence
Image Super-Resolution (ISR) Image Image
Style Transfer (ST) Image Image
Vectorization Comic Panels/Images Vector Image
1 (Sec.6) Augmentation (Sec. 6.2) Depth Estimation Comic Panels/Images Depth Map
Object Detection Tag/s + Image Bounding Boxes
Character Re-identification (Character ID) Multiple Panels Tag
Grounding (Sec. 7.1) Grounding (IG) [Prompt] + Image Bounding Boxes
Character-Balloon Association (Speaker ID) Character + Balloons Tag
Dialog transcription Image Text
Analysis (Sec. 7.2) Translation Image Text
2 (Sec.7) Segmentation (Sec. 7.3) Instance Segmentation (IS) [Prompt] + Image Segments
Image-Text Retrieval (IR) Text Image
Text-Image Retrieval (TR) Image Text
Retrieval (Sec. 8.1) Composed Image Retrieval (CIR) Text + Image Image
Image Inpainting (II) Text + [prompt] + Image Image
3 (Sec.8) Modification (Sec. 8.2) Image Editing via Text (IET) Text + Image Image
Visual Question Answering (VQA) Image + Text Text
Visual Dialog (VisDial) Image + Dialog + Text Text
4 (Sec.9) Understanding (Sec. 9.1) Visual Reasoning (VR) Image + Text Text
Image-2-Text Generation (I2T) Image Text
Grounded Image Captioning (GIC) Image Text + Bbox
Text-2-Image Generation (T2I) Text Image
Scene Graph Generation for Captioning Comic Panel Scene Graph
Generation (Sec. 10.1) Sound Generation from Single Panel Single Comic Panel Sound/Audio
3D Model Generation from Images (3DGI) Collection of Images 3D Model
Video Generation from Text (VGT) Complex Long Text Video
5 (Sec.10) Synthesis (Sec. 10.2) Narrative-Based Complex Scene Generation (NCSG) Detailed Narrative Text Series of Images
2.2 Relevant Surveys
To the best of our knowledge, this is the first survey that reviews comics papers with a task-oriented approach for Vision-Language. Several relevant surveys have been conducted, examining broader aspects of visual art [44] and, more specifically, comics [45, 6, 46]. Notably, Augereau et al. [6] categorized comic research into three primary areas: (i) content analysis, (ii) content generation and adaptation, and (iii) user interaction, which remain relevant entry points for understanding this field’s research and applications. Since those surveys were published, foundational models have emerged, enabling significant advancements, particularly due to their ability to generalize across tasks without task-specific training. These models have unlocked new possibilities in comics research, allowing automatic annotation and analysis, among other tasks.
Building on these advancements, our survey focuses on two key aspects: 1) the recent progress in Comics Understanding tasks and datasets, and 2) the application of Vision-Language tasks to the unique characteristics of comics.
3 Comics Foundations
Our preliminary analysis prioritizes understanding the structural intricacies of comics. The term “comics” broadly denotes the entire medium in its uncountable form, while “a comic” refers to a singular entity within this medium, such as a comic book or strip. The structure of comic books has been a subject of extensive study, examining aspects like their commonalities, layout structures [47], content arrangement, narrative transitions [48, 27, 49, 50, 51], and stylistic narrative variations [52, 53, 54].
Figure 3: Anatomy of Comic Page Elements: This illustration delineates the various components found on comic pages including object detection, linkages between elements, the designated reading order of texts and panels, as well as key textual features like character names—all arranged in a logical sequence throughout the panels and pages.
Typically, comics are presented in single or double-page formats, influenced by the artist’s style. They narrate stories across various genres or series, such as superheroes and science fiction tales, sometimes interlinking through crossovers. Notably, artistic consistency within a series or genre, often achieved through collaborative efforts, is a key feature. Each comic page (as illustrated in Figure 3) encapsulates individual scenes within panels of diverse types - from splash panels for opening scenes to widescreen panels for expansive narratives. These panels convey interactions and dialogues through speech balloons, thought bubbles, and expressive onomatopoeia, enriching the narrative experience. Artistic variations in style and color also play a significant role in comics, ranging from detailed realism to abstract forms, with the color palette setting the mood and atmosphere.
Despite rigorous analyses [55], no significant differences have been identified in terms of attention, subjectivity, and viewpoint between Manga and American comics. Hence, these distinctions are better explored within their distinct historical, cultural, and artistic contexts:
American Comics: With a rich history dating back to the 1930s, American comics are celebrated for their colorful and high-quality presentations, primarily known for their superhero narratives. They adhere to a top-to-bottom, left-to-right reading format and are characterized by a dialogue-rich and brisk storytelling style.
Manga: Emerging in post-World War II Japan and influenced by diverse sources, including Japanese traditional art and American comics, manga is recognized for its unique black-and-white style, though colored variants exist. Covering a broad range of genres, Manga is read right-to-left, top-to-bottom. This style is marked by a distinctive “Japanese Visual Language” that is consistent across genres [56].
Comic Formalisms. Comics are defined as sequential visual narratives that utilize a combination of images and text within distinct frames (panels) to convey a story or information. Each comic, or manga, can be represented as a document $D$. Each document $D$ consists of a series of pages $P=\{p_{1},p_{2},\ldots,p_{n}\}$. A page $p_{i}$ can be described by its layout and the elements it contains: (i) a spatial arrangement of panels $\{c_{1},c_{2},\ldots,c_{m}\}$, also known as Layout; (ii) each panel $c_{i}$ is a bounded area containing images and possibly text; panels are the primary unit for storytelling in comics; (iii) gutters: the space between panels, often used by artist to leave readers imagining the missing piece of the story, which helps in sequencing and timing; and (iv) text elements, including balloons (speech or thought), captions, and sound effects (onomatopoeias). Each panel $c_{i}$ contains characters (depicted through various visual styles and expressions), backgrounds (visual scenery behind the main characters or action), and foregrounds (objects and elements in front of the characters).
4 Tasks and Datasets
This section summarizes the commonly used Comics datasets, as detailed in Table II. Some of these datasets are either intended for training or evaluation, depending on their sizes.
TABLE II: Overview of Comic/Manga Datasets and Tasks, their information (availability, published year, source), and properties (languages, number of comic/manga books and pages). The table rows are repeated according to the supported Tasks. Accessibility is indicated with for no longer existing datasets, indicates existing but not accessible, and means existing and accessible. The link [proj] directs to the project websites, while [data] directs to dataset websites. For CoMix, mix means that it inherits from a mixture of four datasets [57, 8, 58, 41]. For VLRC, Analysis indicates annotations are about statistical analysis of panel size, color, layout, etc. However, the dataset does not provide images and annotations have been gathered directly on physical books.
Task From Name Year Access. Language Origin # books # pages
Image Classification Sequencity [59] 2017 EN, JP - - 140000
BAM! [60] 2017 - - - 2500000
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
[8] EmoRecCom [62] [proj] [data] 2021 EN 1938-1954 - -
Object Detection Fahad18 [63] [proj] 2012 - - - 586
eBDtheque [57] [proj][data] 2013 EN, FR, JP 1905-2012 25 100
[57] sun70 [64] 2013 FR - 6 60
COMICS [8] [proj] [data] 2017 EN 1938-1954 3948 198657
BAM! [60] 2017 - - - 2500000
[61] JC2463 [34] 2017 JP - 14 2463
AEC912 [34] 2017 EN, FR - - 912
GCN [65] [proj] [data] 2017 EN, JP 1978-2013 253 38000
[59] Sequencity612 [59] 2017 EN, JP - - 612
[57] SSGCI [66] [proj] [data] 2016 EN, FR, JP 1905-2012 - 500
[61] Comics3w [67] [proj] 2017 JP, EN - 103 29845
comics2k [68] [proj] [data] 2018 - - - -
DCM772 [58] [proj] [data] 2018 EN 1938-1954 27 772
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
BCBId [69] [proj] [data] 2022 BN - 64 3327
[61] COO [4] [proj] [data] 2022 JP 1970-2010 109 10602
[8] COMICS-Text+ [70] [proj] [data] 2022 EN 1938-1954 3948 198657
PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Re-Identification Fahad18 [63] [proj] 2012 - - - 586
Ho42 [71] 2013 - - - 42
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Linking eBDtheque [57] [proj][data] 2013 EN, FR, JP 1905-2012 25 100
[57] sun70 [64] 2013 FR - 6 60
GCN [65] [proj] [data] 2017 EN, JP 1978-2013 253 38000
[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Segmentation [59] Sequencity4k [38] 2020 EN, FR, JP - - 4479
Dialog generation PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
Analysis VLRC [47] [proj] [data] 2023 JP, FR, EN, 6+ 1940-present 376 7773
4.1 Annotations overview
The available current comics datasets were introduced for different tasks. They vary in the year of release, the language and the origin. This is shown in Table II. From the table, it is noticeable that the available datasets vary significantly in annotation types, making them cover different tasks.
eBDtheque [57] offers comprehensive annotations in high quality for Character, Balloon, and Panel detection, along with Text-Character association. This makes it usable for object detection and linking tasks but it comes with a drawback: it is composed of only 100 pages. In contrast, the COMICS dataset [8] focuses on a set of tasks called closure, based on automatic panel and text detection over a large number of pages. However, it was introduced with poor annotations. Only recently some works [70] have improved OCR text detection on the COMICS dataset. In the same vein, [72] shows that improving OCR accuracy (using Textract) could lead to marginal accuracy gain in closure tasks for the COMICS dataset. Both [70, 72] released the novel OCR annotations to the research community.
It is to note also that some of the dataset with high-quality annotations and large number of pages are not accessible. For instance, BCBId [69], though promising high-quality annotations, falls short in delivery. Similarly, VLRC [47], while rich in annotations, lacks digital applicability due to its focus on physical books.
Manga109 [61] is the first to provide high-level annotations, covering more than 20k pages. The dataset has been improved over several iterations, further expanding the scope of the dataset by adding annotations for onomatopoeias [4] and dialog [73]. The actual state of Manga109 is the result of many years of annotation and improvement by subsequent developments. More recently, following the Manga109 wave, PopManga [41] has been proposed as an English-language alternative, including annotations for recognition and linking, but on a smaller scale (2k images).
In summary, while datasets for Comics Understanding exist, there is a need for more approaches to standardize and unify their annotations. The Comics Dataset Framework [2] proposed a unification of detection annotations as well as a standardization of metrics and settings used to train and measure models. The work was subsequently extended in CoMix [42], where mainly English-language comics (and manga) were selected and the annotations extended to the task of single-page dialogue transcription. In this work, performances of several state-of-the-art models, including Magi [41], CLIP [14], DINOv2 [39], and GPT4 [74], were measured, for the first time.
However, as highlighted in the first column of Table II, the complexity of annotations and tasks is limited. Object detection (panel, characters, texts), text-to-character linking, and re-identifications are integral tasks to understanding the narrative and dialogue in comics. These tasks, along with more recent developments, will be examined in the context of our taxonomy in the following Section.
4.2 Datasets overview
The domain of Comics Understanding encompasses a range of datasets focused on processing and examining comic media. A review of various datasets, as detailed in Table II, reveals that several are no longer accessible (marked with ) or have restricted access (). Those available () often require approval from respective research groups in order to obtain them. This is a common practice in the field of “sharing copyrighted materials” as in fact they can be used only for research purposes in the majority of cases. If the copyright does not allow for redistribution, it is a common practice to release annotations and a script to download and structure the dataset accordingly (as for [41]). As we can see from Table II, many datasets share common origins (column “From”). Notably, these datasets vary significantly in terms of style, quality, and size. The common languages of these datasets are English, French, and Japanese and none of the available data seems to be collected from recent books (after 2010), probably due to copyright limits. We found the above limitation to be one of the main issues in the comic domain: (almost) every work proposes its own (sometimes private) dataset on which results are drawn with little comparison, making these works hard to replicate. Datasets are limited in size and the model’s code and weights are rarely shared. Luckily, this has started changing recently [2, 41, 42].
TABLE III: Number of single instance annotation (declared) in papers. The “Obj. cls” stands for Object-based classification, and refers to the number of classes available if one wants to classify the detected instances in fine-grained classes.
Mode Name Panel Char. Face Textbox Onomat. Obj. cls
Color Fahad18 [63] - 18 - - - 18
Color Ho42 [71] 200 100 - - - -
Color eBDtheque [57] 850 1550 - 1092 - 5
Color SSGCI [66] - 50 - - - 1
Color COMICS [8, 70] 1229664 - - 2498657 - 2
BW Comics3w [75] 159529 - - - - 1
BW JC2463 [34] - - 15801 - - 2
Color AEC912 [34] - - 8184 - - 2
Color DCM772 [58] 4470 10757 5438 - - 4
BW Manga109 [61, 35, 4] 103900 157152 118715 147918 61465 4
Color EmoRecCom [62] 10199 - - - - 8
BW PopManga [41] - 18783 - 20843 - 2
mix CoMix [42] 22176 48903 32625 37923 - 4
In Table III we report, for every dataset, the declared numbers of annotations available. On the right-most column, we report the number of classes annotated, which correspond to different class objects for the detection, or the number of classes for classification annotations. For instance, in Fahad18 there are 18 characters, thus the number of classes is 18. However, in eBDtheque, there are 5 classes, but only panel, character, and face annotations. This is because the characters are separated into three sub-categories: human-like, animal-like, and object-like. Here, we describe the datasets chronologically.
eBDtheque: The eBDtheque [^1] http://ebdtheque.univ-lr.fr/registration [57] is a collection of 100 pages over 20 books, selected from J-comi222previously http://www.j-comi.jp, now no longer available., mainly in the French language (with also English and Japanese comics). [76] led a huge effort in labeling 850 panels, 1092 balloons, 1550 characters, and 4691 text lines, provided by domain experts. For balloons, they provide instance segmentation and classification ( speech, thought, narration, illustrative).
Manga109: The Manga109333http://www.manga109.org/index_en.php [61] dataset contains 109 manga volumes from 93 different authors. Authors successfully obtained the copyrighted material from manga drawers with a “restrictive term-of-use” which implies redistribution only with the scope of research and the possibility to use them in publication with proper referencing the author. These mangas were published between the 1970s and 2010s and are categorized into 12 different genres such as fantasy, humor, sports, etc. This dataset has been further extended by COO [4] which added onomatopoeias polygons annotations and links among truncated onomatopoeias. In fact, on average there are 5.8 onomatopoeias per page in Manga109, while COMICS has much fewer (0.09). Another set of annotations comes from Manga-Dialog [73] which added a text-character link for every speaking text box.
COMICS: The COMICS [^4] https://obj.umiacs.umd.edu/comics/index.html [8] contains 1,229,664 panels paired with automatic textbox transcriptions using an old-OCR system. The 3,948 books are gathered from the Digital Comics Museum (American golden-age comics books) published between 1938 and 1954. The dataset includes ground truth (manually annotated) labeled data such as the rectangular bounding boxes of panels on 500 pages and 1,500 textboxes. Subsequent works have expanded original annotations with better OCR transcriptions [70, 72]. In particular, [70] developed a pipeline for OCR processing and labeling of comic books from COMICS, while [72] extracted all OCR transcripts from COMIC using new-generation OCR textract [^5] Used through API at https://docs.aws.amazon.com/textract/. Lastly, [77] includes strips of dialog prediction based on “persona” information.
DCM772: The DCM772666https://git.univ-lr.fr/crigau02/dcm_dataset [58] contains 772 (pages) images from 27 golden-age comic books from the same COMICS source. Annotations, however, are more precise as they contain panels, the associated characters bounding boxes, and also face boxes.
Figure 4: Example of different comic datasets, from Black-and-White to Color, from comics-style to Manga-style.
4.3 Datasets for Evaluation
Across the datasets presented in Table II, two recent datasets are provided as evaluation sets. In particular, PopManga proposes two splits (validation and test) while CoMix enhances annotations for various existing datasets among which also PopManga.
PopManga: The PopManga [^7] https://github.com/ragavsachdeva/Magi [41] dataset contains English manga titles from the most popular mangas. The dataset contains two test splits: test-seen with 1,136 pages and test-unseen with 789 pages, obtained from 15 and 10 books respectively. The test sets have been annotated with Text (both textboxes and onomatopoeias annotated as text), and Characters.
CoMix: The CoMix [^8] https://github.com/emanuelevivoli/comix-dataset [42] dataset contains English and French comics and manga titles from the most popular datasets: eBDtheque, COMICS, DCM, and PopManga. The dataset contains two splits: validation (available with annotations) and the held-out test split (available only through the evaluation server [^9] Accessible at Robust Reading Competition website). It contains 3.8k images across 100 books, and annotations for multi-task, from detection and linking to dialog generation.
5 Taxonomy: LoCU
Developing a comprehensive taxonomy for comics within the Vision-Language (VL) domain requires a careful overview of existing work in VL surveys and tasks. Our approach synthesizes insights from various benchmarks and taxonomies, forming the foundation for our unique classification – the Layers of Comics Understanding (LoCU).
Relevant Taxonomies. Initial inspiration stems from the breadth and depth of multimodal questions covered in recent benchmarks like MMMU [78], highlighting the necessity for a taxonomy that accounts for modal complexity and task-specific intricacies in comics. This need is further echoed in the domain-specific taxonomy proposed for video understanding in [79], particularly their dimensional analysis approach, which informs our consideration of spatial and temporal domains in comics. Shifting focus to VL multimodal evaluation, [80] reframes VL tasks as Visual Question Answering (VQA), emphasizing a range of abilities from grounding to numerical calculation. While this framework offers valuable categorizations, it lacks a comprehensive structure for VL tasks, particularly those pertinent to comics. Similarly, the multimodal challenges outlined in [81] offer insights into the representation, translation, alignment, fusion, and co-learning challenges in multimodal learning. These challenges guide our approach to defining taxonomic classes and subclasses within the VL domain, though with a more task-centric focus. Drawing further on [82], we observed the classification of VL tasks into grounding, retrieval, understanding, and generation, with specific adjustments and additions to accommodate comic-specific tasks. Our taxonomy expands on this structure, adding categories such as tagging, augmentation, analysis, segmentation, modification, and synthesis, addressing the wider range of tasks observed in comic studies.
In Table I, we present 31 Vision-Language multimodal tasks, categorized according to the modalities involved, forming the backbone of the LoCU framework. This taxonomy, spanning 10 distinct task groups across five layers, aims to capture the multifaceted nature of Comics Understanding. Layer 0, representing fundamental image processing and viewing capabilities, sets the stage for increasingly complex interactions in the subsequent layers. More details are provided in Appendix B. Taxonomy. This structured approach allows for a comprehensive and nuanced understanding of the range and complexity of tasks within the comics domain, laying a foundation for future exploration and innovation in comic analysis and understanding.
6 Layer 1: Tagging and Augmentation
The first layer comprises tasks with unimodal input (single image or sequence of images) and unimodal output (Images or Text): Tagging and Augmentation.
6.1 Tagging
Tagging tasks predominantly involve classification outputs, albeit with varying input types.
6.1.1 Image Classification
Definition: The task of Image Classification requires classifying an image into one of several predefined categories. In the comics domain, the task can be applied at panel or page level. In both cases, the classes vary from artist names, comic/manga styles or the image type (whether it’s an ads or front page or story page).
Early efforts in this domain, such as those by Chu et al. [31], focused on classifying manga panels by artistic style using Support Vector Machines (SVM) and manga-specific feature vectors. Their feature vectors are built on edge-detected lines creating a 20-dimensional vector from elements such as angle, orientation, the density of segments, etc. Hiroe et al. [83] employed a novel method using SVMs to classify comic books based on the frequency of exclamation marks, demonstrating marks-per-page histogram as an innovative approach to content approximation.
Daiku et al. [84] expanded the scope to page-level classification within manga stories using Convolutional Neural Networks (CNN), distinct from Page Stream Segmentation (discussed in Subsection 6.1.4). Jiang et al. [85] introduced the Consensus Style Centralizing Auto-Encoder (CSCAE) for style classification, particularly distinguishing between Shonen (for boy) and Shojo (for girl) manga styles using robust style feature representations. They used low-rank and group sparsity constraints for consensus for the classification of manga character faces. Lastly, in [86] Terauchi et al. presented an approach for classifying manga based on the author’s unique style, employing a Variational Autoencoder (VAE) for this purpose. They also proposed a “four-scene comics story dataset” for this classification task into Mow, Seinen, and Shonen touches (styles).
In a more recent work [87], the authors tackled the challenge of (multi-)genre classification at the story level in Manga. They introduced a Panel-Page-aware Comic genre classification model, which takes page sequences as input and produces class-wise probabilities, relying on the attention mechanism to jointly attend page features and panel boxes, then pooled into a transformer encoder. The transformer classification token is merged with a Graph Convolution Network (GCN) over the labels graph, which predicts the probability distribution of the labels for the comic stories.
6.1.2 Emotion Classification
Definition: Emotion Classification in comics involves categorizing a single panel or an entire page (looking both at the image and the text) into predefined emotional categories, enhancing the understanding of the narrative’s emotional context.
A seminal study by Tanaka et al. [88] explored the interplay between speech balloon shapes and emotions in comics. They analyzed the relationship between balloon styles and the emotions expressed within them, using the Manga109 dataset as a foundation for their study.
The EmoReCon challenge at ICDAR 2021 [62] marked a significant advancement in this field. Participants were tasked with classifying 8 emotions from comic panels, using a dataset specifically designed for this purpose [62]. Notably, one of the winning teams employed a three-level fusion approach (early, mid, and late), integrating EfficientNetB3 [89] and RoBERTa [90] for visual and textual embeddings, respectively. This mid-fusion strategy outperformed other architectural configurations, signifying the potential of multimodal fusion. The second winning team proposed an early fusion approach with ResNet50 image features and OCR-extracted text as input for a transformer encoder BERT-style. In this case, the prediction is drawn from a weighted average over the layers’ cls tokens, with comparable results with the previous method.
Not only have the emotions of comic characters been analyzed, but in [91] authors studied readers’ emotions through physiological signals. Through the relationship between a change in the reader’s physiological signal and the page of the reading, authors classify the manga across three classes: comedy, romance, and horror.
A recent work [92] introduced two noteworthy datasets derived from Manga109: MangaAD+ and MangaEmo+. While MangaAD+ is a curated collection of 600 pages featuring unannotated elements such as shop names and ads - in Manga109 only the balloon text and onomatopoeias are annotated -, MangaEmo+ extends this dataset by annotating each page with a range of 8 emotions for multi-label classification. They employed a modified version of Faster R-CNN, trained on augmented text data, for atypical text detection. The subsequent emotion classification task was approached using EfficientNet, which extracted visual features combined with the above custom text detection model. However, neither of the two datasets has been made publicly available.
6.1.3 Action Detection
Definition: Action Detection in comics is characterized by the analysis of sequences of consecutive panels. This task involves processing multiple images to classify the depicted sequence into one of several predefined action categories.
Unique to Action Detection is its multimodal input nature, where each panel in the sequence contributes to a holistic understanding of the action. Unlike single image analysis, this task requires the model to interpret a series of images, each adding context and continuity to the narrative. Despite being popular for Videos [93], this task has not yet been explored in comics apart from image-cloze task [8] where the model is challenged to select the correct next panel from two options. This task, while elementary, underscores the potential for more sophisticated Action Detection in comics, a domain where sequential imagery plays a pivotal role.
6.1.4 Page Stream Segmentation
Definition: Page Stream Segmentation (PSS) in comics entails segmenting a comic book into different stories or sections. In its simplistic version, the task can be seen as categorizing each page with a tag among cover, regular page, ads, or last page. In its more advanced form, it includes extracting salient information from the stories such as title, authors, drawer, price, characters, and synopsis.
The concept of PSS, while not new in document analysis, has seen various methodologies employed over time. Traditional approaches, such as those utilizing Support Vector Machines (SVM) for single page classification [94], have evolved with the advent of Convolutional Neural Networks (CNNs) [95], and more recently, attention-based multimodal fusion techniques [96]. These methods have been pivotal in processing and understanding page layouts and structures in various document formats.
Despite its longstanding presence in document analysis, PSS remains largely uncharted in the comics domain. Comics present a unique challenge for PSS due to their complex structure; a single volume can interweave multiple stories, interspersed with advertisements and narrative text. This intricacy demands a nuanced approach to segment and understand the diverse content within comic books.
A potential advancement in PSS for comics involves not just segmenting the pages but also tagging these segments with metadata like story titles, authors, and synopses. This represents a promising field of study, extending traditional image classification into a more intricate, story-centric context.
6.2 Augmentation
Augmentation tasks, in the context of comics, involve image manipulation techniques aimed at enhancing or altering the visual presentation. The primary tasks identified under this category are Super-Resolution (SR), Style-Transfer (ST), Vectorization, and Depth Estimation. Each of these tasks revolves around processing a single image input to produce an enhanced or stylistically altered output.
6.2.1 Super-Resolution
Definition: The Super-Resolution task in comics primarily targets the enhancement of image quality, especially for images that are compressed, photograph-based, or scanned.
Recent advancements in manga super-resolution (SR) focus on preserving visual integrity during enhancement. Yao et al. [97] employ deep learning to classify and tailor SR models for manga screentone, ensuring semantic preservation. Dai et al. [98] introduce the Structured Fusion Attention Network (SFAN), which optimizes feature extraction and reconstruction quality through attention modules, achieving superior performance on datasets like Manga109. These methods enhance image quality while maintaining the artistic nuances critical to manga, effectively balancing technical efficiency with artistic integrity.
6.2.2 Style-Transfer
Definition: Style-transfer in comics encompasses a range of image-to-image modifications, including photo-to-comic transformation, image-to-image translation, and colorization, characterized by similar input and output attributes.
Photo-to-Comic Transformation: Photo-to-comic has mainly involved mangas and involves replacing colors and textures with halftone patterns. Traditionally, it often consists of manually selecting and adjusting the colors. Methods have evolved from basic pattern generation to sophisticated techniques capturing structural and tonal similarities. Classic approaches relied on algorithmic solutions and were limited in their ability to handle multiple screentone. Recent advancements by Zhang et al. [99] utilize deep learning for semantic region segmentation in cartoon images, followed by screentone application. However, these methods are primarily effective for color-based inputs, with limited applicability to line-based images.
Image-to-Image Translation: Topal et al. [100] explored style transfer as a transfer for the detection task. Specifically, they used style transfer networks like CycleGAN and CartoonGAN to train a detection network, which was then fine-tuned on specific comic datasets for character and face detection. This approach, to use style transfer techniques to enhance performances in other tasks, is not new [68, 101].
Figure 5: Colorization vs. Screening
Colorization: Colorization in comics, specifically in manga, represents the task of transforming black-and-white images into full-color pages. It is a complex process that requires a balance of artistic understanding and technical precision.
Furusawa et al. [102] introduced a comprehensive approach to the colorization of manga. Their method allows for user interaction, where color dots can be added to guide the colorization process. This approach not only automates the colorization but also gives users the flexibility to influence the final output, ensuring that the automatic colorization aligns with the intended aesthetic vision. In a different approach, Hensman et al. [103] presented a model based on Conditional Generative Adversarial Networks (cGANs). Their model, specifically tuned for each reference image, colorizes manga panels to mimic colorization styles. This method implies training a new GAN for each reference image, making it versatile for different styles but requiring significant computational resources for each new style adaptation. Subsequently, in [104] authors investigated screentone synthesis, a critical aspect of manga art that involves translating line drawings into screentone-patterned images. They utilized Manga109 for comparison with standard image-to-image translation methods like pix2pix [105]. Their approach, although similar in architecture to U-Net used in previous studies [106], stands out by explicitly considering screentone labels, thus preserving the integrity of screentone patterns in the final output. Moving towards a more advanced solution, Golyadkin et al. [107] introduced a two-stage architecture for manga page colorization. The first stage involves drafting the coloring using the Pixel2Style2Pixel architecture [108], followed by a conditional GAN in the second stage to refine the output. This method elegantly fuses color drafts with color hints from users, generating high-quality colorizations. They demonstrated improvements both qualitatively and quantitatively, indicating a significant advancement in manga colorization technology. Building upon previous methods, Jiramahapokee et al. [109] further refined the colorization process. They leveraged a multi-encoder Variational Autoencoder (VAE) for improved color region consistency and combined it with the shading strengths of other tools like manga-colorization-v2. Their approach also incorporated CIELAB [110] interpolation to enhance color saturation and authenticity, leading to outputs that are not only aesthetically pleasing but also consistent and detailed in their shading quality.
Screening: Screening, or adding screentone to black-and-white manga drawings, has been a focus of several studies. A notable contribution by Wu et al. [111] introduced a two-stage architecture for this process. The first stage involves generating grayscale shading using a six-layer Swin-transformer [112], which then serves as input for the second stage, a reference-based screentone generation module. This algorithmic process extracts screentone patterns from a reference manga and applies them to line drawings, considering the shading nuances provided by the first stage.
These advancements in colorization represent a significant contribution to the domain of comics, particularly manga, where colorization can profoundly impact the narrative and aesthetic appeal of the stories.
6.2.3 Vectorization
Definition: Vectorization in the context of comics, particularly manga, involves transforming raster images into vector formats. This process includes operations like screening, structural lines extraction, and converting images into vector graphics, each with unique methodologies and applications.
Structural Lines Extraction: This process, structural lines extraction, focuses on removing screentones to leave only the defining lines of figures and shapes. It can be seen as the reverse operation of the screening, which corresponds to adding the screentone to a structural line image. In this task, Li et al. [106] developed a CNN-based method using a U-Net architecture with residual blocks to effectively extract structural lines from manga pages. This method is adaptable for various applications, including manga retargeting and colorization, and can handle diverse patterns and textures, resulting in clean skeleton sketches.
Manga Vectorization: Various works [113, 114, 115, 116] have proposed different approaches for this task. In particular, Yao et al. [117] developed a method involving adaptive binarization and screentone detection, refining borders, estimating lighting, and compensating for missing strokes, leading to high-quality, resolution-independent rendering. A recent innovative approach, MARVEL by Su et al. [118], employs Reinforcement Learning to train an agent in selecting drawing primitives that recreate raster manga images in vector format. While effective, this approach works with squared manga patches and may result in visual gaps in certain renderings. Moreover, this method is not convenient in shape rendering, as it uses primitives and cannot manipulate them. MARVEL represents an interesting exploration into the use of machine learning for vectorization, highlighting both the potential and the complexities of this task.
6.2.4 Depth Estimation
Definition: Depth Estimation within the comic domain refers to the process of predicting the spatial distance between objects and the viewer in a two-dimensional comic panel. It serves a crucial purpose in the reconfiguration of comic panels and unlocks potential developments in digital artistry.
Up to now, this necessity arose particularly when adapting comics for various digital media, such as smartphones, tablets, and e-readers, where the original layout may not fit the display dimensions optimally. The primary methodologies for reconfiguring comic panels involve: (i) Adjusting panel proportions and rescaling them to fit the new medium without compromising the narrative or visual integrity. (ii) Strategically cropping panels while ensuring that key objects within each panel are preserved and remain visible in the adapted format. Both techniques rely heavily on depth estimation to identify and prioritize the most significant elements in each panel.
One significant challenge in this domain is the lack of ground truth data for comics. To address this, Bhattacharjee et al. [7] adopted an innovative approach using Image-to-Image style-transfer. They transformed comic images to resemble natural images, upon which depth estimation techniques, involving Laplacian edges and feature-based GANs, were applied to predict depth effectively. Building on this, a subsequent study [119], based on their earlier work [120], introduced a Swin Transformer for multitasking, including both depth estimation and object segmentation. This multi-decoder-head architecture, while effective, presents challenges in scalability due to the complexity of accommodating multiple tasks.
Depth Estimation plays a pivotal role in adapting comics to the varying formats of digital consumption. By ensuring that key visual elements are retained and appropriately displayed, these methodologies not only enhance the reader’s experience but also preserve the artistic intent of the original work.
6.3 Satellite Tasks
Some of the tasks that fit under the Augmentation umbrella have not found a space in our taxonomy, e.g. the task of de-warping. The task involves correcting distortions typically found in scanned comic pages. Addressing the de-warping challenge, a recent study by Garai et al. [121] proposed a novel mathematical model to describe the warping process. This model estimates distortion factors based on the boundaries of panels in a comic document image, aiming to rectify these warps effectively. While their approach yielded promising qualitative results, particularly in the image areas, challenges persist in handling text distortion, where some distortions proved unrecoverable.
7 Layer 2: Grounding, Analysis, and Segmentation
The second layer of our LoCU framework comprises three distinct task groups: grounding, analysis, and segmentation. These tasks dig deeper into the intricate components of comics, focusing on detailed elements like panels, text, and characters and their ordering and associations.
7.1 Grounding
Grounding in comics involves identifying and classifying specific elements within a comic’s panel, such as text, characters, and objects. Under this task name umbrella, many individual tasks belong such as Detection, Character Re-identification, and sentence Grouding.
7.1.1 Detection
Definition: In Detection task, given an image and a set of class names (or tags), the goal is to find all instances of those classes in the image and locate them with bounding boxes.
Panels. Initial studies in digital Manga, such as those by Ponsard et al. [122] and Arai et al. [123], emphasized automated segmentation and panel detection, underscoring the importance of maintaining reading order. Arai et al. [124] further developed these ideas with a focus on real-time detection and extraction techniques, which evolved into more sophisticated SVG-based approaches as explored in [125]. [126] contributed advanced methods for panel extraction using region growing techniques and mathematical morphology, combined with speech balloon detection.
Text. Jomaa et al. [127] employed a tracking algorithm for panel extraction, while speech balloons were identified using Robert’s edge detection operator and classified with thresholding. Liu et al. [128] focused on balloon extraction based on text character locations. Pal et al. [129] performed text extraction using SIFT local features, Space Pyramid Matching (SPM), and SVM on eBDtheque and Bangla comics images. In [130] authors contributed to text extraction in Manga109 combining region proposal generation with SVM classifiers. Chu et al. [131] and Qin et al. [34] modified existing architectures like Faster R-CNN for detecting text and faces in comics. In particular, [131] proposed additional aspect ratio regions, while [34] built two comics datasets, namely JC2463 (Manga black and white style) and AEC912 (colored American comics style) to test face detection evaluation.
Face. Interestingly, in [132] authors benchmarked Fast R-CNN, Faster R-CNN and SSD in panel, character, face, and balloon detection. They used the Manga109 dataset, discovering that Fast R-CNN works better in panels when there are clear boundaries, Faster R-CNN when boundaries are not well defined, and SSD needs already cut panels as it is limited for full-page complex mangas.
So far, many methods have used classical detection architectures and applied them to the comic domain to extract elements such as panels, characters, text, etc. A new direction is drawn by [133] challenged the conventional four-edge polygon representation of comic panels, suggesting segmentation-based approaches using U-Net for more accurate extraction. Here, authors performed extraction through segmentation, employing a U-Net that classifies pixels into background, panels, and borders.
Onomatopoeia. Recently, the focus has shifted to onomatopoeia detection, a relatively unexplored area in comic analysis. The release of the COO dataset by Baek et al. [4] has paved the way for new research in this domain. Authors proposed a modified version of M4C [134] designed to link truncated onomatopoeias together. Following this, Yang et al. [92] proposed text augmentation techniques for onomatopoeias using EfficientNet.
Sharma et al. [135] furthered this research by introducing their dataset for panel and character detection, employing a Faster R-CNN model. However, they did not provide comparative analysis with other models.
Benchmarks. Lately, Dutta et al. [136, 69] introduced the BCBId dataset, benchmarking existing datasets (eBDtheque, Manga109, and DCM) and employing YOLO architectures for panel and character detection. They presented comparisons with existing methods, showing enhanced performance in detecting panels and characters across multiple datasets. However, it is unclear if these results are truly comparable. This is due to the fact that in previous (and contemporary) works, authors don’t provide neither information about the settings used (in terms of train-val-split), nor about hyperparameters and model settings to replicate the weights. Only recently, Vivoli et al. [2] fairly compares models across various datasets and styles, unifying annotations for multiple detection classes and providing model weights, code, and data to replicate.
7.1.2 Character Re-Identification
Definition: Character Re-Identification, also known as character retrieval in comics analysis, has evolved from focusing on visual features alone to incorporating textual information alongside character images.
Initial studies by [63, 29] employed techniques like PHOG and SIFT keypoints for character retrieval. Iwata et al. [137], inspired by [138], adapted these methods specifically for manga characters, considering characters rather than entire page ROIs. Authors further explored applications of these techniques in [139]. However, recently, Li et al. [5] noted the limitations of handcrafted features, particularly their lack of robustness to variations in pose, expression, and character shape.
In [5], authors tackled manga character retrieval and verification. While retrieval is a known task 8.1, verification refers to the task of classifying whether two images (a pair) belong to the same character or not. These two tasks are challenging since the manga character images have a long-tailed distribution and large quality variations. While addressing challenges posed by long-tailed distribution and quality variations in manga characters, they proposed a novel dual loss approach, combining dual ring loss and dual adaptive re-weighting loss, to overcome biases introduced by imbalanced training data.
In animated cartoons, where characters undergo shape and color changes, [140] introduced an approach based on Self-Supervised Learning (SSL) and Multi-object Tracking (MOT) to identify character clusters and link character proposals, facilitating character re-identification across different scenes.
Zhang et al. [141] identified challenges unique to manga, such as similarities in faces but differences in bodies. They proposed the Face-body and Spatial-temporal Associated Clustering method (FSAC), which uses face-body graphs and spatial-temporal relationship correction to refine clustering and overcome artistic exaggerations and deformations, through the design of a temporal-spatial-related triplet loss to fine-tune the clustering.
Recognizing the challenges of occlusion in manga, Zhang et al. [142] developed Occlusion-Aware Manga Character Re-identification (OAM-ReID) with self-paced contrastive learning. By synthesizing data with occluded speech balloons and incomplete bodies, their framework becomes more effective in learning representative features of characters. They introduce a combination of contrastive, occlusion, and re-identification losses, training a vanilla ViT and TransReID as baselines on 10 Manga109 books.
Soykan et al. [143] leveraged SimCLR [144] with various data augmentation techniques for character re-identification, calling it “Identity-aware SimCLR”, focusing on both local (face) and global (body) representations. Similarly, Sachdeva et al. [41] proposed a multi-task Transformer model (Relationformer [145]) for character clustering and linking detected character bounding boxes at the page level.
Finally, [146] introduced an improved YOLOv5 and XGBoosting approach for object detection, achieving state-of-the-art results across multiple datasets (StyleObject7K, ClipArt1K, Watercolor2K, and Comic2K). Their method highlights the effectiveness of well-annotated training data and extensive data augmentation in detection tasks.
7.1.3 Sentence-based Grounding
Definition: Sentence-based grounding is a Vision-Language task that links textual content to specific visual elements within a panel. In comics, this task translates into grounding the sentence elements both into a single panel and a full page. Moreover, the complexity of the task varies depending on what is the sentence to be grounded: a panel description or a dialogue or narrative description.
In image and video Multimedia processing various works tackle the task with custom CNNs and LSTMs [147] approaches, in an unsupervised manner [148] or with custom multimodal fusion approaches [149, 40].
However, the only work that used a similar approach (GroundingDINO) in a zero-shot setting is [41] which used it to automatically annotate the training set.
7.2 Analysis
Analysis tasks involve a deeper examination of the relationship between text and visual elements in comics, including character-text association, panel sorting, dialog transcription, and translation.
7.2.1 Text-character association
Definition: The task of text-character association in comics aims to link textual content, usually dialogue, to the corresponding speaking character. This task often involves a two-step process: first identifying panels, text boxes, and character boxes, and then associating the text with its speaker.
Rigaud et al. [150] tackled this task by using geometric graph analysis and anchor point selection. They developed a distance-based algorithm, both global (across the page) and single-frame (within a panel), relying on Euclidean distances and balloon tail information.
Advancing from Rigaud et al.’s work, Nguyen et al. [151] employed a multitask learning approach with a custom Mask R-CNN (adding a new “PairPool” head for the association prediction) which they called ComicMTL. Their method simultaneously addressed panel and character detection, balloon segmentation, text recognition, and text-speaker association, demonstrating its effectiveness on the DCM772 and eBDtheque datasets.
In [152], authors introduced a novel perspective by considering the continuity of comics. Their work focused on estimating the reading order of frames and texts, vital for associating text with off-panel speakers. However, their algorithmic approach raised concerns regarding scalability and generalizability.
Li et al. [73] extended the Manga109 dataset with the Manga109 dialog dataset, enabling more precise speaker detection. They classified text-character links into “easy” (character in the panel) and “hard” (character not in the panel) configurations, using Faster R-CNN for detection and subsequent association classification. The challenges in this approach, when balloons are spoken alternatively, highlight the potential role of NLP models in handling complex dialogue scenarios.
In a recent study, Sachdeva et al. [41] proposed a multi-task Transformer-based model (Relationformer [145]) with specialized MLP heads for detection and linking tasks. While their model offers an “end-to-end” solution, it operates in two iterations: first for detection and then for linking characters and text. However, their model’s comparison with state-of-the-art methods, particularly in detection, is limited, highlighting the need for more comprehensive benchmarking. In fact, the model does not compare with previous state-of-the-art models (YOLO-based) for detection, but only with publicly available GroundingDino [40], in a zero-shot setting.
The character-text association is a critical but challenging task in comics analysis, requiring a nuanced understanding of narrative structure and visual-textual interplay. Future research may benefit from integrating advanced NLP techniques to better interpret complex dialogue scenarios. Additionally, there is a growing need for unified benchmarks and replicable methodologies, as indicated by the diverse approaches in the LoCU framework, to facilitate comparative analysis and further advancements in this field.
7.2.2 Panel Sorting
Definition: Sorting individual comic panels into their correct narrative order presents a unique challenge, both for humans and automated systems.
Ueno et al. [153] conducted a study revealing that while people can relatively easily identify the first and last panels in a sequence, accuracy diminishes with the presentation of all panels from four-panel strips. This suggests the complexity involved in discerning narrative continuity in comics.
The same study attempted to automate panel sorting using an AlexNet-like architecture. However, the CNN’s performance was poor, indicating the difficulty of the task and the need for more advanced or specialized algorithms in panel sorting.
7.2.3 Dialog transcription
Definition: Dialog transcription in comics aims to create an automatic transcript of dialogues, identifying the speaker and the sequence of speech.
The process typically involves several steps: (i) detecting panels, text boxes, and character boxes, (ii) associating text with speakers, and (iii) sorting text boxes according to their reading order to generate a coherent dialogue transcript.
This task has been addressed by [73] for Japanese Mangas, proposing the needed annotations for Manga109, and more recently by [41] for English Mangas called PopManga. However, a few limitations arise from these approaches: sorting the text boxes is always done as a postprocessing operation, depending only on algorithmic approaches. Cases in which balloons alternate among close panels are inherently mistaken. Recently, Sachdeva et al. [43] proposed Magiv2 which tackled the problem of dialog transcription adding also names (from a name bank) allowing it to operate on the whole collection of pages.
In CoMix [42], a first benchmark for this task was proposed, authors enhanced the previous annotations and proposed a new metric for character naming (based on ANLS) and dialog transcription (based on Hungarian matching and edit distance).
7.3 Segmentation
Segmentation in comic analysis, particularly instance segmentation, presents a nuanced approach to identifying and isolating various elements within a comic panel.
7.3.1 Instance segmentation
Definition: Instance segmentation in comics extends beyond the scope of typical object detection. Rather than just identifying and classifying elements within bounding boxes, this task involves generating precise pixel-wise masks for each instance.
By focusing on segmentation instead of detection, we can achieve more accurate panel separation and representation. This precision is especially beneficial when reconfiguring or reformatting comics for different layouts. It ensures that each panel is cleanly isolated, without the remnants of adjacent panels or the original layout, which is a common issue with less precise bounding boxes.
One of the key challenges in instance segmentation within comics is dealing with the diversity of artistic styles and the complexity of layouts. Comics often feature intricate designs, overlapping elements, and varying levels of detail, which can complicate the segmentation process. As described in [47], there are dozens of configurations in which panels appear to respect each other (pure grid, vertical and horizontal staggering, bleed, inset, overlap, etc.). Addressing the task of separating panels and understanding their relative positions within Cohn nomenclature is hard with bounding boxes when delimitations are not precise.
Future research in comic instance segmentation could explore advanced deep-learning models that are adept at handling these complexities. The development of specialized algorithms that can adapt to the unique characteristics of different comic styles would significantly advance the field. There’s also potential for integrating instance segmentation with other tasks like character-text association and panel sorting to create a more comprehensive comic analysis system.
Across the presented datasets, no one proposes semantic segmentations.
7.4 Future Tasks
The Layer of Comics Understanding is a general framework with which we aim to collect various Vision-Language tasks. Therefore, it includes also tasks that are currently not explored in comics, such as Translation.
7.4.1 Translation
Definition: The task of translation in comics refers to translating the textual elements that appear in a panel or a page, according to the scene. It goes beyond simple text translation, intertwining linguistic, graphical, and spatial challenges. It necessitates an integrated approach that respects the visual and narrative structure of the comic.
While translation in comics may seem like a straightforward NLP problem, it involves unique complications. Text in comics appears in various forms such as balloons, onomatopoeias, and scene text, each requiring a different approach. In Balloon Text, translated text must fit within the space of the original speech balloon, preserving the visual integrity of the panel. Complications arise when the text to be translated contains elements like sarcasm, irony, jokes, or idiomatic expressions. In Onomatopoeias, both linguistic translation and graphic replacement are often required to maintain the style consistent with the target language. Lastly, in Scene Text, translating involves both linguistic translation and potentially replacing the text within the scene, which can be a complex graphic design task.
8 Layer 3: Retrieval and Modification
Layer 3 in the Layers of Comics Understanding focuses into advanced image and text understanding through tasks such as Retrieval (uni-modal and multi-modal) and Modification.
8.1 Retrieval
8.1.1 Unimodal Retrieval
Definition: Unimodal retrieval in comics involves searching for a specific media element (image or text) within a database, using a query of the same format (image or text, respectively).
Pioneering works in this area have explored various aspects of comic retrieval. Wu et al. [154] developed a comic retrieval system that leveraged OCR and external text.
One significant article [138] focused on retrieving full pages or books containing similar elements. This approach is particularly relevant in copyright violation contexts. The authors implemented feature vectors for each image, storing them in a database, and used these vectors to find the most similar documents. Their feature representation method was based on the Histogram of Oriented Gradients (HOG), which outperformed the traditional SIFT method. Notably, they applied HOG not to the entire images but selectively to Regions Of Interest (ROIs) identified in the comics, such as faces or specific markings.
Building on this approach, authors in [137] utilized a similar methodology for manga character retrieval. They focused on identifying and extracting distinctive features from characters, enhancing the ability to retrieve specific characters from large databases.
8.1.2 Cross-modal retrieval
Definition: Cross-modal retrieval in comics encompasses tasks where the query and the retrieval results are from different modalities, such as using sketches or text queries for retrieving a comic image.
Sketch-based Retrieval. Matsui et al. [32] introduced a framework for retrieving manga images from sketches using the Fine Multi-scale Edge Orientation Histogram (FMEOH). This method indexed different-sized squares on a page and enabled efficient retrieval from sketch queries. Further advancing this area, authors [155] proposed a two-step system that first processes manga and then retrieves images based on sketches. This framework was one of the first to employ the Manga109 dataset and included margin labeling, objectness-based edge orientation histograms, and approximate nearest neighbor search. Narita et al. [156] explored CNN-based feature extraction for sketches and manga images (without and with screentone, respectively), demonstrating improvements over algorithmic approaches.
Multimodal Learning and Character Retrieval. Nguyen et al. [157] proposed a multitask multimodal approach for character image and text learning, aiming at tasks such as character retrieval and emotion recognition. Their method focused on leveraging both visual and verbal information in manga content, in a CLIP-like style aiming at performing various character tasks (e.g. detection, retrieval, emotion classification), trained uniquely with visual information and verbal information in manga image content. Later, in [158] authors introduced ComicLib, a dataset for comic sketch research, providing benchmarks across various tasks like colorization, generation, and retrieval. They conducted extensive comparison experiments with other datasets (e.g. QuickDraw [159]) to provide a benchmark for ComicLib on common tasks like colorization (Style2Paints [160]), generation (DCGAN [161]), retrieval (image-to-image with ResNet), detection (with Faster R-CNN [162], YOLOv3 [163], and SSD [164]) and recognition. However, the dataset’s availability is unclear.
Text-to-Image Retrieval. Shen et al. [165] presented a system for text-to-image content retrieval within Manga frames. This multi-staged system integrates object detection, text recognition, and a vision-text encoder to facilitate efficient search and retrieval of dialogues and scenes. Their method is composed of an object detection model for identifying text and frame bounding boxes (DETR-based [166]), a Vision Encoder-Decoder model for text recognition (DiT [167]), a text encoder for embedding text (multilingual DistilBERT [168]), and a vision-text encoder with unified embedding space (CLIP [14]). They experimented with Japanese and GPT-4101010Using the website https://platform.openai.com/translate/rephrased English sentences, on the annotated panels from “Dollgun Book”. Despite being a composition of existing systems on panel level, on a small distribution of data, the work is a first direction to more complex Comic retrieval.
8.2 Modification
In the realm of comics, modification tasks such as image inpainting and image editing play a crucial role in streamlining the storytelling process and enhancing creative expression.
8.2.1 Image Inpainting and Editing
Definition: Image inpainting and editing in the context of comics involve various techniques and tools that allow for the alteration or enhancement of comic panels, facilitating more dynamic and engaging storytelling.
One innovative tool in this space is CodeToon, presented by [169]. CodeToon is a comic authoring tool that facilitates a code-driven storytelling process. It leverages two mechanisms: story ideation from code through metaphorical interpretation and automatic comic generation from these ideated stories. Significant attention has been given to the interface design to address common challenges faced by comic creators, such as generating stories directly from code snippets and seamlessly transitioning these stories into comic format.
Another notable contribution is ComicScript by [170], also known as interactive data comics [^11] Accessible at https://interactivedatacomics.github.io/.. ComicScript is a suite of operations designed in collaboration with practitioners to create and modify comics. These operations allow for the addition and removal of panels, support branching narratives, perspective changes, and detailed exploration. They also facilitate data manipulation and interaction within the comics, enhancing personalization and reader engagement. To validate the utility of ComicScript, professional illustrators, designers, and data comics enthusiasts were recruited to craft interactive comics. Their experiences provided valuable insights into the authoring workflow and the potential of interactive comics as a storytelling medium.
8.3 Future Tasks
8.3.1 Personalized Image retrieval
Definition: Personalized Image Retrieval (PIR) focuses on developing person-centric models that can efficiently retrieve specific comic panels based on a compound query involving both image and text inputs [171]. PIR aims to identify and retrieve images or videos that correspond to a compound query, which typically includes an image of a person’s face combined with a textual scene or action description.
In the realm of comics, this translates to the capability of retrieving panels where a specific character appears, given an image of the character, the character’s name, or a description of the character. This task becomes increasingly complex when the retrieval criteria include context-specific actions (e.g., a character eating pizza or jumping out of a window) or interactions with other characters.
9 Layer 4: Advanced Visual-Language Understanding
9.1 Understanding
This layer investigates tasks that require a deeper integration of visual elements and language, advancing beyond basic identification or retrieval to more complex forms of comprehension and reasoning within comics.
9.1.1 Visual Question Answering
Definition: Given an image-question pair, Visual Question Answer requires answering a question based on the image. Most studies treat VQA as a classification problem on a predefined answer set, but its general definition would be an open answer where no specific set of options is provided.
The first notable attempt in the domain of the comic is by [172] with ComicVQA. However, their approach, more akin to retrieval based on specific “Doraemon” comics datasets, diverges from the typical VQA framework. Later, Sumi et al. [173] presented a system, ComicsQA, with a unique twist, converting user QA interactions into comic stories, reflecting the user’s situation and offering solutions.
The creation of a comprehensive VQA dataset specifically for comics, encompassing single-panel, single-page, and multi-page formats, remains an open challenge. In a recent survey [174], Zeng et al. show the massive work that has been carried out on VQA on multi-modal data across different mediums (movies, comics, etc.) and Multimodal Machine Comprehension [175, 176]. This, together with multiple datasets across various mediums [177, 178, 179, 180, 181, 181] are laying the groundwork for advancing VQA also in comics.
9.2 Future Tasks
In the context of Understanding, a number of Vision Language tasks could be of interest in comics, both for pertaining and fine-tuning.
9.2.1 Visual Entailment
Definition: Visual entailment (VE) in comics is an area yet to be deeply explored. This task involves assessing whether a given image semantically entails the accompanying text. The challenge lies in interpreting the visual narrative in conjunction with textual elements to ascertain semantic consistency or correlation. It can be seen as a simplified version of the text-closure task where we aim at correctly classifying whether the image is represented by the given text, however, depending on the provided text, VE could quickly lead to complex scenarios.
9.2.2 Visual Dialog
Definition: The task of Visual Dialogs (VisDial) corresponds to answering a specific question, having an image, a question about the image, and the previous dialog history about the image.
In comics, Visual dialog is an area still in its infancy and requires constructing a narrative dialogue based on visual cues within the comic panels. This task demands a comprehensive understanding of the storyline, character interactions, and visual symbolism to generate contextually relevant dialogues.
9.2.3 Visual Reasoning
Definition: Visual reasoning in comics extends beyond simple question answering to include the ability to interpret and analyze the visual content in depth. This task necessitates understanding the objects, their interactions, and the underlying narrative structure within the comic panels. It’s a sophisticated blend of visual comprehension and logical deduction, aimed at uncovering deeper layers of meaning in the comics.
10 Layer 5: Generation and Synthesis
The last layer explores the creative frontier in comics analysis, where the synthesis and generation of comics from various media sources play a pivotal role.
10.1 Generation
The field of Generation in comics has seen lots of attraction on Manga style comics rather than American ones. Their popularity has attracted the research community towards creating comics from various media such as video, images, and text descriptions.
10.1.1 Comics generation from other media.
Definition: The task of generating comics from other media comprehends an umbrella of tasks with various input types (video, audio, charts, and comics itself) whose output is a comic panel, a sequence of panels, a page, or a book.
One of the earliest attempts at comic generation from other media is the “Comic Live Chat” by Matsuda et al. [182], which transformed video meetings into comics. This approach involved selecting keyframes and rendering the dialogues of meeting participants in a comic format. Tanapichet et al. [183] proposed a system for creating comic strips from cartoon animations using optical flow techniques, demonstrating a novel approach to narrative translation from one visual medium to another. Hoashi et al. [184] worked on creating manga previews by detecting panels and text balloons, while in [185, 186] authors explored generating manga faces and caricatures from real images. Cao et al. [187] and Herranz et al. [188] introduced methods for comic-like video summarization. These systems developed pipelines for layout proposal, image reconfiguration, and balloon creation, treating comic creation as a series of steps from keyframe selection to final panel layout. On the same task, to automatically create comics from video content, Jing et al. [189] pioneered video transformation using TV show subtitles and a speaker detection algorithm to generate comic layouts optimized through an energy-based metric. This approach was a significant step forward compared to earlier methods like [190].
Giving it more time, a groundbreaking approach by Pesko et al. [191] employed neural style transfer, creating visually compelling comic representations from video-sourced data [^12] GitHub repository: https://github.com/maciej3031/comixify. While their startup has since become closed-sourced, their initial approach signifies a significant leap in comic generation technology. Yang et al. [192] presented a system to generate comics from movies, employing a multi-page layout framework and emotion-aware balloon generation, showcasing the potential of comprehensive systems in comic generation. To extract keyframes from movies, they used GIST [193] similarity between two frames. To detect the speaker, they perform lip motion analysis and use that to associate balloons with talking characters. Finally, the movie frames are rendered in a comic-layout page after the adaption to comic style. Moreover, in [194] authors represent conversational videos to comics “fixed-layout” representation, developing a system that learns guidance field which provides a prior prediction of the possible positions of word balloons while making the word balloons not overlap with other nonverbal information (e.g. hand gestures, visual clues in the background, etc). Their input is raw video, thus the text dialog is captured from audio and rendered in the location chosen from their method.
Comics have been used also in education [195, 196, 197] to teach and explain in an easier way different concepts. An example of this is [198] where authors proposed a system that crafts data stories from a collection of user-created charts, using a comic-style panel to imply the underlying sequence and logic of data-driven narratives.
10.1.2 Image-to-Text generation
Definition: Image-to-Text is a general umbrella term that covers tasks from captioning, to textbox prediction in comics, masking a specific part of the dialog.
Motivated by supporting the Blind or Low Vision community (BLV) - also known as People with Visual Impairements (PVI) - Ramaprasad et al. [199] developed a two-step method to create natural language descriptions of comic strips. This method first extracts information about panels, characters, and text using computer vision techniques, followed by the use of a multimodal large language model (MLLM) to generate descriptions. This work stands as one of the few promoting MLLM specifically for comic strip captioning, despite using “out of the box” architectures like Grounding DINO [40] to extract arbitrary elements, CLIP [14] to match character images with name-descriptions, and LLaVa [200] as MLLM.
Vivoli et al. [72] proposed an advanced version of the text-cloze task with image-to-text generation, increasing the challenge level. They adopted the VL-T5 model [16] using two configurations with similar performances but different model sizes: with BLIP-2 visual backbone (1.3B) or a ResNet-50 fine-tuned with SimCLR on comics (0.2B). They surpassed previous state of the art in the standard “text-cloze” task, and propose benchmarks in the generation version of the task. Authors in [77] presented a similar task but incorporated character descriptions and transcriptions, emphasizing a language model-centric approach. Guo et al. [201] introduced the multi-modal manga complement task. This innovative task combines visual sequences from comic pages with corresponding text dialogues, challenging the model to complete the narrative appropriately. They propose an effective method with CLIP as Feature Encoding, a custom cross-attention module called Fine-grained Visual Prompt Generation, and a transformer-based encoder-decoder architecture called Dialog complement. In a recent work [41], authors proposed a Relationformer-based architecture that generates dialog for manga pages. Their approach includes detection, matching, and sorting algorithms to produce an ordered dialog that aligns with the visual narrative. However, all these works are limited by the dialog appearing in the comic. No work has explored producing a multi-page description of the scene, the actions, and the surrounding environment of every single panel.
10.1.3 Text-to-Image generation
Definition: The Text-to-Image task comprehends generating an image, or a sequence of images, from a text description.
In the field of Text-to-Image generation, one of the earliest endeavors was by Jin et al. [202], who successfully created anime-style character faces using a GAN trained on a curated dataset of anime faces, showcasing the potential of GANs in generating stylized images. Inoue et al. [68] introduced novel methods for domain transfer and pseudo-labeling in image generation, marking significant advancements in the area. StoryGAN [203] emerged as a groundbreaking initiative in story visualization, employing a sequential conditional GAN framework equipped with a context encoder to dynamically track story flow and a story-level discriminator. Enhancements to StoryGAN have included a dual learning framework to bolster semantic consistency between stories and images by refining captions [204] and a recursive architecture to handle structured text inputs via constituency parsing trees [205]. In [206], a deep learning pipeline was proposed for generating synthetic graphic novels using GPT-2 for text and StyleGAN2 for image synthesis, trained on a large collection of manga pages. Proven-Bessel et al. [207] developed a text-to-image GAN for generating single-panel comics in the style of Dilbert, showcasing new possibilities for comic generation from text. Another approach involved FastGAN [208], enhanced with a condition vector for generating high-quality images from small datasets, such as manga faces drawn from Osamu Tezuka’s works [209], which showed improved FID scores compared to the original FastGAN. In [210], a multilingual text-to-image model was developed for creating webtoons, demonstrating the adaptability of GANs to multilingual contexts. However, despite various training experiments, the webtoon fine-tuned GAN underperformed compared to fine-tuning Korean MSCOCO, with the latter still producing abstract images that noticeably differed from authentic comics.
Significant progress has also been observed in story visualization through diffusion models. Innovations like StoryDALL-E [211] employ autoregressive Transformers for extensive model tuning and adaptive adjustments. Everaert et al. [212] discovered the ability to adapt Stable Diffusion to various styles, including comics, by adjusting the initial latent tensor, showing how diffusion models can contribute to comic generation. Despite diffusion models, both on pixel and on latent space, being not good at generating text, the comic panels generated are in line with COMIC data distribution. The topic of story visualization as a subfield of text-to-image generation is very active, moving also towards integration of more powerful LLMs [213], with the main challenges being the consistency of characters and style across multiple images (panels).
Lastly, as we saw previously for cross-modal retrieval tasks, using existing models as building blocks to automatically annotate new data is a good option, especially when the task doesn’t exist or annotations are expensive. In [214], authors used an LLM (ChatGPT) to generate storylines and dialogue of a “possible continuation” of the well-known manga One Piece. Then, they fine-tuned stable diffusion with LoRA and ControlNet to adapt to the manga layout, colorization, and character aesthetic. They finally generated the comic using the fine-tuned stable diffusion, rendered it in panels, and added the LLM dialogs to balloons and narration boxes.
10.1.4 Sound Generation
Definition: Sound generation in the comics domain stands for generating sound effects for a single panel, or in the transaction from one panel to another.
In [215], authors proposed a system that automatically narrates the input comic page of “Clubhouse Rascals” by synthesizing realistic speeches for each character. The synthesized speeches carry the identity properties (gender and age) and emotional conditions (e.g., happy, sad, angry) of the characters inferred from the comic input. They perform a series of subsequential steps including detection of panels, speech balloons, text, character, and character’s face, as well as character re-identification, text-character association, and dialog generation. They classify every character’s face with gender, age, and emotional state, and generate the comic speech synthesis given all this information as input to a waveform reconstruction. They used the Manga109 dataset for all experiments.
In a later study [216], authors explored whether having scene description and sound effects could improve the user experience of comics for People with Visual Impairments (PVI), discovering that both types are welcomed with the first (scene description) helping fr concentration and understanding of the situation, and the second one (sound effect) perceived as more immersive and realistic book-reading experience.
10.2 Synthesis
The group of Synthesis tasks, compared to the “generation” one, refers to the creation of more complex, structured, and often temporally extended outputs, such as creating a full graphic novel from a complex storyline or merging multiple elements (characters, scenes, dialogues) in a consistent way across multiple pages.
10.2.1 3D generation from images
Definition: The task of generating 3D models from 2D comic illustrations presents unique challenges. Unlike conventional portrait illustrations, comics often involve stylized and exaggerated features, adding complexity to this already intricate process.
While there are advancements in using 3D character assets for tasks like pose estimation [217], re-targetting [218, 219], and character reposing [220], the field lacks scalable training resources due to the unavailability of suitable 3D character assets. In light of these issues, Chen et al. [221] addressed this gap by formalizing the stylized reconstruction task with the introduction of the AnimeRecon benchmark and the Vroid dataset. The Vroid dataset, in particular, provides a wealth of 3D assets for scalable training in this domain. A significant part of their work involves solving the challenge of contour removal from illustrations, which is a crucial step in achieving accurate 3D reconstruction from stylized comic images.
While there are numerous works on 3D reconstruction from line drawings in general [222, 223], the specific application in comics is underexplored, likely due to the absence of suitable datasets and ground truth references for comics.
It is worth noting that the exploration of 3D reconstruction from comic images holds great potential. It not only enriches the visual experience of comics but also opens up possibilities for animations, interactive media, and virtual reality applications within the comic domain.
10.2.2 Video generation
Definition: Video generation in comic refers to the synthesis of video content from static comic panels. This process represents a significant leap in merging traditional comic art with dynamic multimedia.
Recent advancements have seen significant progress in generating videos from both images [224, 225] and text [226], although accurately defining metrics and benchmarks remains challenging [79]. These developments are particularly notable in the broader Vision and Language community, showcasing the fast-paced evolution of the field.
The domain of comics, however, is still in the early stages of exploring video generation capabilities. In one of the initial attempts, Cao et al. [227] introduced a framework to animate manga panels. This method involves classifying panels based on motion and emotion and using these classifications to animate backgrounds and characters. The final video creation considers both intra-panel animations and transitions between panels. Later, Gupta et al. [228] further advanced this concept with the Comic-to-Video Network (C2VNet). This framework creates panel-audio videos by cropping comic panels and adding a highlighting effect to speech balloons, simulating characters’ dialogues. Alongside this, they introduced the IMCDB, an annotated dataset of Indian Mythological Comics in English, facilitating research in this area.
While these efforts mark initial steps, they still fall short of the more advanced capabilities seen in general Vision-Language video generation, such as producing fully animated video clips with sound from a single image. The potential for integrating comics with video and audio elements presents an exciting opportunity for enhancing storytelling and audience engagement in the comic medium.
Future research in this area could lead to new forms of interactive and immersive comic experiences, blending traditional art with contemporary digital technologies.
10.2.3 Narrative-based Comic Generation
Definition: The task involves generating a comic (whether a panel, strip, or page) from a plain text description. The term “Narrative-based” highlights that the text not only describes the scene but also conveys the story’s narration.
The pioneering efforts in the realm of narrative-based comic generation commenced with the introduction of StoryGAN by Li et al. [229], which marked a significant advancement in generating image sequences from storylines using a sequential conditional GAN framework. This innovation incorporated a Context Encoder, which dynamically tracked the story flow, laying the groundwork for future developments in this area.
Building on this foundation, Maharana et al. [230] enhanced visual quality and coherence by incorporating a dual learning framework that utilized video captioning, a copy-transform mechanism, and MART-based transformers. This approach significantly improved the semantic alignment between the story and generated images, thus enriching the story visualization process.
In terms of character coherence and continuity, Chen et al. [231] focused on maintaining consistent character portrayal throughout the narrative. They adapted Vector-Quantized Variational Autoencoders (VQ-VAE) with a text-to-visual-token architecture, ensuring character coherence in the visual narratives. Furthering this direction, Maharana et al. [211] introduced the concept of story continuation, where the generated visual story is conditioned on a source image, allowing narratives to incorporate new characters more fluidly. They also introduced the DiDeMoSV dataset, which provided a platform for exploring these complex narrative structures.
Recent advancements in story visualization are exemplified by Pan et al. [232], with presenting AR-LDM, a latent diffusion model that significantly raised the standards for visual quality in natural image datasets like VIST. This model was auto-regressively conditioned on history captions and generated images, capturing complex interactions between frames. Rahman et al. [233] introduced a framework with a visual memory module that implicitly captured actor and background context, generating frames that were not only of high visual quality but also consistent with the story. Peng et al. [234] proposed PCSG, a diffusion-based text-to-image synthesis framework with controllable plugins for character consistency, scene layout specification, and character pose specification, further enhancing the personalized aspects of story visualization.
These emerging trends and future directions in narrative-based comic generation showcase a notable shift towards more sophisticated, context-aware, and visually coherent storytelling techniques. The integration of elements like character consistency, scene dynamics, and actor-background interaction is pivotal in creating compelling visual narratives. Future research in this area can further refine these techniques, potentially leading to automated comic creation tools that adapt to various narrative styles and complexities, opening new horizons in the field of comics and visual storytelling.
10.3 Future Tasks
10.3.1 Comics to Scene graph
Definition: Transforming comics into scene graphs is an emerging area of research. This task involves dissecting comic panels to identify and represent the relationships between different elements within the panel in a graph format.
The last work that tackles this task is [41], which employs a Relationformer architecture [145], explicitly designed for image-to-graph tasks. However, the authors convert the full page into different types of nodes (panels, text, characters) and only consider two types of arches: characters-characters (re-identification) and text-characters (speaker identification). The inherently complexity of the comics, such as the elements in the image (e.g. the background and foreground objects) have not been considered yet. The task has yet to be explored to fine-grained graph creation from single panels, full pages, or a book.
11 CONCLUSION
In reviewing the diverse and innovative efforts in the realm of comic research and technology, it becomes evident that this field is in a state of dynamic and exciting evolution. The array of tasks and applications that researchers have embarked upon reflects not only the complexity of comics as a medium but also their immense potential as a bridge between artistic expression and technological advancement.
We have covered numerous works and categorized them into distinct layers based on a proposed taxonomy. We have aimed to distill the key topics, methods, challenges, and emerging directions that can shape future research in Comics Understanding.
In the future, we will continue to monitor advancements in this field, and we will update our findings in the following repository: https://github.com/emanuelevivoli/awesome-comics-understanding.
A. Methodology details
In this section, we will explore the process of finding the state-of-the-art in Comics and Manga. We utilized the software “Publish or Perish” [^13] Available at: https://harzing.com/resources/publish-or-perish, which comprehends the following Search sources: “Semantic Scholar”, “Google Scholar”, “Web of Science” and “Scopus”. As every source possesses its own interface and query structure, we designed a “SQL-like” query:
  • Title: “comic*”, “manga”
  • Keywords: (“comic*” OR “manga”) AND (“machine*” OR “learning*” OR “ml” OR “computer*” OR “science*” OR “artificial*”)
For every source, we limited the results to the last 1000 from 2010 to 2024. For every paper (out of the 3500 obtained), we used the “Semantic Scholar API” [^14] API at link: https://www.semanticscholar.org/product/api to query the title and obtain the unique DOI. We finally filtered unique papers, automatically removing duplicated copies (DOI) copies, obtaining 2276 papers. This duplicate-removal process has been further implemented also in “zotero” [^15] Zotero software available at: https://forums.zotero.org, fetching data from the websites and using javascript code [^16] Zotero deduplication script. Finally, we define our inclusion/exclusion criteria as follows:
Inclusion. The study regards:
  • (I1)
    Comics and Manga;
  • (I2)
    Computer science;
  • (I3)
    Deep Learning;
  • (I4)
    new Dataset and/or Task.
Exclusion. The study is:
  • (E1)
    not written in English;
  • (E2)
    not available;
  • (E3)
    a duplicate or extension of an already included study;
  • (E4)
    a secondary or tertiary study;
  • (E5)
    in the form of editorials, tutorials, books, extended abstracts, etc.;
  • (E6)
    a non-scientific publication or grey literature;
  • (E7)
    cited at least 2 times (if previous 2022);
  • (E7)
    not published in AI conferences.
After this process, we ended with 216 papers. The last phase of “snowballing” is applied at this point: whi le reading and analyzing the papers we discovered important references that were not in the collection because they were not selected by our query or were discarded by I/E criteria. However, this process is needed to include important papers that are highly cited from our collection. This step enlarged the collection to the final number of 353 papers. Finally, we manually check all the references and include the one that suits the best our taxonomy and scope.
B. Taxonomy
Task classification within the LoCU is informed by both the modalities and dimensions of task inputs and outputs. We define modalities as Images (ranging from no image to multiple pages), Text (from no text to complex multi-sentence structures), Audio (from no audio to complex sound mixes), and Geometry (encompassing points, lines, bounding boxes, segmentations, and 3D geometries). Dimensional analysis considers Space (from non-spatial to tridimensional) and Time (whether time is a factor). Each task is scored based on its input and output characteristics in these modalities and dimensions (see Tables IV and V), with the cumulative score determining its placement within the LoCU layers.
TABLE IV: Modality Scoring for LoCU Tasks
Modality Category Score
Images No image 0
Single panel 1
Single page (multiple panels) 2
Multiple pages 3
Text No text 0
Single word (or tag) 1
Single sentence 2
Multi sentences 3
Instruct multi-sentences 4
Complex multi-sentence 5
Audio No audio 0
Single sound 1
Multi-elements sound 2
Complex mix of sounds 4
Geometry No 0
Point 1
Line 2
Bounding box 3
Instance segmentation 4
Semantic segmentation (multiple masks) 5
3D object geometry 6
TABLE V: Dimensions Scoring for LoCU Tasks
Dimensions Category Score
Space No 0
2D 1
3D 2
Time Not considered 0
Considered 1
In Table VI are reported the values of modality and dimensions, together with the Layer of Comics Understanding Table I. This shows the modalities and dimensions value assigned to every task’s input and output, in order to obtain an objective complexity score that could sort each of the tasks. The acronyms “I, T, A, G” represent “Images, Text, Audio and Geometry”, respectively. Similarly, the acronyms “S, T” correspond to the Dimensions “Space, and Time”, respectively.
TABLE VI: Detailed view of Modalities and Dimensions for both inputs and outputs.
Mods Dim Mods Dim
Layer Category Tasks Input I T A G S T Output I T A G S T Score
0 View Basic Image Viewing (BIV) Text Command 0 0 0 0 0 0 Image Display 0 0 0 0 0 0 0
Image Classification (I-CLS) Image 1 0 0 0 0 0 Tag 0 1 0 0 0 0 2
Emotion Classification Comic Panels/Images 1 0 0 0 0 0 Emotion Labels 0 1 0 0 0 0 2
Action Detection Multiple Panels 2 0 0 0 1 1 Tag 0 1 0 0 0 0 5
Tagging Page Stream Segmentation (PSS) Images 2 0 0 0 1 0 Tags sequence 0 1 0 0 0 0 4
Image Super-Resolution (ISR) Image 1 0 0 0 1 0 Image 1 0 0 0 1 0 4
Style Transfer (ST) Image 1 0 0 0 1 0 Image 1 0 0 0 1 0 4
Vectorization Comic Panels/Images 1 0 0 0 1 0 Vector image 1 0 0 1 1 0 5
1 Augmentation Depth Estimation Comic Panels/Images 1 0 0 0 1 0 Depth Map 1 0 0 0 1 0 4
Object Detection Tag/s + Image 1 1 0 0 1 0 Bounding Boxes 0 0 0 2 1 0 6
Character Re-identification (Character ID) Multiple Panels 2 1 0 0 1 0 Text [Tagging] 0 1 0 2 0 0 7
Grounding Grounding (IG) [prompt] + Image 1 2 0 0 1 0 Bounding Boxes 0 0 0 2 1 0 7
Character-Balloon Association (Speaker ID) Character + Balloons 2 0 0 2 1 0 Text [Tagging] 0 0 0 2 0 0 7
Dialog transcription Image 1 0 0 0 1 1 Text 0 3 0 0 0 1 7
Analysis Translation Image 1 3 0 0 1 0 Text 0 3 0 0 0 0 8
2 Segmentation Instance Segmentation (IS) [prompt] + Image 1 0 0 2 1 0 Segments 0 0 0 3 1 0 8
Image-Text Retrieval (IR) Text 1 3 0 0 1 0 Image 0 2 0 0 1 0 8
Text-Image Retrieval (TR) Image 3 1 0 0 1 0 Text 2 0 0 0 1 0 8
Retrieval Composed Image Retrieval (CIR) Text + Image 3 2 0 0 1 0 Image 2 0 0 0 0 0 8
Image Inpainting (II) Text + [prompt] + Image 1 2 0 2 1 0 Image 2 0 0 0 1 0 9
3 Modification Image Editing via Text (IET) Text + Image 1 4 0 0 1 0 Image 2 0 0 0 1 0 9
Visual Dialog (VisDial) Image + Dialog + Text 2 4 0 0 1 1 Text 0 3 0 0 1 1 13
Visual Reasoning (VR) Image + Text 2 3 0 0 1 1 Text 0 3 0 0 1 1 12
4 Understanding Visual Question Answering (VQA) Image + Text 2 3 0 0 1 1 Text 0 2 0 0 1 0 10
Image Captioning [img-2-text] (IC) Image 1 0 0 0 2 0 Text 0 3 0 0 2 0 8
Image Generation [text-2-img] (IG) Text 0 5 0 0 1 0 Image 1 0 0 0 2 0 9
Scene Graph Generation for Captioning Comic Panel 1 0 0 0 2 0 Scene Graph 0 3 0 3 2 0 11
Generation Sound Generation from Single Panel Single Comic Panel 1 0 0 0 2 0 Sound/Audio 0 0 4 0 3 1 11
3D Model Generation from Images (3DGI) Collection of Images 3 0 0 0 1 0 3D Model 0 0 0 3 3 0 10
Video Generation from Text (VGT) Complex Long Text 0 5 0 0 1 1 Video 2 1 0 0 1 1 12
5 Synthesis Narrative-Based Complex Scene Generation (NCSG) Detailed Narrative Text 2 5 0 0 1 1 Series of Images 3 1 0 0 1 1 15
[Uncaptioned image] Emanuele Vivoli is pursuing its joint-PhD at the Computer Vision Center (Autonomous University of Barcelona) and at the MICC Lab (University of Florence). His research interests encompass multi-modal learning (vision and language integration), with a focus on comics. His work covers a diverse range of topics, including document and table analysis, understanding comics and manga, and creating datasets and frameworks. If you are interested, read his papers.
[Uncaptioned image] Andrey Barsky is a postdoctoral researcher at the Computer Vision Center, Barcelona, Spain. He received his Ph.D. in 2015 from the University of Nottingham in the UK. His research focuses on computer vision and multimodal learning, as well as robustness and explainability in AI models.
[Uncaptioned image] Mohamed Ali Souibgui is a postdoctoral researcher at Computer Vision Center, Barcelona, Spain. He received the Ph.D. degree in 2022 from the Universitat Autònoma de Barcelona (UAB), Spain. His current research focuses on document image analysis using computer vision and machine learning tools, with the focus on privacy and explainability. He is an author of several journal and peer-reviewed conference papers.
[Uncaptioned image] Artemis Llabrés is a PhD student in the Computer Vision Center. She is part of the Vision, Language, and reading group where she is working on multimodal models for document understanding. In 2024 she was awarded best student paper by the 16th IAPR International Workshop on Document Analysis Systems, for her work titled “Image-text matching for large-scale book collections”.
[Uncaptioned image] Marco Bertini is an Associate Professor in Computer Science at the University of Florence, Italy. He is the director of the Media Integration and Communication Center of the University of Florence. His interests are focused on multimedia and computer vision. On these subjects he has addressed semantic analysis, automatic content indexing, semantic retrieval and video quality improvement, applying these techniques to different domains among which cultural heritage. He is author of more than 30 journal papers and more than 140 peer-reviewed conference papers. He has been involved in 10 EU research projects as WP coordinator and researcher, among which IM3I, euTV, ORUSSI, UMETECH, AI4Media and ReInHerit. He has been general and program co-chair of several conferences on multimedia. He is co-founder of Small Pixels, an academic spin-off working on GenAI solutions to improve video quality and video compression.
[Uncaptioned image] Dimosthenis Karatzas is a professor at the Universitat Autónoma de Barcelona and associate director of the Computer Vision Centre, leading the Vision and Language research group. His research focuses on computer vision and machine learning, with an emphasis on robust reading systems and document image analysis. He received the 2013 IAPR/ICDAR Young Investigator Award. He has created two spin-off companies. He chaired the IAPR TC11 on Reading Systems from 2016 to 2021, he is an IEEE senior member, an ELLIS fellow and co-director of the ELLIS Unit Barcelona, and a member of the Artificial Intelligence Doctoral Academy (AIDA) Research and Industry Board. He launched the Robust Reading Competition portal, which serves more than 50,000 researchers worldwide.

### v2

@@
 One missing piece in Vision and Language: A Survey on Comics Understanding
-Abstract
-Vision-language models have recently evolved into versatile systems capable of high performance across a range of tasks, such as document understanding, visual question answering, and grounding, often in zero-shot settings.
-Comics Understanding, a complex and multifaceted field, stands to greatly benefit from these advances.
-Comics, as a medium, combine rich visual and textual narratives, challenging AI models with tasks that span image classification, object detection, instance segmentation, and deeper narrative comprehension through sequential panels.
-However, the unique structure of comics—characterized by creative variations in style, reading order, and non-linear storytelling—presents a set of challenges distinct from those in other visual-language domains.
-In this survey, we present a comprehensive review of Comics Understanding  from both dataset and task perspectives.
-Our contributions are fivefold: (1) We analyze the structure of the comics medium, detailing its distinctive compositional elements; (2) We survey the widely used datasets and tasks in comics research, emphasizing their role in advancing the field; (3) We introduce the Layer of Comics Understanding  (LoCU) framework, a novel taxonomy that redefines vision-language tasks within comics and lays the foundation for future work; (4) We provide a detailed review and categorization of existing methods following the LoCU framework; (5) Finally, we highlight current research challenges and propose directions for future exploration, particularly in the context of vision-language models applied to comics.
-This survey is the first to propose a task-oriented framework for comics intelligence and aims to guide future research by addressing critical gaps in data availability and task definition.
+Abstract.
+Vision-language models have recently evolved into versatile systems capable of high performance across various tasks, often in zero-shot settings.
+Comics Understanding, a complex and multifaceted field, stands to benefit from these advances greatly.
+As a medium, comics combine rich visual and textual narratives, challenging AI models with tasks that span image classification, object detection, instance segmentation, and deeper narrative comprehension through sequential panels.
+However, the unique structure of comics—marked by stylistic variation, reading order complexity, and non-linear storytelling—poses distinct challenges.
+In this survey, we provide a comprehensive review of Comics Understanding from both dataset and task perspectives.
+Our contributions include: (1) analyzing the unique structure and elements of the comics medium; (2) surveying key datasets and tasks in comics research; (3) introducing the Layer of Comics Understanding (LoCU) framework, a novel taxonomy for redefining vision-language tasks in comics; (4) categorizing existing methods using the LoCU framework; and (5) identifying research challenges and future directions for applying vision-language models to comics.
+This survey pioneers a task-oriented framework for comics intelligence, aiming to guide future research by addressing gaps in data and task definition.
 A project associated with this survey is available at https://github.com/emanuelevivoli/awesome-comics-understanding.
-Index Terms:
-Survey, Comics, Manga, Layers of Understanding, Comics Understanding, Vision-Language tasks
-1 Introduction
-Comics serve as a sophisticated medium that combines visual and textual elements to tell stories [1].
+Comics understanding, Vision and Language tasks, Comics, Manga, Sequential Visual Art
+$c$s: Computing methodologies Information extractionccs: Computing methodologies Computer vision representationsccs: Computing methodologies Image representationsccs: Computing methodologies Computer graphicsccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processing
+1. Introduction
+Comics serve as a sophisticated medium that combines visual and textual elements to tell stories (Scholarsarchive et al., 2018).
 Applying recent machine learning tools to understand comics is complicated.
@@
 Motivated by these challenges, advanced artificial intelligence (AI) approaches have been increasingly applied to Comics Understanding.
-Given that AI thrives on tackling complex and diverse tasks, many researchers are now focusing on problems like object detection [2], semantic segmentation [3], optical character recognition (OCR) [4], recurrence of characters and objects in varying contexts [5], pose estimation [6], depth estimation [7], and integration of visual and textual narratives [8].
+Given that AI thrives on tackling complex and diverse tasks, many researchers are now focusing on problems like object detection (Vivoli et al., 2024c), semantic segmentation (Bhattacharjee and Aydemir, 2024), optical character recognition (OCR) (Baek et al., 2022), recurrence of characters and objects in varying contexts (Li et al., 2021), pose estimation (Augereau et al., 2018), depth estimation (Bhattacharjee et al., 2021), and integration of visual and textual narratives (Iyyer et al., 2016).
 Comics also include complex narrative elements such as satire and irony and spatial and temporal dynamics within the story that have not been deeply investigated.
-Many of these tasks, individually or in combination, represent the cutting edge of advanced AI, particularly in the area of Multimodal Vision-Language understanding [9, 10, 11, 12, 13].
-Figure 1: Number of publications on Comics and Manga (from Google Scholar). The publications have been filtered by keywords (manga, comics, graphic novels), topics and journals (computer science).
+Many of these tasks, individually or in combination, represent the cutting edge of advanced AI, particularly in the area of Multimodal Vision-Language understanding (Liu et al., 2023a; Li et al., 2024b; Xue et al., 2024; Laurençon et al., 2024; Yao et al., 2024).
+Figure 1. Number of publications on Comics and Manga (from Google Scholar). The publications have been filtered by keywords (manga, comics, graphic novels), topics, and journals (computer science).
 Historically, in machine learning (and deep learning), innovation has been driven by the available data, both in terms of tasks and more capable models.
@@
 The domain of multimodal learning, especially the convergence of vision and language, has witnessed significant advancements.
-These advancements include the shift away from independent uni-modal representations [14, 15] toward the alignment and integration of different modalities [16, 17, 18, 19, 20, 21] and the creation of multimodal embedding space combining aligned uni-modal inputs [22, 23], with early and late fusion.
-Trained on extensive multimodal datasets, these models have shown remarkable abilities in tasks like multimodal retrieval [14, 22, 23], visual question answering, image captioning [17], reasoning [20, 21], and retrieval-augmented multimodal generation [24] as well as image generation [25, 26].
+These advancements include the shift away from independent uni-modal representations (Radford et al., 2021; Jia et al., 2021) toward the alignment and integration of different modalities (Cho et al., 2021; Li et al., 2023b; Singh et al., 2021; Zhu et al., 2023a; Chen et al., 2023b; Wang et al., 2023b) and the creation of multimodal embedding space combining aligned uni-modal inputs (Girdhar et al., 2023; Zhu et al., 2023b), with early and late fusion.
+Trained on extensive multimodal datasets, these models have shown remarkable abilities in tasks like multimodal retrieval (Radford et al., 2021; Girdhar et al., 2023; Zhu et al., 2023b), visual question answering, image captioning (Li et al., 2023b), reasoning (Chen et al., 2023b; Wang et al., 2023b), and retrieval-augmented multimodal generation (Yasunaga et al., 2023) as well as image generation (Gal et al., 2022; Yang et al., 2022).
 The comics domain is uniquely well-suited to driving advancements in these types of multimodal reasoning models.
-Research in comics has extensively explored a range of questions, from the human ability to derive meaning from sequential images [27] to machine interpretation of comic strips, particularly through closure tasks [8].
+Research in comics has extensively explored a range of questions, from the human ability to derive meaning from sequential images (Cohn, 2013b) to machine interpretation of comic strips, particularly through closure tasks (Iyyer et al., 2016).
 With the continuous advancement in AI models and the intensive interest in harvesting vast knowledge from Comics, a lot of work is being published constantly.
@@
 1.
-These contributions shifted the landscape around Comics Understanding, leaving previous surveys outdated [6].
+These contributions shifted the landscape around Comics Understanding, leaving previous surveys outdated (Augereau et al., 2018).
 Nowadays, the Comics Understanding  research community lacks a comprehensive survey that catalogs these papers and the various challenges in the Vision-Language domain.
 To address this need, we introduce in this paper the novel Layer of Comics Understanding (LoCU) framework.
-This taxonomy, illustrated in Table I, delineates tasks by their input/output modalities and the spatio-temporal dimensions necessary for reasoning over comic data, inviting a structured and progressive approach to comics analysis.
+This taxonomy, illustrated in Table 1, delineates tasks by their input/output modalities and the spatio-temporal dimensions necessary for reasoning over comic data, inviting a structured and progressive approach to comics analysis.
 Motivated by the uniquely broad task distribution of the comics medium, this survey aggregates and examines all extant research on machine learning in comics, aimed at highlighting critical unresolved issues in the field.
 It is the first to review the breadth of comics research in a hierarchical structure, and to our knowledge, the first to propose a big-picture taxonomy of comics intelligence from a task-oriented perspective.
-Figure 2: Visualization of the Layers of Comics Understanding tasks. From panel level to multipage, from unimodal to multimodal, from simplest to more complex.
+Figure 2. Visualization of the tasks in Layers of Comics Understanding. From panel level to multipage, from unimodal to multimodal, and from simplest to more complex.
 Within this framework, the survey enumerates the state of the art in Comics Understanding tasks, including image classification, object detection, semantic segmentation, style transfer, multimodal retrieval, synthesis, and generation.
@@
 Section 4 introduces existing tasks and the commonly used datasets in Comics Understanding, also focusing on evaluation procedures.
-In Section 5 we propose the Layer of Comics Understanding taxonomy (LoCU), and for each layer a collection of tasks grouped by difficulty (see Annex LABEL:annex:values).
+In Section 5 we propose the Layer of Comics Understanding taxonomy (LoCU), and for each layer a collection of tasks grouped by difficulty.
 In Section 6 we start with Layer 1: Tagging and Augmentation, Section 7 with Layer 2: Grounding, Analysis, and Segmentation, in Section 8 with Layer 3: Retrieval and Modification, Section 9 with Layer 4: Understanding, and in Section 10 with Layer 5: Generation and Synthesis.
 Finally, we conclude with Section 11.
-2 Background
+2. Background
 This section first presents the development of the comics research field based on the learning paradigm and how it evolves towards more complex tasks.
 Then, we introduce the main limitations encountered by the research community, as well as discuss several related surveys to highlight the scope and contributions of this work.
-2.1 The Comics Research Epochs
+2.1. The Comics Research Epochs
 The development of the comics field can be broadly divided into three phases, including (1) Traditional Machine Learning, (2) Deep Learning, and (3) Modern Foundational Models.
 In what follows, we introduce, compare, and analyze the three paradigms and how they shaped Comics advancements in detail.
-2.1.1 Traditional Machine Learning
-Early machine learning approaches heavily depended on feature engineering, utilizing hand-crafted features [28, 29, 30], which were capable of, e.g., classify the style of a manga using SVMs [31], retrieve a manga character from a sketch with edge orientation histograms [32] or segment comic balloons with energy functions and active contour models [33].
+Traditional Machine Learning.
+Early machine learning approaches heavily depended on feature engineering, utilizing hand-crafted features (Chou and Shih, 2011; Amirshahi et al., 2012; Sun and Kise, 2013), which were capable of, e.g., classify the style of a manga using SVMs (Chu and Chao, 2014), retrieve a manga character from a sketch with edge orientation histograms (Matsui et al., 2014) or segment comic balloons with energy functions and active contour models (Rigaud et al., 2013).
 However, these paradigms require domain experts to craft effective features for specific tasks, which do not cope well with complex tasks and have poor generalizability.
 Moreover, the feature engineering obstacles the extension of the Comics Understanding to more tasks.
-2.1.2 Deep Learning approaches
+Deep Learning approaches.
 With the development of deep learning, researchers have achieved major improvements by using end-to-end trainable deep neural networks (DNNs).
@@
 Applying these paradigms to comics came as a natural step.
-Models became better at detection [34, 35] and segmentation [36, 37, 38], and more complex tasks started being investigated and solved (e.g.
-visual and textual closure [8]).
+Models became better at detection (Qin et al., 2017; Ogawa et al., 2018) and segmentation (Dubray and Laubrock, 2019; Dutta et al., 2021; Nguyen et al., 2020), and more complex tasks started being investigated and solved (e.g.
+visual and textual closure (Iyyer et al., 2016)).
 However, the turn from traditional machine learning toward deep learning raises a new grand challenge: the laborious collection of large-scale, task-specific, and crowd-labeled data in DNN training.
@@
 In the domain of comics, where data is scarce, protected by copyright, and challenging to annotate, these constraints have significantly hindered the development of more generalized, multi-task models for Comics Understanding.
-2.1.3 Modern Foundational models
-The era of foundational models marks a significant evolution in the application of AI to comics, with models that can learn from vast amounts of unstructured data to perform a variety of complex tasks [14, 39, 40].
+Modern Foundational models.
+The era of foundational models marks a significant evolution in the application of AI to comics, with models that can learn from vast amounts of unstructured data to perform a variety of complex tasks (Radford et al., 2021; Oquab et al., 2023; Liu et al., 2023c).
 These models relax previous limitations by “solving” simple yet important tasks as a zero-shot counter effect, e.g.
-panel, character, and text-box detection [41].
-By leveraging these foundational models, more advanced tasks can be proposed and takled [42, 43], opening the doors to yet unexplored more complex ones.
-TABLE I: Overview of Vision-Language Tasks of the Layers of Comics Understanding. The ranking is based on input and output modalities and dimensions, as illustrated in Appendix B. Taxonomy.
+panel, character, and text-box detection (Sachdeva and Zisserman, 2024).
+By leveraging these foundational models, more advanced tasks can be proposed and takled (Vivoli et al., 2024a; Sachdeva et al., 2024), opening the doors to yet unexplored more complex ones.
+Table 1. Overview of Vision-Language Tasks of the Layers of Comics Understanding. The ranking is based on input and output modalities and dimensions, as illustrated in Supplementary Materials.
 Layer Category Task Input Output
@@
 5 (Sec.10) Synthesis (Sec. 10.2) Narrative-Based Complex Scene Generation (NCSG) Detailed Narrative Text Series of Images
-2.2 Relevant Surveys
-To the best of our knowledge, this is the first survey that reviews comics papers with a task-oriented approach for Vision-Language.
-Several relevant surveys have been conducted, examining broader aspects of visual art [44] and, more specifically, comics [45, 6, 46].
+2.2. Relevant Surveys
+To the best of our knowledge, this is the first survey that reviews comics understanding publications with a task-oriented approach for Vision-Language.
+Several relevant surveys have been conducted, examining broader aspects of visual art (Santos et al., 2021) and, more specifically, comics (Augereau et al., 2017, 2018; Nairat, 2021).
 Notably, Augereau et al.
-[6] categorized comic research into three primary areas: (i) content analysis, (ii) content generation and adaptation, and (iii) user interaction, which remain relevant entry points for understanding this field’s research and applications.
-Since those surveys were published, foundational models have emerged, enabling significant advancements, particularly due to their ability to generalize across tasks without task-specific training.
-These models have unlocked new possibilities in comics research, allowing automatic annotation and analysis, among other tasks.
-Building on these advancements, our survey focuses on two key aspects: 1) the recent progress in Comics Understanding tasks and datasets, and 2) the application of Vision-Language tasks to the unique characteristics of comics.
-3 Comics Foundations
+(Augereau et al., 2018) categorized comic research into three primary areas: (i) content analysis, (ii) content generation and adaptation, and (iii) user interaction, which remain relevant entry points for understanding this field’s research and applications.
+In contrast, building on recent advancements, our survey focuses on two key aspects: 1) the recent progress in Comics Understanding tasks and datasets, and 2) the application of Vision-Language tasks to the unique characteristics of comics.
+3. Comics Foundations
 Our preliminary analysis prioritizes understanding the structural intricacies of comics.
 The term “comics” broadly denotes the entire medium in its uncountable form, while “a comic” refers to a singular entity within this medium, such as a comic book or strip.
-The structure of comic books has been a subject of extensive study, examining aspects like their commonalities, layout structures [47], content arrangement, narrative transitions [48, 27, 49, 50, 51], and stylistic narrative variations [52, 53, 54].
-Figure 3: Anatomy of Comic Page Elements: This illustration delineates the various components found on comic pages including object detection, linkages between elements, the designated reading order of texts and panels, as well as key textual features like character names—all arranged in a logical sequence throughout the panels and pages.
+The structure of comic books has been a subject of extensive study, examining aspects like their commonalities, layout structures (Cohn, 2023), content arrangement, narrative transitions (Cohn, 2013a, b; Cohn et al., 2014; Cohn, 2014, 2016), and stylistic narrative variations (Cohn et al., 2017; Ang, 2021; Klomberg et al., 2022).
+Figure 3.
+Anatomy of Comic Page Elements: This illustration delineates the various components found on comic pages including object detection, linkages between elements, the designated reading order of texts and panels, as well as key textual features like character names—all arranged in a logical sequence throughout the panels and pages.
 Typically, comics are presented in single or double-page formats, influenced by the artist’s style.
@@
 Each comic page (as illustrated in Figure 3) encapsulates individual scenes within panels of diverse types - from splash panels for opening scenes to widescreen panels for expansive narratives.
-These panels convey interactions and dialogues through speech balloons, thought bubbles, and expressive onomatopoeia, enriching the narrative experience.
+These panels convey interactions and dialogues through speech or thought balloons, and expressive onomatopoeia, enriching the narrative experience.
 Artistic variations in style and color also play a significant role in comics, ranging from detailed realism to abstract forms, with the color palette setting the mood and atmosphere.
-Despite rigorous analyses [55], no significant differences have been identified in terms of attention, subjectivity, and viewpoint between Manga and American comics.
+Despite rigorous analyses (Cohn, 2011), no significant differences have been identified in terms of attention, subjectivity, and viewpoint between Manga and American comics.
 Hence, these distinctions are better explored within their distinct historical, cultural, and artistic contexts:
@@
 Covering a broad range of genres, Manga is read right-to-left, top-to-bottom.
-This style is marked by a distinctive “Japanese Visual Language” that is consistent across genres [56].
-Comic Formalisms.
-Comics are defined as sequential visual narratives that utilize a combination of images and text within distinct frames (panels) to convey a story or information.
-Each comic, or manga, can be represented as a document $D$.
-Each document $D$ consists of a series of pages $P=\{p_{1},p_{2},\ldots,p_{n}\}$.
-A page $p_{i}$ can be described by its layout and the elements it contains: (i) a spatial arrangement of panels $\{c_{1},c_{2},\ldots,c_{m}\}$, also known as Layout; (ii) each panel $c_{i}$ is a bounded area containing images and possibly text; panels are the primary unit for storytelling in comics; (iii) gutters: the space between panels, often used by artist to leave readers imagining the missing piece of the story, which helps in sequencing and timing; and (iv) text elements, including balloons (speech or thought), captions, and sound effects (onomatopoeias).
-Each panel $c_{i}$ contains characters (depicted through various visual styles and expressions), backgrounds (visual scenery behind the main characters or action), and foregrounds (objects and elements in front of the characters).
-4 Tasks and Datasets
-This section summarizes the commonly used Comics datasets, as detailed in Table II. Some of these datasets are either intended for training or evaluation, depending on their sizes.
-TABLE II: Overview of Comic/Manga Datasets and Tasks, their information (availability, published year, source), and properties (languages, number of comic/manga books and pages).
-The table rows are repeated according to the supported Tasks.
-Accessibility is indicated with for no longer existing datasets, indicates existing but not accessible, and means existing and accessible.
-The link [proj] directs to the project websites, while [data] directs to dataset websites.
-For CoMix, mix means that it inherits from a mixture of four datasets [57, 8, 58, 41].
-For VLRC, Analysis indicates annotations are about statistical analysis of panel size, color, layout, etc.
-However, the dataset does not provide images and annotations have been gathered directly on physical books.
-Task From Name Year Access. Language Origin # books # pages
-Image Classification Sequencity [59] 2017 EN, JP - - 140000
-BAM! [60] 2017 - - - 2500000
-[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
-[8] EmoRecCom [62] [proj] [data] 2021 EN 1938-1954 - -
-Object Detection Fahad18 [63] [proj] 2012 - - - 586
-eBDtheque [57] [proj][data] 2013 EN, FR, JP 1905-2012 25 100
-[57] sun70 [64] 2013 FR - 6 60
-COMICS [8] [proj] [data] 2017 EN 1938-1954 3948 198657
-BAM! [60] 2017 - - - 2500000
-[61] JC2463 [34] 2017 JP - 14 2463
-AEC912 [34] 2017 EN, FR - - 912
-GCN [65] [proj] [data] 2017 EN, JP 1978-2013 253 38000
-[59] Sequencity612 [59] 2017 EN, JP - - 612
-[57] SSGCI [66] [proj] [data] 2016 EN, FR, JP 1905-2012 - 500
-[61] Comics3w [67] [proj] 2017 JP, EN - 103 29845
-comics2k [68] [proj] [data] 2018 - - - -
-DCM772 [58] [proj] [data] 2018 EN 1938-1954 27 772
-[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
-BCBId [69] [proj] [data] 2022 BN - 64 3327
-[61] COO [4] [proj] [data] 2022 JP 1970-2010 109 10602
-[8] COMICS-Text+ [70] [proj] [data] 2022 EN 1938-1954 3948 198657
-PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
-mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
-Re-Identification Fahad18 [63] [proj] 2012 - - - 586
-Ho42 [71] 2013 - - - 42
-[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
-PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
-mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
-Linking eBDtheque [57] [proj][data] 2013 EN, FR, JP 1905-2012 25 100
-[57] sun70 [64] 2013 FR - 6 60
-GCN [65] [proj] [data] 2017 EN, JP 1978-2013 253 38000
-[61] Manga109 [61, 35] [proj][data] 2018 JP 1970-2010 109 21142
-PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
-mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
-Segmentation [59] Sequencity4k [38] 2020 EN, FR, JP - - 4479
-Dialog generation PopManga [41] [proj] [data] 2024 EN 1990-2020 25 1925
-mix CoMix [42] [proj] [data] 2024 EN, FR 1938-2023 100 3800
-Analysis VLRC [47] [proj] [data] 2023 JP, FR, EN, 6+ 1940-present 376 7773
-4.1 Annotations overview
+This style is marked by a distinctive “Japanese Visual Language” that is consistent across genres (Cohn and Ehly, 2016).
+4. Tasks and Datasets
+This section summarizes the commonly used Comics datasets, as detailed in Table 2. Some of these datasets are either intended for training or evaluation, depending on their sizes.
+[…]
+From Name Year Acc. Lang. Origin # books # pages IC OD RI L S DG A
+[…]
+[…]
+(Guérin et al., 2013) sun70 (Sun et al., 2013) 2013 FR - 6 60 x x
+Ho42 (Ho et al., 2013) 2013 - - - 42 x
+[…]
+Sequencity (Nguyen et al., 2017) 2017 GB, JP - - 140000 x
+BAM! (Wilber et al., 2017) 2017 - - - 2500000 x x
+[…]
+(Fujimoto et al., 2016) JC2463 (Qin et al., 2017) 2017 JP - 14 2463 x
+AEC912 (Qin et al., 2017) 2017 GB, FR - - 912 x
+[…]
+(Nguyen et al., 2017) Sequencity612 (Nguyen et al., 2017) 2017 GB, JP - - 612 x
+[…]
+[…]
+[…]
+[…]
+(Nguyen et al., 2017) Sequencity4k (Nguyen et al., 2020) 2020 GB, FR, JP - - 4479 x
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+4.1. Annotations overview
 The available current comics datasets were introduced for different tasks.
-They vary in the year of release, the language and the origin.
-This is shown in Table II.
-From the table, it is noticeable that the available datasets vary significantly in annotation types, making them cover different tasks.
-eBDtheque [57] offers comprehensive annotations in high quality for Character, Balloon, and Panel detection, along with Text-Character association.
+From Table 2 it is noticeable that the available datasets vary significantly in size, languages, origin, and -more importantly- annotations, making them cover different tasks.
+eBDtheque (Guérin et al., 2013) offers comprehensive annotations in high quality for Character, Balloon, and Panel detection, along with Text-Character association.
 This makes it usable for object detection and linking tasks but it comes with a drawback: it is composed of only 100 pages.
-In contrast, the COMICS dataset [8] focuses on a set of tasks called closure, based on automatic panel and text detection over a large number of pages.
+In contrast, the COMICS dataset (Iyyer et al., 2016) focuses on a set of tasks called closure, based on automatic panel and text detection over a large number of pages.
 However, it was introduced with poor annotations.
-Only recently some works [70] have improved OCR text detection on the COMICS dataset.
-In the same vein, [72] shows that improving OCR accuracy (using Textract) could lead to marginal accuracy gain in closure tasks for the COMICS dataset.
-Both [70, 72] released the novel OCR annotations to the research community.
+Only recently some works (Soykan et al., 2022) have improved OCR text detection on the COMICS dataset.
+In the same vein, (Vivoli et al., 2024d) shows that improving OCR accuracy (using Textract) could lead to marginal accuracy gain in closure tasks for the COMICS dataset.
+Both (Soykan et al., 2022; Vivoli et al., 2024d) released the novel OCR annotations to the research community.
 It is to note also that some of the dataset with high-quality annotations and large number of pages are not accessible.
-For instance, BCBId [69], though promising high-quality annotations, falls short in delivery.
-Similarly, VLRC [47], while rich in annotations, lacks digital applicability due to its focus on physical books.
-Manga109 [61] is the first to provide high-level annotations, covering more than 20k pages.
-The dataset has been improved over several iterations, further expanding the scope of the dataset by adding annotations for onomatopoeias [4] and dialog [73].
-The actual state of Manga109 is the result of many years of annotation and improvement by subsequent developments.
-More recently, following the Manga109 wave, PopManga [41] has been proposed as an English-language alternative, including annotations for recognition and linking, but on a smaller scale (2k images).
+For instance, BCBId (Dutta et al., 2022), though promising high-quality annotations, falls short in delivery.
+Similarly, VLRC (Cohn, 2023), while rich in annotations, lacks digital applicability due to its focus on physical books.
+Manga109 (Fujimoto et al., 2016) is the first to provide high-level annotations, covering more than 20k pages.
+The dataset has been improved over several iterations, further expanding the scope of the dataset by adding annotations for onomatopoeias (Baek et al., 2022) and dialog (Li et al., 2023a).
+The current state of Manga109 is the result of many years of annotation and improvement by subsequent developments.
+More recently, following the Manga109 wave, PopManga (Sachdeva and Zisserman, 2024) has been proposed as an English-language alternative, including annotations for recognition and linking, but on a smaller scale (2k images).
 In summary, while datasets for Comics Understanding exist, there is a need for more approaches to standardize and unify their annotations.
-The Comics Dataset Framework [2] proposed a unification of detection annotations as well as a standardization of metrics and settings used to train and measure models.
-The work was subsequently extended in CoMix [42], where mainly English-language comics (and manga) were selected and the annotations extended to the task of single-page dialogue transcription.
-In this work, performances of several state-of-the-art models, including Magi [41], CLIP [14], DINOv2 [39], and GPT4 [74], were measured, for the first time.
-However, as highlighted in the first column of Table II, the complexity of annotations and tasks is limited.
-Object detection (panel, characters, texts), text-to-character linking, and re-identifications are integral tasks to understanding the narrative and dialogue in comics.
-These tasks, along with more recent developments, will be examined in the context of our taxonomy in the following Section.
-4.2 Datasets overview
+The Comics Dataset Framework (Vivoli et al., 2024c) proposed a unification of detection annotations as well as a standardization of metrics and settings used to train and measure models.
+The work was subsequently extended in CoMix (Vivoli et al., 2024a), where mainly English-language comics (and manga) were selected and the annotations extended to the task of single-page dialogue transcription.
+In this work, performances of several state-of-the-art models, including Magi (Sachdeva and Zisserman, 2024), CLIP (Radford et al., 2021), DINOv2 (Oquab et al., 2023), and GPT4 (OpenAI, 2023), were measured, across various datasets, for the first time.
+However, as highlighted in the first column of Table 2, the complexity of annotations and tasks is limited.
+While object detection and linking are integral tasks to understanding the narrative and dialogue in comics, we must advance toward more complex tasks found in Vision-Language research.
+These will be discussed within our taxonomy in the next section.
+4.2. Datasets overview
 The domain of Comics Understanding encompasses a range of datasets focused on processing and examining comic media.
-A review of various datasets, as detailed in Table II, reveals that several are no longer accessible (marked with ) or have restricted access ().
-Those available () often require approval from respective research groups in order to obtain them.
+A review of various datasets, as detailed in Table 2, reveals that several are no longer accessible (marked with ) or have restricted access ().
+Those available ( ✓) often require
+Figure 4. Example of different comic datasets, from Black-and-White to Color, from comics-style to Manga-style.
+approval from respective research groups to obtain them.
 This is a common practice in the field of “sharing copyrighted materials” as in fact they can be used only for research purposes in the majority of cases.
-If the copyright does not allow for redistribution, it is a common practice to release annotations and a script to download and structure the dataset accordingly (as for [41]).
-As we can see from Table II, many datasets share common origins (column “From”).
+If the copyright does not allow for redistribution, it is a common practice to release annotations and a script to download and structure the dataset accordingly (as for (Sachdeva and Zisserman, 2024)).
+As we can see from Table 2, many datasets share common origins (column “From”).
 Notably, these datasets vary significantly in terms of style, quality, and size.
@@
 Datasets are limited in size and the model’s code and weights are rarely shared.
-Luckily, this has started changing recently [2, 41, 42].
-TABLE III: Number of single instance annotation (declared) in papers.
-The “Obj.
-cls” stands for Object-based classification, and refers to the number of classes available if one wants to classify the detected instances in fine-grained classes.
-Mode Name Panel Char. Face Textbox Onomat. Obj. cls
-Color Fahad18 [63] - 18 - - - 18
-Color Ho42 [71] 200 100 - - - -
-Color eBDtheque [57] 850 1550 - 1092 - 5
-Color SSGCI [66] - 50 - - - 1
-Color COMICS [8, 70] 1229664 - - 2498657 - 2
-BW Comics3w [75] 159529 - - - - 1
-BW JC2463 [34] - - 15801 - - 2
-Color AEC912 [34] - - 8184 - - 2
-Color DCM772 [58] 4470 10757 5438 - - 4
-BW Manga109 [61, 35, 4] 103900 157152 118715 147918 61465 4
-Color EmoRecCom [62] 10199 - - - - 8
-BW PopManga [41] - 18783 - 20843 - 2
-mix CoMix [42] 22176 48903 32625 37923 - 4
-In Table III we report, for every dataset, the declared numbers of annotations available.
+Luckily, this has started changing recently (Vivoli et al., 2024c; Sachdeva and Zisserman, 2024; Vivoli et al., 2024a).
+[…]
+Mode / Name Panel Char. Face Text Cls
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+In Table 3 we report, for every dataset, the declared numbers of annotations available.
 On the right-most column, we report the number of classes annotated, which correspond to different class objects for the detection, or the number of classes for classification annotations.
@@
 This is because the characters are separated into three sub-categories: human-like, animal-like, and object-like.
-Here, we describe the datasets chronologically.
-eBDtheque: The eBDtheque [^1] http://ebdtheque.univ-lr.fr/registration [57] is a collection of 100 pages over 20 books, selected from J-comi222previously http://www.j-comi.jp, now no longer available., mainly in the French language (with also English and Japanese comics).
-[76] led a huge effort in labeling 850 panels, 1092 balloons, 1550 characters, and 4691 text lines, provided by domain experts.
-For balloons, they provide instance segmentation and classification ( speech, thought, narration, illustrative).
-Manga109: The Manga109333http://www.manga109.org/index_en.php [61] dataset contains 109 manga volumes from 93 different authors.
-Authors successfully obtained the copyrighted material from manga drawers with a “restrictive term-of-use” which implies redistribution only with the scope of research and the possibility to use them in publication with proper referencing the author.
-These mangas were published between the 1970s and 2010s and are categorized into 12 different genres such as fantasy, humor, sports, etc.
-This dataset has been further extended by COO [4] which added onomatopoeias polygons annotations and links among truncated onomatopoeias.
-In fact, on average there are 5.8 onomatopoeias per page in Manga109, while COMICS has much fewer (0.09).
-Another set of annotations comes from Manga-Dialog [73] which added a text-character link for every speaking text box.
-COMICS: The COMICS [^4] https://obj.umiacs.umd.edu/comics/index.html [8] contains 1,229,664 panels paired with automatic textbox transcriptions using an old-OCR system.
-The 3,948 books are gathered from the Digital Comics Museum (American golden-age comics books) published between 1938 and 1954.
-The dataset includes ground truth (manually annotated) labeled data such as the rectangular bounding boxes of panels on 500 pages and 1,500 textboxes.
-Subsequent works have expanded original annotations with better OCR transcriptions [70, 72].
-In particular, [70] developed a pipeline for OCR processing and labeling of comic books from COMICS, while [72] extracted all OCR transcripts from COMIC using new-generation OCR textract [^5] Used through API at https://docs.aws.amazon.com/textract/.
-Lastly, [77] includes strips of dialog prediction based on “persona” information.
-DCM772: The DCM772666https://git.univ-lr.fr/crigau02/dcm_dataset [58] contains 772 (pages) images from 27 golden-age comic books from the same COMICS source.
+Here, we describe the datasets chronologically and provide an overview of the image style in Figure 4.
+eBDtheque: The eBDtheque [^1] http://ebdtheque.univ-lr.fr/registration (Guérin et al., 2013) includes 100 pages from 20 books, mainly in French, with some English and Japanese comics.
+(Rigaud, 2014) labeled 850 panels, 1,092 balloons, 1,550 characters, and 4,691 text lines, offering expert annotations.
+Balloon labels include instance segmentation and classification (speech, thought, narration, illustrative).
+Manga109: The Manga109222http://www.manga109.org/index_en.php (Fujimoto et al., 2016) dataset includes 109 manga volumes from 93 authors, with copyrighted material obtained under a “restrictive term-of-use”, allowing use for research and publications with proper citation.
+The mangas, published between the 1970s and 2010s, span 12 genres.
+The dataset was extended by COO (Baek et al., 2022), which added onomatopoeia annotations (5.8 per page on average), and Manga-Dialog (Li et al., 2023a), which linked text boxes to characters.
+COMICS: The COMICS [^3] https://obj.umiacs.umd.edu/comics/index.html (Iyyer et al., 2016) dataset contains 1,229,664 panels with automatic textbox transcriptions from golden-age American comics (1938-1954).
+It includes 3,948 books from the Digital Comics Museum and manually annotated bounding boxes for 500 pages and 1,500 textboxes.
+Later works improved OCR transcriptions (Soykan et al., 2022; Vivoli et al., 2024d), with pipelines developed for better OCR processing and transcript extraction.
+(Agrawal et al., 2023) added dialog prediction based on ”persona” information.
+DCM772: The DCM772444https://git.univ-lr.fr/crigau02/dcm_dataset (Nguyen et al., 2018) contains 772 (pages) images from 27 golden-age comic books from the same COMICS source.
 Annotations, however, are more precise as they contain panels, the associated characters bounding boxes, and also face boxes.
-Figure 4: Example of different comic datasets, from Black-and-White to Color, from comics-style to Manga-style.
-4.3 Datasets for Evaluation
-Across the datasets presented in Table II, two recent datasets are provided as evaluation sets.
+4.3. Datasets for Evaluation
+Across the datasets presented in Table 2, two recent datasets are provided as evaluation sets.
 In particular, PopManga proposes two splits (validation and test) while CoMix enhances annotations for various existing datasets among which also PopManga.
-PopManga: The PopManga [^7] https://github.com/ragavsachdeva/Magi [41] dataset contains English manga titles from the most popular mangas.
+PopManga: The PopManga [^5] https://github.com/ragavsachdeva/Magi (Sachdeva and Zisserman, 2024) dataset contains English manga titles from the most popular mangas.
 The dataset contains two test splits: test-seen with 1,136 pages and test-unseen with 789 pages, obtained from 15 and 10 books respectively.
 The test sets have been annotated with Text (both textboxes and onomatopoeias annotated as text), and Characters.
-CoMix: The CoMix [^8] https://github.com/emanuelevivoli/comix-dataset [42] dataset contains English and French comics and manga titles from the most popular datasets: eBDtheque, COMICS, DCM, and PopManga.
+CoMix: The CoMix [^6] https://emanuelevivoli.github.io/CoMix-dataset (Vivoli et al., 2024a) dataset contains English and French comics and manga titles from the most popular datasets: eBDtheque, COMICS, DCM, and PopManga.
 The dataset contains two splits: validation (available with annotations) and the held-out test split (available only through the evaluation server [^7] Accessible at Robust Reading Competition website).
 It contains 3.8k images across 100 books, and annotations for multi-task, from detection and linking to dialog generation.
-5 Taxonomy: LoCU
+5. Taxonomy: LoCU
 Developing a comprehensive taxonomy for comics within the Vision-Language (VL) domain requires a careful overview of existing work in VL surveys and tasks.
@@
 Relevant Taxonomies.
-Initial inspiration stems from the breadth and depth of multimodal questions covered in recent benchmarks like MMMU [78], highlighting the necessity for a taxonomy that accounts for modal complexity and task-specific intricacies in comics.
-This need is further echoed in the domain-specific taxonomy proposed for video understanding in [79], particularly their dimensional analysis approach, which informs our consideration of spatial and temporal domains in comics.
-Shifting focus to VL multimodal evaluation, [80] reframes VL tasks as Visual Question Answering (VQA), emphasizing a range of abilities from grounding to numerical calculation.
-While this framework offers valuable categorizations, it lacks a comprehensive structure for VL tasks, particularly those pertinent to comics.
-Similarly, the multimodal challenges outlined in [81] offer insights into the representation, translation, alignment, fusion, and co-learning challenges in multimodal learning.
-These challenges guide our approach to defining taxonomic classes and subclasses within the VL domain, though with a more task-centric focus.
-Drawing further on [82], we observed the classification of VL tasks into grounding, retrieval, understanding, and generation, with specific adjustments and additions to accommodate comic-specific tasks.
-Our taxonomy expands on this structure, adding categories such as tagging, augmentation, analysis, segmentation, modification, and synthesis, addressing the wider range of tasks observed in comic studies.
-In Table I, we present 31 Vision-Language multimodal tasks, categorized according to the modalities involved, forming the backbone of the LoCU framework.
+Inspiration for this taxonomy comes from multimodal benchmarks like MMMU (Yue et al., 2023), which highlight the need for a structure addressing modal complexity and task-specific intricacies in comics.
+The dimensional analysis in video understanding from (Fan et al., 2024) also informs our spatial and temporal considerations for comics.
+In VL evaluation, (Fu et al., 2023) reframes tasks as Visual Question Answering (VQA), but lacks a comprehensive structure for comics-related tasks.
+Challenges in multimodal learning from (Baltrušaitis et al., 2017) guide our classification of VL tasks, with a task-centric focus.
+(Li et al., 2022) provides a foundational classification of VL tasks into grounding, retrieval, understanding, and generation.
+Their work is crucial in shaping our approach, and we build upon their framework by extending these categories to account for comic-specific tasks.
+In particular, we add categories like tagging, augmentation, analysis, segmentation, modification, and synthesis, which better capture the diversity of tasks observed in comic studies and expand the structure to fully represent the unique challenges in comics understanding.
+In Table 1, we present 31 Vision-Language multimodal tasks, categorized according to the modalities involved, forming the backbone of the LoCU framework.
 This taxonomy, spanning 10 distinct task groups across five layers, aims to capture the multifaceted nature of Comics Understanding.
 Layer 0, representing fundamental image processing and viewing capabilities, sets the stage for increasingly complex interactions in the subsequent layers.
-More details are provided in Appendix B.
-Taxonomy.
+More details are provided in Supplementary Materials.
 This structured approach allows for a comprehensive and nuanced understanding of the range and complexity of tasks within the comics domain, laying a foundation for future exploration and innovation in comic analysis and understanding.
-6 Layer 1: Tagging and Augmentation
+6. Layer 1: Tagging and Augmentation
 The first layer comprises tasks with unimodal input (single image or sequence of images) and unimodal output (Images or Text): Tagging and Augmentation.
-6.1 Tagging
-Tagging tasks predominantly involve classification outputs, albeit with varying input types.
-6.1.1 Image Classification
+6.1. Tagging
+Figure 5. Illustration of Tagging tasks.
+Tagging tasks predominantly involve classification outputs, albeit with varying input types (Fig. 5).
+6.1.1. Image Classification
 Definition: The task of Image Classification requires classifying an image into one of several predefined categories.
-In the comics domain, the task can be applied at panel or page level.
-In both cases, the classes vary from artist names, comic/manga styles or the image type (whether it’s an ads or front page or story page).
+In the domain of comics, the task can be applied at the panel or page level.
+In both cases, the classes vary from artist names, comic/manga styles, or image type (whether it’s an ad, a front page, or a story page).
 Early efforts in this domain, such as those by Chu et al.
-[31], focused on classifying manga panels by artistic style using Support Vector Machines (SVM) and manga-specific feature vectors.
+(Chu and Chao, 2014), focused on classifying manga panels by artistic style using Support Vector Machines (SVM) and manga-specific feature vectors.
 Their feature vectors are built on edge-detected lines creating a 20-dimensional vector from elements such as angle, orientation, the density of segments, etc.
 Hiroe et al.
-[83] employed a novel method using SVMs to classify comic books based on the frequency of exclamation marks, demonstrating marks-per-page histogram as an innovative approach to content approximation.
+(Hiroe and Hotta, 2017) employed a novel method using SVMs to classify comic books based on the frequency of exclamation marks, demonstrating marks-per-page histogram as an innovative approach to content approximation.
 Daiku et al.
-[84] expanded the scope to page-level classification within manga stories using Convolutional Neural Networks (CNN), distinct from Page Stream Segmentation (discussed in Subsection 6.1.4).
+(Daiku et al., 2017) expanded the scope to page-level classification within manga stories using Convolutional Neural Networks (CNN), distinct from Page Stream Segmentation (discussed in Subsection 6.1.4).
 Jiang et al.
-[85] introduced the Consensus Style Centralizing Auto-Encoder (CSCAE) for style classification, particularly distinguishing between Shonen (for boy) and Shojo (for girl) manga styles using robust style feature representations.
+(Jiang et al., 2018) introduced the Consensus Style Centralizing Auto-Encoder (CSCAE) for style classification, particularly distinguishing between Shonen (for boy) and Shojo (for girl) manga styles using robust style feature representations.
 They used low-rank and group sparsity constraints for consensus for the classification of manga character faces.
-Lastly, in [86] Terauchi et al.
+Lastly, in (Terauchi et al., 2019) Terauchi et al.
 presented an approach for classifying manga based on the author’s unique style, employing a Variational Autoencoder (VAE) for this purpose.
 They also proposed a “four-scene comics story dataset” for this classification task into Mow, Seinen, and Shonen touches (styles).
-In a more recent work [87], the authors tackled the challenge of (multi-)genre classification at the story level in Manga.
-They introduced a Panel-Page-aware Comic genre classification model, which takes page sequences as input and produces class-wise probabilities, relying on the attention mechanism to jointly attend page features and panel boxes, then pooled into a transformer encoder.
+In a more recent work (Xu et al., 2023), the authors tackled the challenge of (multi-)genre classification at the story level in Manga, introducing a Panel-Page-aware Comic genre classification model that takes page sequences as input and produces class-wise probabilities.
+Their proposal relyied on the attention mechanism to jointly attend page features and panel boxes, then pooled into a transformer encoder.
 The transformer classification token is merged with a Graph Convolution Network (GCN) over the labels graph, which predicts the probability distribution of the labels for the comic stories.
-6.1.2 Emotion Classification
+6.1.2. Emotion Classification
 Definition: Emotion Classification in comics involves categorizing a single panel or an entire page (looking both at the image and the text) into predefined emotional categories, enhancing the understanding of the narrative’s emotional context.
-A seminal study by Tanaka et al.
-[88] explored the interplay between speech balloon shapes and emotions in comics.
-They analyzed the relationship between balloon styles and the emotions expressed within them, using the Manga109 dataset as a foundation for their study.
-The EmoReCon challenge at ICDAR 2021 [62] marked a significant advancement in this field.
-Participants were tasked with classifying 8 emotions from comic panels, using a dataset specifically designed for this purpose [62].
-Notably, one of the winning teams employed a three-level fusion approach (early, mid, and late), integrating EfficientNetB3 [89] and RoBERTa [90] for visual and textual embeddings, respectively.
-This mid-fusion strategy outperformed other architectural configurations, signifying the potential of multimodal fusion.
-The second winning team proposed an early fusion approach with ResNet50 image features and OCR-extracted text as input for a transformer encoder BERT-style.
-In this case, the prediction is drawn from a weighted average over the layers’ cls tokens, with comparable results with the previous method.
-Not only have the emotions of comic characters been analyzed, but in [91] authors studied readers’ emotions through physiological signals.
-Through the relationship between a change in the reader’s physiological signal and the page of the reading, authors classify the manga across three classes: comedy, romance, and horror.
-A recent work [92] introduced two noteworthy datasets derived from Manga109: MangaAD+ and MangaEmo+.
-While MangaAD+ is a curated collection of 600 pages featuring unannotated elements such as shop names and ads - in Manga109 only the balloon text and onomatopoeias are annotated -, MangaEmo+ extends this dataset by annotating each page with a range of 8 emotions for multi-label classification.
-They employed a modified version of Faster R-CNN, trained on augmented text data, for atypical text detection.
-The subsequent emotion classification task was approached using EfficientNet, which extracted visual features combined with the above custom text detection model.
-However, neither of the two datasets has been made publicly available.
-6.1.3 Action Detection
+Tanaka et al.
+(Tanaka et al., 2015) conducted a seminal study on the relationship between speech balloon shapes and emotions in comics, using the Manga109 dataset to analyze how different balloon styles convey emotions.
+The EmoReCon challenge at ICDAR 2021 (Nguyen et al., 2021b) pushed the field forward by tasking participants with classifying 8 emotions from comic panels using a specially designed dataset.
+One winning team employed a three-level fusion strategy (early, mid, and late), combining EfficientNetB3 (Tan and Le, 2019) for visual embeddings and RoBERTa (Liu et al., 2019) for textual embeddings.
+Their mid-fusion approach outperformed other methods, showcasing the power of multimodal fusion.
+Another top team used an early fusion method, integrating ResNet50 image features with OCR-extracted text processed by a transformer encoder.
+They predicted emotions using a weighted average of cls tokens, achieving results comparable to the previous team.
+Beyond character emotions, Sanches et al.
+(Sanches et al., 2016) analyzed readers’ emotions by tracking physiological signals.
+By monitoring changes in the reader’s physiological responses, they classified manga into three genres: comedy, romance, and horror.
+Yang et al.
+(Yang and Chu, 2023) introduced two new datasets, MangaAD+ and MangaEmo+, based on Manga109.
+MangaAD+ adds unannotated elements like shop names and ads, while MangaEmo+ includes annotations for 8 emotions on each page for multi-label classification.
+Using a modified Faster R-CNN for text detection and EfficientNet for visual features, they tackled the emotion classification task, though neither dataset has been made publicly available.
+6.1.3. Action Detection
 Definition: Action Detection in comics is characterized by the analysis of sequences of consecutive panels.
@@
 Unlike single image analysis, this task requires the model to interpret a series of images, each adding context and continuity to the narrative.
-Despite being popular for Videos [93], this task has not yet been explored in comics apart from image-cloze task [8] where the model is challenged to select the correct next panel from two options.
-This task, while elementary, underscores the potential for more sophisticated Action Detection in comics, a domain where sequential imagery plays a pivotal role.
-6.1.4 Page Stream Segmentation
-Definition: Page Stream Segmentation (PSS) in comics entails segmenting a comic book into different stories or sections.
-In its simplistic version, the task can be seen as categorizing each page with a tag among cover, regular page, ads, or last page.
-In its more advanced form, it includes extracting salient information from the stories such as title, authors, drawer, price, characters, and synopsis.
-The concept of PSS, while not new in document analysis, has seen various methodologies employed over time.
-Traditional approaches, such as those utilizing Support Vector Machines (SVM) for single page classification [94], have evolved with the advent of Convolutional Neural Networks (CNNs) [95], and more recently, attention-based multimodal fusion techniques [96].
-These methods have been pivotal in processing and understanding page layouts and structures in various document formats.
-Despite its longstanding presence in document analysis, PSS remains largely uncharted in the comics domain.
-Comics present a unique challenge for PSS due to their complex structure; a single volume can interweave multiple stories, interspersed with advertisements and narrative text.
-This intricacy demands a nuanced approach to segment and understand the diverse content within comic books.
-A potential advancement in PSS for comics involves not just segmenting the pages but also tagging these segments with metadata like story titles, authors, and synopses.
-This represents a promising field of study, extending traditional image classification into a more intricate, story-centric context.
-6.2 Augmentation
-Augmentation tasks, in the context of comics, involve image manipulation techniques aimed at enhancing or altering the visual presentation.
+Despite being popular for Videos (Zhu et al., 2020), this task has not yet been explored in comics apart from image-cloze task (Iyyer et al., 2016) where the model is challenged to select the correct next panel from two options.
+In the context of clozure, many works have proposed valid solutions, starting from LSTM-based original solution of (Iyyer et al., 2016), the BERT-style multimodal and multigranularity embedding of (Soykan et al., 2024a), and the Multimodal T5-based approach of (Vivoli et al., 2024d).
+However, while elementary, this task underscores the potential for more sophisticated Panel sequence analysis (and Action Detection) in comics, a domain where sequential imagery plays a pivotal role.
+Recently, (Ikuta et al., 2024) propose panel-level benchmarks about text and character recognition, as well as visual-cloze tasks in manga-style comics benchmarking recent open-source MLLMs models.
+6.1.4. Page Stream Segmentation
+Definition: Page Stream Segmentation (PSS) in comics involves dividing a comic book into different sections or stories.
+In its simplest form, this entails categorizing pages with tags like cover, regular page, ads, or last page.
+In a more advanced form, PSS extracts key information such as story titles, authors, illustrators, prices, characters, and synopses.
+Though PSS is not a new concept in document analysis, it has seen various methods evolve over time.
+Traditional approaches like Support Vector Machines (SVM) for single-page classification (Gordo et al., 2013) have given way to Convolutional Neural Networks (CNNs) (Wiedemann and Heyer, 2017) and more recent attention-based multimodal fusion techniques (Demirtacs et al., 2022), which have significantly improved the understanding of page layouts and structures in documents.
+However, PSS remains largely unexplored in the comics domain, where the task is complicated by the unique structure of comics.
+A single volume may contain multiple interwoven stories, ads, and narrative text, creating a more complex segmentation challenge.
+This complexity requires more sophisticated methods for segmenting and understanding the diverse content of comic books.
+An emerging area of interest in PSS for comics is not just segmenting the pages but also tagging each segment with metadata such as story titles, authors, and synopses.
+Figure 6. Illustration of Augmentation tasks.
+This expands traditional image classification into a more detailed, story-oriented approach, offering a promising field of study.
+6.2. Augmentation
+Augmentation tasks, in the context of comics, involve image manipulation techniques aimed at enhancing or altering the visual presentation (Fig.
+6).
 The primary tasks identified under this category are Super-Resolution (SR), Style-Transfer (ST), Vectorization, and Depth Estimation.
 Each of these tasks revolves around processing a single image input to produce an enhanced or stylistically altered output.
-6.2.1 Super-Resolution
+6.2.1. Super-Resolution
 Definition: The Super-Resolution task in comics primarily targets the enhancement of image quality, especially for images that are compressed, photograph-based, or scanned.
@@
 Yao et al.
-[97] employ deep learning to classify and tailor SR models for manga screentone, ensuring semantic preservation.
+(Yao et al., 2023) employ deep learning to classify and tailor SR models for manga screentone, ensuring semantic preservation.
 Dai et al.
-[98] introduce the Structured Fusion Attention Network (SFAN), which optimizes feature extraction and reconstruction quality through attention modules, achieving superior performance on datasets like Manga109.
+(Dai et al., 2022) introduce the Structured Fusion Attention Network (SFAN), which optimizes feature extraction and reconstruction quality through attention modules, achieving superior performance on datasets like Manga109.
 These methods enhance image quality while maintaining the artistic nuances critical to manga, effectively balancing technical efficiency with artistic integrity.
-6.2.2 Style-Transfer
+6.2.2. Style-Transfer
 Definition: Style-transfer in comics encompasses a range of image-to-image modifications, including photo-to-comic transformation, image-to-image translation, and colorization, characterized by similar input and output attributes.
@@
 Recent advancements by Zhang et al.
-[99] utilize deep learning for semantic region segmentation in cartoon images, followed by screentone application.
+(Zhang et al., 2021) utilize deep learning for semantic region segmentation in cartoon images, followed by screentone application.
 However, these methods are primarily effective for color-based inputs, with limited applicability to line-based images.
 Image-to-Image Translation: Topal et al.
-[100] explored style transfer as a transfer for the detection task.
+(Topal et al., 2022) explored style transfer as a transfer for the detection task.
 Specifically, they used style transfer networks like CycleGAN and CartoonGAN to train a detection network, which was then fine-tuned on specific comic datasets for character and face detection.
-This approach, to use style transfer techniques to enhance performances in other tasks, is not new [68, 101].
-Figure 5: Colorization vs. Screening
-Colorization: Colorization in comics, specifically in manga, represents the task of transforming black-and-white images into full-color pages.
-It is a complex process that requires a balance of artistic understanding and technical precision.
+This approach -using style transfer techniques to enhance performances in other tasks- has also been explored in different applications (Inoue et al., 2018; Deng et al., 2020).
+In particular, the multi-task learning method for dense predictions in comic panels has been used as a medium for comic panel reconfiguration (Bhattacharjee et al., 2023) and depth estimation (Bhattacharjee et al., 2021).
+Colorization: Colorization in comics, particularly manga, involves transforming black-and-white images into full-color pages, requiring both artistic precision and technical expertise.
 Furusawa et al.
-[102] introduced a comprehensive approach to the colorization of manga.
-Their method allows for user interaction, where color dots can be added to guide the colorization process.
-This approach not only automates the colorization but also gives users the flexibility to influence the final output, ensuring that the automatic colorization aligns with the intended aesthetic vision.
-In a different approach, Hensman et al.
-[103] presented a model based on Conditional Generative Adversarial Networks (cGANs).
-Their model, specifically tuned for each reference image, colorizes manga panels to mimic colorization styles.
-This method implies training a new GAN for each reference image, making it versatile for different styles but requiring significant computational resources for each new style adaptation.
-Subsequently, in [104] authors investigated screentone synthesis, a critical aspect of manga art that involves translating line drawings into screentone-patterned images.
-They utilized Manga109 for comparison with standard image-to-image translation methods like pix2pix [105].
-Their approach, although similar in architecture to U-Net used in previous studies [106], stands out by explicitly considering screentone labels, thus preserving the integrity of screentone patterns in the final output.
-Moving towards a more advanced solution, Golyadkin et al.
-[107] introduced a two-stage architecture for manga page colorization.
-The first stage involves drafting the coloring using the Pixel2Style2Pixel architecture [108], followed by a conditional GAN in the second stage to refine the output.
-This method elegantly fuses color drafts with color hints from users, generating high-quality colorizations.
-They demonstrated improvements both qualitatively and quantitatively, indicating a significant advancement in manga colorization technology.
-Building upon previous methods, Jiramahapokee et al.
-[109] further refined the colorization process.
-They leveraged a multi-encoder Variational Autoencoder (VAE) for improved color region consistency and combined it with the shading strengths of other tools like manga-colorization-v2.
-Their approach also incorporated CIELAB [110] interpolation to enhance color saturation and authenticity, leading to outputs that are not only aesthetically pleasing but also consistent and detailed in their shading quality.
+(Furusawa et al., 2017) introduced a method that allows users to add color dots to guide an automated process, giving users the flexibility to influence the final output while ensuring the system aligns with their vision.
+Hensman et al.
+(Hensman and Aizawa, 2017) used Conditional GANs (cGANs) to colorize manga panels by tailoring the model to each reference image, allowing versatile style adaptations but requiring a new GAN for every image, increasing computational demands.
+Tsubota et al.
+(Tsubota et al., 2019) explored
+Figure 7. Colorization vs. Screening
+screentone synthesis, essential for manga, by comparing image translation techniques like pix2pix (Isola et al., 2016) on the Manga109 dataset.
+Their approach, using a U-Net architecture (Li et al., 2017), stands out by explicitly considering screentone labels, preserving the screentone patterns integral to manga art.
+Golyadkin et al.
+(Golyadkin and Makarov, 2023) introduced a two-stage method for manga page colorization: Pixel2Style2Pixel (Richardson et al., 2020) for color drafts, followed by a conditional GAN to refine the output.
+This method elegantly combines user-provided color hints and draft colorizations, yielding high-quality results that surpass previous methods.
+Jiramahapokee et al.
+(Jiramahapokee, 2023) further refined the process using a multi-encoder Variational Autoencoder (VAE) to ensure color consistency and integrated CIELAB interpolation (Zeyen et al., 2018) to enhance color saturation and authenticity.
+Their approach produced detailed and visually consistent colorizations, particularly in shading quality.
 Screening: Screening, or adding screentone to black-and-white manga drawings, has been a focus of several studies.
 A notable contribution by Wu et al.
-[111] introduced a two-stage architecture for this process.
-The first stage involves generating grayscale shading using a six-layer Swin-transformer [112], which then serves as input for the second stage, a reference-based screentone generation module.
+(Wu et al., 2023) introduced a two-stage architecture for this process.
+The first stage involves generating grayscale shading using a six-layer Swin-transformer (Liu et al., 2021a), which then serves as input for the second stage, a reference-based screentone generation module.
 This algorithmic process extracts screentone patterns from a reference manga and applies them to line drawings, considering the shading nuances provided by the first stage.
 These advancements in colorization represent a significant contribution to the domain of comics, particularly manga, where colorization can profoundly impact the narrative and aesthetic appeal of the stories.
-6.2.3 Vectorization
+6.2.3. Vectorization
 Definition: Vectorization in the context of comics, particularly manga, involves transforming raster images into vector formats.
@@
 In this task, Li et al.
-[106] developed a CNN-based method using a U-Net architecture with residual blocks to effectively extract structural lines from manga pages.
+(Li et al., 2017) developed a CNN-based method using a U-Net architecture with residual blocks to effectively extract structural lines from manga pages.
 This method is adaptable for various applications, including manga retargeting and colorization, and can handle diverse patterns and textures, resulting in clean skeleton sketches.
-Manga Vectorization: Various works [113, 114, 115, 116] have proposed different approaches for this task.
+Manga Vectorization: Various works (Kopf and Lischinski, 2011; Selinger, 2003; Ma et al., 2022) have proposed different approaches for this task.
 In particular, Yao et al.
-[117] developed a method involving adaptive binarization and screentone detection, refining borders, estimating lighting, and compensating for missing strokes, leading to high-quality, resolution-independent rendering.
+(Yao et al., 2017) developed a method involving adaptive binarization and screentone detection, refining borders, estimating lighting, and compensating for missing strokes, leading to high-quality, resolution-independent rendering.
 A recent innovative approach, MARVEL by Su et al.
-[118], employs Reinforcement Learning to train an agent in selecting drawing primitives that recreate raster manga images in vector format.
+(Su et al., 2024), employs Reinforcement Learning to train an agent in selecting drawing primitives that recreate raster manga images in vector format.
 While effective, this approach works with squared manga patches and may result in visual gaps in certain renderings.
@@
 MARVEL represents an interesting exploration into the use of machine learning for vectorization, highlighting both the potential and the complexities of this task.
-6.2.4 Depth Estimation
+6.2.4. Depth Estimation
 Definition: Depth Estimation within the comic domain refers to the process of predicting the spatial distance between objects and the viewer in a two-dimensional comic panel.
 It serves a crucial purpose in the reconfiguration of comic panels and unlocks potential developments in digital artistry.
-Up to now, this necessity arose particularly when adapting comics for various digital media, such as smartphones, tablets, and e-readers, where the original layout may not fit the display dimensions optimally.
-The primary methodologies for reconfiguring comic panels involve: (i) Adjusting panel proportions and rescaling them to fit the new medium without compromising the narrative or visual integrity.
-(ii) Strategically cropping panels while ensuring that key objects within each panel are preserved and remain visible in the adapted format.
-Both techniques rely heavily on depth estimation to identify and prioritize the most significant elements in each panel.
-One significant challenge in this domain is the lack of ground truth data for comics.
+To date, this need has arisen primarily when adapting comics for digital media like smartphones, tablets, and e-readers, where the original layout may not fit the screen optimally.
+Key methods for reconfiguring panels include: (i) Adjusting panel proportions and rescaling them to fit the new medium while maintaining narrative and visual integrity, and (ii) Cropping panels to preserve essential elements.
+Both approaches rely on depth estimation to identify and prioritize key components.
+A major challenge in this area is the lack of ground truth data for comics.
 To address this, Bhattacharjee et al.
-[7] adopted an innovative approach using Image-to-Image style-transfer.
-They transformed comic images to resemble natural images, upon which depth estimation techniques, involving Laplacian edges and feature-based GANs, were applied to predict depth effectively.
-Building on this, a subsequent study [119], based on their earlier work [120], introduced a Swin Transformer for multitasking, including both depth estimation and object segmentation.
-This multi-decoder-head architecture, while effective, presents challenges in scalability due to the complexity of accommodating multiple tasks.
-Depth Estimation plays a pivotal role in adapting comics to the varying formats of digital consumption.
-By ensuring that key visual elements are retained and appropriately displayed, these methodologies not only enhance the reader’s experience but also preserve the artistic intent of the original work.
-6.3 Satellite Tasks
+(Bhattacharjee et al., 2021) employed Image-to-Image style-transfer, transforming comic images to resemble natural ones, enabling depth estimation using Laplacian edges and feature-based GANs.
+Building on this, a follow-up study (Bhattacharjee et al., 2023), based on their earlier work (Bhattacharjee et al., 2022), introduced a Swin Transformer for multitasking, including depth estimation and object segmentation.
+However, the multi-decoder-head architecture poses scalability challenges due to task complexity.
+Depth estimation is crucial for adapting comics to various digital formats, ensuring key visual elements are retained while preserving the original artistic intent.
+6.3. Satellite Tasks
 Some of the tasks that fit under the Augmentation umbrella have not found a space in our taxonomy, e.g.
@@
 Addressing the de-warping challenge, a recent study by Garai et al.
-[121] proposed a novel mathematical model to describe the warping process.
+(Garai et al., 2022) proposed a novel mathematical model to describe the warping process.
 This model estimates distortion factors based on the boundaries of panels in a comic document image, aiming to rectify these warps effectively.
 While their approach yielded promising qualitative results, particularly in the image areas, challenges persist in handling text distortion, where some distortions proved unrecoverable.
-7 Layer 2: Grounding, Analysis, and Segmentation
+7. Layer 2: Grounding, Analysis, and Segmentation
 The second layer of our LoCU framework comprises three distinct task groups: grounding, analysis, and segmentation.
 These tasks dig deeper into the intricate components of comics, focusing on detailed elements like panels, text, and characters and their ordering and associations.
-7.1 Grounding
-Grounding in comics involves identifying and classifying specific elements within a comic’s panel, such as text, characters, and objects.
+7.1. Grounding
+Figure 8. Illustration of Grounding tasks.
+Grounding in comics involves identifying and classifying specific elements within a comic’s panel, such as text, characters, and objects (Fig.
+8).
 Under this task name umbrella, many individual tasks belong such as Detection, Character Re-identification, and sentence Grouding.
-7.1.1 Detection
+7.1.1. Detection
 Definition: In Detection task, given an image and a set of class names (or tags), the goal is to find all instances of those classes in the image and locate them with bounding boxes.
@@
 Initial studies in digital Manga, such as those by Ponsard et al.
-[122] and Arai et al.
-[123], emphasized automated segmentation and panel detection, underscoring the importance of maintaining reading order.
+(Ponsard, 2009) and Arai et al.
+(Arai and Tolle, 2010), emphasized automated segmentation and panel detection, underscoring the importance of maintaining reading order.
 Arai et al.
-[124] further developed these ideas with a focus on real-time detection and extraction techniques, which evolved into more sophisticated SVG-based approaches as explored in [125].
-[126] contributed advanced methods for panel extraction using region growing techniques and mathematical morphology, combined with speech balloon detection.
+(Arai and Tolle, 2011) further developed these ideas with a focus on real-time detection and extraction techniques, which evolved into more sophisticated SVG-based approaches as explored in (Su et al., 2011).
+(Ho et al., 2012) contributed advanced methods for panel extraction using region growing techniques and mathematical morphology, combined with speech balloon detection.
 Text.
 Jomaa et al.
-[127] employed a tracking algorithm for panel extraction, while speech balloons were identified using Robert’s edge detection operator and classified with thresholding.
+(Jomaa et al., 2015) employed a tracking algorithm for panel extraction, while speech balloons were identified using Robert’s edge detection operator and classified with thresholding.
 Liu et al.
-[128] focused on balloon extraction based on text character locations.
+(Liu et al., 2015b) focused on balloon extraction based on text character locations.
 Pal et al.
-[129] performed text extraction using SIFT local features, Space Pyramid Matching (SPM), and SVM on eBDtheque and Bangla comics images.
-In [130] authors contributed to text extraction in Manga109 combining region proposal generation with SVM classifiers.
+(Pal et al., 2016) performed text extraction using SIFT local features, Space Pyramid Matching (SPM), and SVM on eBDtheque and Bangla comics images.
+In (Aramaki et al., 2016) authors contributed to text extraction in Manga109 combining region proposal generation with SVM classifiers.
 Chu et al.
-[131] and Qin et al.
-[34] modified existing architectures like Faster R-CNN for detecting text and faces in comics.
-In particular, [131] proposed additional aspect ratio regions, while [34] built two comics datasets, namely JC2463 (Manga black and white style) and AEC912 (colored American comics style) to test face detection evaluation.
+(Chu and Yu, 2018) and Qin et al.
+(Qin et al., 2017) modified existing architectures like Faster R-CNN for detecting text and faces in comics.
+In particular, (Chu and Yu, 2018) proposed additional aspect ratio regions, while (Qin et al., 2017) built two comics datasets, namely JC2463 (Manga black and white style) and AEC912 (colored American comics style) to test face detection evaluation.
 Face.
-Interestingly, in [132] authors benchmarked Fast R-CNN, Faster R-CNN and SSD in panel, character, face, and balloon detection.
+Interestingly, in (Yanagisawa et al., 2018) authors benchmarked Fast R-CNN, Faster R-CNN and SSD in panel, character, face, and balloon detection.
 They used the Manga109 dataset, discovering that Fast R-CNN works better in panels when there are clear boundaries, Faster R-CNN when boundaries are not well defined, and SSD needs already cut panels as it is limited for full-page complex mangas.
 So far, many methods have used classical detection architectures and applied them to the comic domain to extract elements such as panels, characters, text, etc.
-A new direction is drawn by [133] challenged the conventional four-edge polygon representation of comic panels, suggesting segmentation-based approaches using U-Net for more accurate extraction.
+A new direction is drawn by (Nguyen et al., 2019b) challenged the conventional four-edge polygon representation of comic panels, suggesting segmentation-based approaches using U-Net for more accurate extraction.
 Here, authors performed extraction through segmentation, employing a U-Net that classifies pixels into background, panels, and borders.
@@
 The release of the COO dataset by Baek et al.
-[4] has paved the way for new research in this domain.
-Authors proposed a modified version of M4C [134] designed to link truncated onomatopoeias together.
+(Baek et al., 2022) has paved the way for new research in this domain.
+Authors proposed a modified version of M4C (Hu et al., 2019) designed to link truncated onomatopoeias together.
 Following this, Yang et al.
-[92] proposed text augmentation techniques for onomatopoeias using EfficientNet.
+(Yang and Chu, 2023) proposed text augmentation techniques for onomatopoeias using EfficientNet.
 Sharma et al.
-[135] furthered this research by introducing their dataset for panel and character detection, employing a Faster R-CNN model.
+(Sharma and Kukreja, 2023) furthered this research by introducing their dataset for panel and character detection, employing a Faster R-CNN model.
 However, they did not provide comparative analysis with other models.
@@
 Lately, Dutta et al.
-[136, 69] introduced the BCBId dataset, benchmarking existing datasets (eBDtheque, Manga109, and DCM) and employing YOLO architectures for panel and character detection.
+(Dutta and Biswas, 2019; Dutta et al., 2022) introduced the BCBId dataset, benchmarking existing datasets (eBDtheque, Manga109, and DCM) and employing YOLO architectures for panel and character detection.
 They presented comparisons with existing methods, showing enhanced performance in detecting panels and characters across multiple datasets.
@@
 Only recently, Vivoli et al.
-[2] fairly compares models across various datasets and styles, unifying annotations for multiple detection classes and providing model weights, code, and data to replicate.
-7.1.2 Character Re-Identification
+(Vivoli et al., 2024c) fairly compare models across various datasets and styles, unifying annotations for multiple detection classes and providing model weights, code, and data to replicate.
+7.1.2. Character Re-Identification
 Definition: Character Re-Identification, also known as character retrieval in comics analysis, has evolved from focusing on visual features alone to incorporating textual information alongside character images.
-Initial studies by [63, 29] employed techniques like PHOG and SIFT keypoints for character retrieval.
+Initial studies by Khan et al.
+(Khan et al., 2012) and Amirshahi et al.
+(Amirshahi et al., 2012) employed techniques like PHOG and SIFT keypoints for character retrieval.
 Iwata et al.
-[137], inspired by [138], adapted these methods specifically for manga characters, considering characters rather than entire page ROIs.
-Authors further explored applications of these techniques in [139].
-However, recently, Li et al.
-[5] noted the limitations of handcrafted features, particularly their lack of robustness to variations in pose, expression, and character shape.
-In [5], authors tackled manga character retrieval and verification.
-While retrieval is a known task 8.1, verification refers to the task of classifying whether two images (a pair) belong to the same character or not.
-These two tasks are challenging since the manga character images have a long-tailed distribution and large quality variations.
-While addressing challenges posed by long-tailed distribution and quality variations in manga characters, they proposed a novel dual loss approach, combining dual ring loss and dual adaptive re-weighting loss, to overcome biases introduced by imbalanced training data.
-In animated cartoons, where characters undergo shape and color changes, [140] introduced an approach based on Self-Supervised Learning (SSL) and Multi-object Tracking (MOT) to identify character clusters and link character proposals, facilitating character re-identification across different scenes.
-Zhang et al.
-[141] identified challenges unique to manga, such as similarities in faces but differences in bodies.
-They proposed the Face-body and Spatial-temporal Associated Clustering method (FSAC), which uses face-body graphs and spatial-temporal relationship correction to refine clustering and overcome artistic exaggerations and deformations, through the design of a temporal-spatial-related triplet loss to fine-tune the clustering.
-Recognizing the challenges of occlusion in manga, Zhang et al.
-[142] developed Occlusion-Aware Manga Character Re-identification (OAM-ReID) with self-paced contrastive learning.
-By synthesizing data with occluded speech balloons and incomplete bodies, their framework becomes more effective in learning representative features of characters.
-They introduce a combination of contrastive, occlusion, and re-identification losses, training a vanilla ViT and TransReID as baselines on 10 Manga109 books.
+(Iwata et al., 2014), inspired by Sun et al.
+(Sun and Kise, 2011), adapted these methods for manga characters, focusing on character-level identification rather than entire page regions, and further explored their applications in (Iwata et al., 2015).
+However, Li et al.
+(Li et al., 2021) highlighted the limitations of these handcrafted features, particularly their sensitivity to pose, expression, and shape variations.
+Li et al.
+(Li et al., 2021) addressed manga character retrieval and verification, the latter being the task of classifying whether two images belong to the same character.
+They introduced a dual loss approach—combining dual ring loss and dual adaptive re-weighting loss—to address long-tailed distributions and quality variations in manga data, improving robustness.
+In animated cartoons, where characters frequently change shape and color, Nir et al.
+(Nir et al., 2022) proposed a Self-Supervised Learning (SSL) and Multi-object Tracking (MOT) approach to cluster and link character proposals, facilitating character re-identification across scenes.
+Similarly, Zhang et al.
+(Zhang et al., 2022) tackled the challenges of character re-identification in manga, where faces are similar but body features differ.
+They introduced FSAC (Face-body and Spatial-temporal Associated Clustering), which refines clustering through face-body graphs and spatial-temporal relationships, using a triplet loss to handle artistic exaggerations.
+To address occlusion in manga, Zhang et al.
+(Zhang and Chu, 2023) developed Occlusion-Aware Manga Character Re-identification (OAM-ReID), using self-paced contrastive learning.
+Their model synthesizes occluded data, such as speech balloons and incomplete bodies, and combines contrastive, occlusion, and re-identification losses, training a ViT and TransReID on 10 Manga109 books.
 Soykan et al.
-[143] leveraged SimCLR [144] with various data augmentation techniques for character re-identification, calling it “Identity-aware SimCLR”, focusing on both local (face) and global (body) representations.
+(Soykan et al., 2023) introduced ”Identity-aware SimCLR” for character re-identification, leveraging SimCLR (Chen et al., 2020) with data augmentations to capture both local (face) and global (body) features.
 Similarly, Sachdeva et al.
-[41] proposed a multi-task Transformer model (Relationformer [145]) for character clustering and linking detected character bounding boxes at the page level.
-Finally, [146] introduced an improved YOLOv5 and XGBoosting approach for object detection, achieving state-of-the-art results across multiple datasets (StyleObject7K, ClipArt1K, Watercolor2K, and Comic2K).
-Their method highlights the effectiveness of well-annotated training data and extensive data augmentation in detection tasks.
-7.1.3 Sentence-based Grounding
+(Sachdeva and Zisserman, 2024) proposed Relationformer (Shit et al., 2022), a multi-task Transformer model for character clustering and linking bounding boxes at the page level.
+Finally, Ahmad et al.
+(Ahmad and Schich, 2023) introduced an enhanced YOLOv5 combined with XGBoost for object detection, achieving state-of-the-art performance across multiple datasets, emphasizing the importance of well-annotated data and augmentation techniques in improving detection tasks.
+7.1.3. Sentence-based Grounding
 Definition: Sentence-based grounding is a Vision-Language task that links textual content to specific visual elements within a panel.
@@
 Moreover, the complexity of the task varies depending on what is the sentence to be grounded: a panel description or a dialogue or narrative description.
-In image and video Multimedia processing various works tackle the task with custom CNNs and LSTMs [147] approaches, in an unsupervised manner [148] or with custom multimodal fusion approaches [149, 40].
-However, the only work that used a similar approach (GroundingDINO) in a zero-shot setting is [41] which used it to automatically annotate the training set.
-7.2 Analysis
-Analysis tasks involve a deeper examination of the relationship between text and visual elements in comics, including character-text association, panel sorting, dialog transcription, and translation.
-7.2.1 Text-character association
+In image and video Multimedia processing various works tackle the task with custom CNNs and LSTMs (Li et al., 2019) approaches, in an unsupervised manner (Parcalabescu and Frank, 2020) or with custom multimodal fusion approaches (Shen et al., 2023b; Liu et al., 2023c).
+However, the only work that used a similar approach (GroundingDINO) in a zero-shot setting is (Sachdeva and Zisserman, 2024) which used it to automatically annotate the training set.
+7.2. Analysis
+Figure 9. Illustration of Analysis tasks.
+Analysis tasks involve a deeper examination of the relationship between text and visual elements in comics, including character-text association, panel sorting, dialog transcription, and translation (Fig.
+9).
+7.2.1. Text-character association
 Definition: The task of text-character association in comics aims to link textual content, usually dialogue, to the corresponding speaking character.
@@
 Rigaud et al.
-[150] tackled this task by using geometric graph analysis and anchor point selection.
-They developed a distance-based algorithm, both global (across the page) and single-frame (within a panel), relying on Euclidean distances and balloon tail information.
-Advancing from Rigaud et al.’s work, Nguyen et al.
-[151] employed a multitask learning approach with a custom Mask R-CNN (adding a new “PairPool” head for the association prediction) which they called ComicMTL.
-Their method simultaneously addressed panel and character detection, balloon segmentation, text recognition, and text-speaker association, demonstrating its effectiveness on the DCM772 and eBDtheque datasets.
-In [152], authors introduced a novel perspective by considering the continuity of comics.
-Their work focused on estimating the reading order of frames and texts, vital for associating text with off-panel speakers.
-However, their algorithmic approach raised concerns regarding scalability and generalizability.
+(Rigaud et al., 2015) addressed the text-character association task using geometric graph analysis and anchor point selection, developing a distance-based algorithm that utilized both global (page-level) and single-frame (panel-level) information, relying on Euclidean distances and balloon tail cues.
+Building on this, Nguyen et al.
+(Nguyen et al., 2019a) introduced ComicMTL, a multitask learning approach based on a modified Mask R-CNN with an added ”PairPool” head for association prediction.
+Their method tackled multiple tasks—panel and character detection, balloon segmentation, text recognition, and text-speaker association—demonstrating effectiveness on DCM772 and eBDtheque datasets.
+Omori et al.
+(Omori et al., 2022) took a novel approach by focusing on the continuity of comics, particularly reading order, to associate text with off-panel speakers.
+However, their approach raised scalability and generalizability concerns.
 Li et al.
-[73] extended the Manga109 dataset with the Manga109 dialog dataset, enabling more precise speaker detection.
-They classified text-character links into “easy” (character in the panel) and “hard” (character not in the panel) configurations, using Faster R-CNN for detection and subsequent association classification.
-The challenges in this approach, when balloons are spoken alternatively, highlight the potential role of NLP models in handling complex dialogue scenarios.
-In a recent study, Sachdeva et al.
-[41] proposed a multi-task Transformer-based model (Relationformer [145]) with specialized MLP heads for detection and linking tasks.
-While their model offers an “end-to-end” solution, it operates in two iterations: first for detection and then for linking characters and text.
-However, their model’s comparison with state-of-the-art methods, particularly in detection, is limited, highlighting the need for more comprehensive benchmarking.
-In fact, the model does not compare with previous state-of-the-art models (YOLO-based) for detection, but only with publicly available GroundingDino [40], in a zero-shot setting.
-The character-text association is a critical but challenging task in comics analysis, requiring a nuanced understanding of narrative structure and visual-textual interplay.
-Future research may benefit from integrating advanced NLP techniques to better interpret complex dialogue scenarios.
-Additionally, there is a growing need for unified benchmarks and replicable methodologies, as indicated by the diverse approaches in the LoCU framework, to facilitate comparative analysis and further advancements in this field.
-7.2.2 Panel Sorting
+(Li et al., 2023a) expanded the Manga109 dataset by introducing the Manga109 Dialog dataset, refining speaker detection.
+They categorized text-character links into “easy” (character in the panel) and “hard” (character off-panel), employing Faster R-CNN for detection and association classification.
+Despite progress, handling alternating speaker dialogue remains challenging, pointing to the potential of NLP models for complex scenarios.
+Sachdeva et al.
+(Sachdeva and Zisserman, 2024) recently proposed a multi-task Transformer model, Relationformer (Shit et al., 2022), with specialized MLP heads for detection and linking.
+Their two-step, end-to-end solution first detects and then links characters and text, though their model’s benchmarking, especially against YOLO-based models, remains limited, focusing only on GroundingDino (Liu et al., 2023c) in a zero-shot setting.
+Lastly, in (Soykan et al., 2024b) built on top of Comic-MTL (Nguyen et al., 2021a) a Mask R-CNN based model for detecting and linking textboxes and characters, but did not fully compared with previous models.
+Character-text association is a crucial yet challenging task in comics analysis, requiring an understanding of narrative and visual-textual interplay.
+Future research could benefit from advanced NLP techniques to manage complex dialogue scenarios, and more unified benchmarks and replicable methodologies are needed, as highlighted by the LoCU framework, to support comparative analysis and innovation in this field.
+7.2.2. Panel Sorting
 Definition: Sorting individual comic panels into their correct narrative order presents a unique challenge, both for humans and automated systems.
 Ueno et al.
-[153] conducted a study revealing that while people can relatively easily identify the first and last panels in a sequence, accuracy diminishes with the presentation of all panels from four-panel strips.
+(Ueno and Isahara, 2017) conducted a study revealing that while people can relatively easily identify the first and last panels in a sequence, accuracy diminishes with the presentation of all panels from four-panel strips.
 This suggests the complexity involved in discerning narrative continuity in comics.
@@
 However, the CNN’s performance was poor, indicating the difficulty of the task and the need for more advanced or specialized algorithms in panel sorting.
-7.2.3 Dialog transcription
+7.2.3. Dialog transcription
 Definition: Dialog transcription in comics aims to create an automatic transcript of dialogues, identifying the speaker and the sequence of speech.
 The process typically involves several steps: (i) detecting panels, text boxes, and character boxes, (ii) associating text with speakers, and (iii) sorting text boxes according to their reading order to generate a coherent dialogue transcript.
-This task has been addressed by [73] for Japanese Mangas, proposing the needed annotations for Manga109, and more recently by [41] for English Mangas called PopManga.
+This task has been addressed by (Li et al., 2023a) for Japanese Mangas, proposing the needed annotations for Manga109, and more recently by (Sachdeva and Zisserman, 2024) for English Mangas called PopManga.
 However, a few limitations arise from these approaches: sorting the text boxes is always done as a postprocessing operation, depending only on algorithmic approaches.
@@
 Recently, Sachdeva et al.
-[43] proposed Magiv2 which tackled the problem of dialog transcription adding also names (from a name bank) allowing it to operate on the whole collection of pages.
-In CoMix [42], a first benchmark for this task was proposed, authors enhanced the previous annotations and proposed a new metric for character naming (based on ANLS) and dialog transcription (based on Hungarian matching and edit distance).
-7.3 Segmentation
-Segmentation in comic analysis, particularly instance segmentation, presents a nuanced approach to identifying and isolating various elements within a comic panel.
-7.3.1 Instance segmentation
+(Sachdeva et al., 2024) proposed Magiv2 which tackled the problem of dialog transcription adding also names (from a name bank) allowing it to operate on the whole collection of pages.
+In CoMix (Vivoli et al., 2024a), a first benchmark for this task was proposed, authors enhanced the previous annotations and proposed a new metric for character naming (based on ANLS) and dialog transcription (based on Hungarian matching and edit distance).
+7.2.4. Translation
+Definition: The task of translation in comics involves not only linguistic challenges but also graphical and spatial ones, requiring an integrated approach that respects the visual and narrative structure of the comic.
+Text appears in various forms such as balloons, onomatopoeias, and scene text, each necessitating different strategies.
+For instance, Balloon Text must fit within the original speech balloon, preserving visual integrity, while Onomatopoeias may require both linguistic translation and graphic replacement to maintain stylistic consistency.
+Recently, several works have addressed these complexities by integrating multimodal information and context-aware techniques.
+(Hinami et al., 2020) were among the first to incorporate context obtained from manga images into a multimodal translation framework, developing a comprehensive system for fully automated manga translation.
+To improve text image translation performance, (Ma et al., 2023) proposed a Multi-Teacher Knowledge Distillation method, effectively distilling knowledge from a pipeline model into an end-to-end Text Image Machine Translation (TIMT) model.
+The advent of Large Language Models (LLMs) has further advanced the field; (Yang et al., [n.
+d.]) explored leveraging LLMs for manga translation tasks, demonstrating the potential of these models in handling complex scenarios involving both text and visual elements.
+Similarly, (Lippmann et al., 2024) proposed a methodology that leverages the vision component of multimodal LLMs to improve translation quality, investigating the impact of translation unit size and context length, and introducing a token-efficient approach along with a new evaluation dataset.
+Inspired by human translation workflows, (Singh et al., 2024) presented an approach to completely automate the process of translating graphic novels, addressing the challenges of automating different aspects of the translation workflow.
+Focusing on character identification and speaker prediction, (Li et al., 2024a) proposed an iterative multimodal framework that employs multimodal information for both tasks, allowing machines to identify characters and predict speaker names based solely on unannotated comic images.
+Furthermore, (Liu et al., 2024) introduced the task of character-centric story generation, presenting the first model capable of generating visual stories with consistently grounded and coreferent character mentions, emphasizing the importance of character consistency and coreference in narrative understanding and translation.
+Despite these tasks being tackled as image-to-text translation, they have origins in image-to-image translation due to possible changes in visual components influenced by cultural elements, as discussed by (Khanuja et al., 2024).
+This indicates that future research may focus more on end-to-end image-to-image translation and transcreation, where both textual and visual elements are adapted to enhance cultural relevance.
+7.3. Segmentation
+Figure 10. Illustration of Segmentation tasks.
+Segmentation in comic analysis, particularly instance segmentation, presents a nuanced approach to identifying and isolating various elements within a comic panel (Fig. 10).
+7.3.1. Instance segmentation
 Definition: Instance segmentation in comics extends beyond the scope of typical object detection.
 Rather than just identifying and classifying elements within bounding boxes, this task involves generating precise pixel-wise masks for each instance.
-By focusing on segmentation instead of detection, we can achieve more accurate panel separation and representation.
-This precision is especially beneficial when reconfiguring or reformatting comics for different layouts.
-It ensures that each panel is cleanly isolated, without the remnants of adjacent panels or the original layout, which is a common issue with less precise bounding boxes.
-One of the key challenges in instance segmentation within comics is dealing with the diversity of artistic styles and the complexity of layouts.
-Comics often feature intricate designs, overlapping elements, and varying levels of detail, which can complicate the segmentation process.
-As described in [47], there are dozens of configurations in which panels appear to respect each other (pure grid, vertical and horizontal staggering, bleed, inset, overlap, etc.).
-Addressing the task of separating panels and understanding their relative positions within Cohn nomenclature is hard with bounding boxes when delimitations are not precise.
-Future research in comic instance segmentation could explore advanced deep-learning models that are adept at handling these complexities.
-The development of specialized algorithms that can adapt to the unique characteristics of different comic styles would significantly advance the field.
-There’s also potential for integrating instance segmentation with other tasks like character-text association and panel sorting to create a more comprehensive comic analysis system.
-Across the presented datasets, no one proposes semantic segmentations.
-7.4 Future Tasks
-The Layer of Comics Understanding is a general framework with which we aim to collect various Vision-Language tasks.
-Therefore, it includes also tasks that are currently not explored in comics, such as Translation.
-7.4.1 Translation
-Definition: The task of translation in comics refers to translating the textual elements that appear in a panel or a page, according to the scene.
-It goes beyond simple text translation, intertwining linguistic, graphical, and spatial challenges.
-It necessitates an integrated approach that respects the visual and narrative structure of the comic.
-While translation in comics may seem like a straightforward NLP problem, it involves unique complications.
-Text in comics appears in various forms such as balloons, onomatopoeias, and scene text, each requiring a different approach.
-In Balloon Text, translated text must fit within the space of the original speech balloon, preserving the visual integrity of the panel.
-Complications arise when the text to be translated contains elements like sarcasm, irony, jokes, or idiomatic expressions.
-In Onomatopoeias, both linguistic translation and graphic replacement are often required to maintain the style consistent with the target language.
-Lastly, in Scene Text, translating involves both linguistic translation and potentially replacing the text within the scene, which can be a complex graphic design task.
-8 Layer 3: Retrieval and Modification
-Layer 3 in the Layers of Comics Understanding focuses into advanced image and text understanding through tasks such as Retrieval (uni-modal and multi-modal) and Modification.
-8.1 Retrieval
-8.1.1 Unimodal Retrieval
+Instance segmentation, as opposed to detection, offers more accurate panel separation and representation in comics, which is crucial when reconfiguring or reformatting them for different layouts.
+This precision ensures that each panel is cleanly isolated, avoiding remnants of adjacent panels or the original layout—a common issue with less precise bounding boxes.
+One of the key challenges in instance segmentation for comics lies in the diversity of artistic styles and the complexity of layouts.
+Comics often feature intricate designs, overlapping elements, and various configurations such as grids, staggered layouts, bleeds, and insets, as noted by Cohn (Cohn, 2023).
+These complex panel arrangements make it difficult to segment panels accurately using bounding boxes, especially when panel boundaries are not clearly defined.
+In a recent work (Kouletou et al., 2024), authors proposed a pipeline of detection models (YOLO) and segmentations (GroundingDINO) for segmenting objects in comic panels.
+Despite the idea being simple and elegant -using in a cascade approach a YOLO detector and prompting SAM with the detected bounding boxes- the approach has been only evaluated for the detection tasks as no Segmentation ground truth exists for the used datasets.
+Moreover, as noticed in (Vivoli et al., 2024c), results are not comparable due to different evaluation datasets/splits/criteria.
+Although current datasets do not include semantic segmentations, the 2024 “AI for Visual Art” workshop (Bhattacharjee and Aydemir, 2024) introduced a challenge on comic semantic segmentation, signaling the growing importance of this task in comic analysis.
+8. Layer 3: Retrieval and Modification
+Layer 3 in the Layers of Comics Understanding focuses on advanced image and text understanding through tasks such as Retrieval (uni-modal and multi-modal) and Modification.
+8.1. Retrieval
+Figure 11. Illustration of Retrieval tasks.
+The task of retrieval, in the domain of the comic, addresses two main retrieval settings in: unimodal and cross-modal retrieval, which do not differ, in practice, from the vision-language retrieval tasks (Fig.
+11).
+These tasks aim to retrieve an instance of the same modality of the query (unimodal) or a different modality from the query (cross-modal), using either text or images as the query.
+An additional task emerged recently in the realm of Vision-Language, called “Composed image retrieval”, which combines in the query both image and text.
+8.1.1. Unimodal Retrieval
 Definition: Unimodal retrieval in comics involves searching for a specific media element (image or text) within a database, using a query of the same format (image or text, respectively).
-Pioneering works in this area have explored various aspects of comic retrieval. Wu et al. [154] developed a comic retrieval system that leveraged OCR and external text.
-One significant article [138] focused on retrieving full pages or books containing similar elements.
+Pioneering works in this area have explored various aspects of comic retrieval.
+Wu et al.
+(Wu, 2010) developed a comic retrieval system that leveraged OCR and external text.
+One significant article (Sun and Kise, 2011) focused on retrieving full pages or books containing similar elements.
 This approach is particularly relevant in copyright violation contexts.
@@
 Notably, they applied HOG not to the entire images but selectively to Regions Of Interest (ROIs) identified in the comics, such as faces or specific markings.
-Building on this approach, authors in [137] utilized a similar methodology for manga character retrieval.
+Building on this approach, authors in (Iwata et al., 2014) utilized a similar methodology for manga character retrieval.
 They focused on identifying and extracting distinctive features from characters, enhancing the ability to retrieve specific characters from large databases.
-8.1.2 Cross-modal retrieval
+8.1.2. Cross-modal retrieval
 Definition: Cross-modal retrieval in comics encompasses tasks where the query and the retrieval results are from different modalities, such as using sketches or text queries for retrieving a comic image.
@@
 Matsui et al.
-[32] introduced a framework for retrieving manga images from sketches using the Fine Multi-scale Edge Orientation Histogram (FMEOH).
+(Matsui et al., 2014) introduced a framework for retrieving manga images from sketches using the Fine Multi-scale Edge Orientation Histogram (FMEOH).
 This method indexed different-sized squares on a page and enabled efficient retrieval from sketch queries.
-Further advancing this area, authors [155] proposed a two-step system that first processes manga and then retrieves images based on sketches.
+Further advancing this area, authors (Matsui et al., 2015) proposed a two-step system that first processes manga and then retrieves images based on sketches.
 This framework was one of the first to employ the Manga109 dataset and included margin labeling, objectness-based edge orientation histograms, and approximate nearest neighbor search.
 Narita et al.
-[156] explored CNN-based feature extraction for sketches and manga images (without and with screentone, respectively), demonstrating improvements over algorithmic approaches.
+(Narita et al., 2017) explored CNN-based feature extraction for sketches and manga images (without and with screentone, respectively), demonstrating improvements over algorithmic approaches.
 Multimodal Learning and Character Retrieval.
 Nguyen et al.
-[157] proposed a multitask multimodal approach for character image and text learning, aiming at tasks such as character retrieval and emotion recognition.
+(Nguyen et al., 2021a) proposed a multitask multimodal approach for character image and text learning, aiming at tasks such as character retrieval and emotion recognition.
 Their method focused on leveraging both visual and verbal information in manga content, in a CLIP-like style aiming at performing various character tasks (e.g.
 detection, retrieval, emotion classification), trained uniquely with visual information and verbal information in manga image content.
-Later, in [158] authors introduced ComicLib, a dataset for comic sketch research, providing benchmarks across various tasks like colorization, generation, and retrieval.
+Later, in (Wei et al., 2022) authors introduced ComicLib, a dataset for comic sketch research, providing benchmarks across various tasks like colorization, generation, and retrieval.
 They conducted extensive comparison experiments with other datasets (e.g.
-QuickDraw [159]) to provide a benchmark for ComicLib on common tasks like colorization (Style2Paints [160]), generation (DCGAN [161]), retrieval (image-to-image with ResNet), detection (with Faster R-CNN [162], YOLOv3 [163], and SSD [164]) and recognition.
+QuickDraw (Ha and Eck, 2017)) to provide a benchmark for ComicLib on common tasks like colorization (Style2Paints (Zhang et al., 2017)), generation (DCGAN (Radford et al., 2015)), retrieval (image-to-image with ResNet), detection (with Faster R-CNN (Ren et al., 2015), YOLOv3 (Redmon et al., 2015), and SSD (Liu et al., 2015a)) and recognition.
 However, the dataset’s availability is unclear.
@@
 Shen et al.
-[165] presented a system for text-to-image content retrieval within Manga frames.
+(Shen et al., 2023a) presented a system for text-to-image content retrieval within Manga frames.
 This multi-staged system integrates object detection, text recognition, and a vision-text encoder to facilitate efficient search and retrieval of dialogues and scenes.
-Their method is composed of an object detection model for identifying text and frame bounding boxes (DETR-based [166]), a Vision Encoder-Decoder model for text recognition (DiT [167]), a text encoder for embedding text (multilingual DistilBERT [168]), and a vision-text encoder with unified embedding space (CLIP [14]).
-They experimented with Japanese and GPT-4101010Using the website https://platform.openai.com/translate/rephrased English sentences, on the annotated panels from “Dollgun Book”.
+Their method is composed of an object detection model for identifying text and frame bounding boxes (DETR-based (Carion et al., 2020)), a Vision Encoder-Decoder model for text recognition (DiT (Touvron et al., 2020)), a text encoder for embedding text (multilingual DistilBERT (Sanh et al., 2019)), and a vision-text encoder with unified embedding space (CLIP (Radford et al., 2021)).
+They experimented with Japanese and GPT-4888Using the website https://platform.openai.com/translate/rephrased English sentences, on the annotated panels from “Dollgun Book”.
 Despite being a composition of existing systems on panel level, on a small distribution of data, the work is a first direction to more complex Comic retrieval.
-8.2 Modification
-In the realm of comics, modification tasks such as image inpainting and image editing play a crucial role in streamlining the storytelling process and enhancing creative expression.
-8.2.1 Image Inpainting and Editing
+8.2. Modification
+Figure 12. Illustration of Modification tasks.
+In the realm of comics, modification tasks such as image inpainting and image editing play a crucial role in streamlining the storytelling process and enhancing creative expression (Fig. 12).
+8.2.1. Image Inpainting and Editing
 Definition: Image inpainting and editing in the context of comics involve various techniques and tools that allow for the alteration or enhancement of comic panels, facilitating more dynamic and engaging storytelling.
-One innovative tool in this space is CodeToon, presented by [169].
-CodeToon is a comic authoring tool that facilitates a code-driven storytelling process.
-It leverages two mechanisms: story ideation from code through metaphorical interpretation and automatic comic generation from these ideated stories.
-Significant attention has been given to the interface design to address common challenges faced by comic creators, such as generating stories directly from code snippets and seamlessly transitioning these stories into comic format.
-Another notable contribution is ComicScript by [170], also known as interactive data comics [^11] Accessible at https://interactivedatacomics.github.io/..
-ComicScript is a suite of operations designed in collaboration with practitioners to create and modify comics.
-These operations allow for the addition and removal of panels, support branching narratives, perspective changes, and detailed exploration.
-They also facilitate data manipulation and interaction within the comics, enhancing personalization and reader engagement.
-To validate the utility of ComicScript, professional illustrators, designers, and data comics enthusiasts were recruited to craft interactive comics.
-Their experiences provided valuable insights into the authoring workflow and the potential of interactive comics as a storytelling medium.
-8.3 Future Tasks
-8.3.1 Personalized Image retrieval
-Definition: Personalized Image Retrieval (PIR) focuses on developing person-centric models that can efficiently retrieve specific comic panels based on a compound query involving both image and text inputs [171].
+In the realm of comic inpainting and editing, CodeToon (Suh and Zhao, 2022) offers a code-driven storytelling tool that automatically generates comics from code snippets through metaphorical interpretation.
+It simplifies the transition from story ideation to comic format, addressing common challenges faced by comic creators.
+ComicScript (Wang et al., 2021) provides a suite of operations for interactive comic creation, allowing users to add/remove panels, change perspectives, and support branching narratives.
+It also integrates data manipulation to enhance engagement.
+Professional creators validated its potential by crafting interactive comics, showcasing the tool’s ability to personalize storytelling.
+Both tools highlight the growing role of computational methods in enhancing comic creation and editing workflows.
+8.3. Future Tasks
+8.3.1. Personalized Image retrieval
+Definition: Personalized Image Retrieval (PIR) focuses on developing person-centric models that can efficiently retrieve specific comic panels based on a compound query involving both image and text inputs (Korbar and Zisserman, 2022).
 PIR aims to identify and retrieve images or videos that correspond to a compound query, which typically includes an image of a person’s face combined with a textual scene or action description.
@@
 This task becomes increasingly complex when the retrieval criteria include context-specific actions (e.g., a character eating pizza or jumping out of a window) or interactions with other characters.
-9 Layer 4: Advanced Visual-Language Understanding
-9.1 Understanding
-This layer investigates tasks that require a deeper integration of visual elements and language, advancing beyond basic identification or retrieval to more complex forms of comprehension and reasoning within comics.
-9.1.1 Visual Question Answering
+9. Layer 4: Advanced Visual-Language Understanding
+In the fourth layer, we propose in the Layer of Comics Understanding a set of tasks related to multimodal understanding, which comprehends often reasoning through a text-rich image given a text prompt and one or more images as input.
+However, the only task that has been explored in comics already is Visual Question Answering, which under its umbrella groups different definitions of VQA in comics.
+9.1. Understanding
+Figure 13. Illustration of Understanding tasks.
+This layer investigates tasks that require a deeper integration of visual elements and language, advancing beyond basic identification or retrieval to more complex forms of comprehension and reasoning within comics (Fig.
+13).
+9.1.1. Visual Question Answering
 Definition: Given an image-question pair, Visual Question Answer requires answering a question based on the image.
 Most studies treat VQA as a classification problem on a predefined answer set, but its general definition would be an open answer where no specific set of options is provided.
-The first notable attempt in the domain of the comic is by [172] with ComicVQA.
+The first notable attempt in the domain of the comic is by (Moriyama et al., 2016) with ComicVQA.
 However, their approach, more akin to retrieval based on specific “Doraemon” comics datasets, diverges from the typical VQA framework.
 Later, Sumi et al.
-[173] presented a system, ComicsQA, with a unique twist, converting user QA interactions into comic stories, reflecting the user’s situation and offering solutions.
+(Sumi, 2017) presented a system, ComicsQA, with a unique twist, converting user QA interactions into comic stories, reflecting the user’s situation and offering solutions.
 The creation of a comprehensive VQA dataset specifically for comics, encompassing single-panel, single-page, and multi-page formats, remains an open challenge.
-In a recent survey [174], Zeng et al.
-show the massive work that has been carried out on VQA on multi-modal data across different mediums (movies, comics, etc.) and Multimodal Machine Comprehension [175, 176].
-This, together with multiple datasets across various mediums [177, 178, 179, 180, 181, 181] are laying the groundwork for advancing VQA also in comics.
-9.2 Future Tasks
-In the context of Understanding, a number of Vision Language tasks could be of interest in comics, both for pertaining and fine-tuning.
-9.2.1 Visual Entailment
+In a recent survey (Zeng et al., 2020), Zeng et al.
+show the massive work that has been carried out on VQA on multi-modal data across different mediums (movies, comics, etc.) and Multimodal Machine Comprehension (Sahu et al., 2021b, a).
+This, together with multiple datasets across various mediums (Singh et al., 2019; Biten et al., 2019; Mathew et al., 2022; Vivoli et al., 2022; Tanaka et al., 2023, 2023) are laying the groundwork for advancing VQA also in comics.
+9.1.2. Visual Entailment
 Definition: Visual entailment (VE) in comics is an area yet to be deeply explored.
@@
 It can be seen as a simplified version of the text-closure task where we aim at correctly classifying whether the image is represented by the given text, however, depending on the provided text, VE could quickly lead to complex scenarios.
-9.2.2 Visual Dialog
+Recently, (Hu et al., 2024) investigated the performances of Multimodal models on comics with contradictory narratives, where each comic consists of two panels that create a humorous contradiction.
+The paper introduces the YesBut benchmark and assesses Multimodal Models’ limited capabilities in this reasoning task.
+9.2. Future Tasks
+In the context of Understanding, a number of Vision Language tasks could be of interest in comics, both for pertaining and fine-tuning.
+9.2.1. Visual Dialog
 Definition: The task of Visual Dialogs (VisDial) corresponds to answering a specific question, having an image, a question about the image, and the previous dialog history about the image.
@@
 This task demands a comprehensive understanding of the storyline, character interactions, and visual symbolism to generate contextually relevant dialogues.
-9.2.3 Visual Reasoning
+9.2.2. Visual Reasoning
 Definition: Visual reasoning in comics extends beyond simple question answering to include the ability to interpret and analyze the visual content in depth.
@@
 It’s a sophisticated blend of visual comprehension and logical deduction, aimed at uncovering deeper layers of meaning in the comics.
-10 Layer 5: Generation and Synthesis
+10. Layer 5: Generation and Synthesis
 The last layer explores the creative frontier in comics analysis, where the synthesis and generation of comics from various media sources play a pivotal role.
-10.1 Generation
+10.1. Generation
+Figure 14. Illustration of Generation tasks.
 The field of Generation in comics has seen lots of attraction on Manga style comics rather than American ones.
-Their popularity has attracted the research community towards creating comics from various media such as video, images, and text descriptions.
-10.1.1 Comics generation from other media.
+Their popularity has attracted the research community towards creating comics from various media such as video, images, and text descriptions (Fig.
+14).
+10.1.1. Comics generation from other media.
 Definition: The task of generating comics from other media comprehends an umbrella of tasks with various input types (video, audio, charts, and comics itself) whose output is a comic panel, a sequence of panels, a page, or a book.
 One of the earliest attempts at comic generation from other media is the “Comic Live Chat” by Matsuda et al.
-[182], which transformed video meetings into comics.
+(Matsuda et al., 2010), which transformed video meetings into comics.
 This approach involved selecting keyframes and rendering the dialogues of meeting participants in a comic format.
 Tanapichet et al.
-[183] proposed a system for creating comic strips from cartoon animations using optical flow techniques, demonstrating a novel approach to narrative translation from one visual medium to another.
+(Tanapichet et al., 2011) proposed a system for creating comic strips from cartoon animations using optical flow techniques, demonstrating a novel approach to narrative translation from one visual medium to another.
 Hoashi et al.
-[184] worked on creating manga previews by detecting panels and text balloons, while in [185, 186] authors explored generating manga faces and caricatures from real images.
+(Hoashi et al., 2011) worked on creating manga previews by detecting panels and text balloons, while in (Deng et al., 2010; Chang and Cheng, 2011) authors explored generating manga faces and caricatures from real images.
 Cao et al.
-[187] and Herranz et al.
-[188] introduced methods for comic-like video summarization.
+(Cao et al., 2012) and Herranz et al.
+(Herranz et al., 2012) introduced methods for comic-like video summarization.
 These systems developed pipelines for layout proposal, image reconfiguration, and balloon creation, treating comic creation as a series of steps from keyframe selection to final panel layout.
 On the same task, to automatically create comics from video content, Jing et al.
-[189] pioneered video transformation using TV show subtitles and a speaker detection algorithm to generate comic layouts optimized through an energy-based metric.
-This approach was a significant step forward compared to earlier methods like [190].
+(Jing et al., 2015) pioneered video transformation using TV show subtitles and a speaker detection algorithm to generate comic layouts optimized through an energy-based metric.
+This approach was a significant step forward compared to earlier methods like (Wang et al., 2012).
 Giving it more time, a groundbreaking approach by Pesko et al.
-[191] employed neural style transfer, creating visually compelling comic representations from video-sourced data [^12] GitHub repository: https://github.com/maciej3031/comixify.
+(Pesko et al., 2018) employed neural style transfer, creating visually compelling comic representations from video-sourced data [^9] GitHub repository: https://github.com/maciej3031/comixify.
 While their startup has since become closed-sourced, their initial approach signifies a significant leap in comic generation technology.
 Yang et al.
-[192] presented a system to generate comics from movies, employing a multi-page layout framework and emotion-aware balloon generation, showcasing the potential of comprehensive systems in comic generation.
-To extract keyframes from movies, they used GIST [193] similarity between two frames.
+(Yang et al., 2021) presented a system to generate comics from movies, employing a multi-page layout framework and emotion-aware balloon generation, showcasing the potential of comprehensive systems in comic generation.
+To extract keyframes from movies, they used GIST (Oliva and Torralba, 2001) similarity between two frames.
 To detect the speaker, they perform lip motion analysis and use that to associate balloons with talking characters.
 Finally, the movie frames are rendered in a comic-layout page after the adaption to comic style.
-Moreover, in [194] authors represent conversational videos to comics “fixed-layout” representation, developing a system that learns guidance field which provides a prior prediction of the possible positions of word balloons while making the word balloons not overlap with other nonverbal information (e.g.
+Moreover, in (Zhang et al., 2023) authors represent conversational videos to comics “fixed-layout” representation, developing a system that learns guidance field which provides a prior prediction of the possible positions of word balloons while making the word balloons not overlap with other nonverbal information (e.g.
 hand gestures, visual clues in the background, etc).
 Their input is raw video, thus the text dialog is captured from audio and rendered in the location chosen from their method.
-Comics have been used also in education [195, 196, 197] to teach and explain in an easier way different concepts.
-An example of this is [198] where authors proposed a system that crafts data stories from a collection of user-created charts, using a comic-style panel to imply the underlying sequence and logic of data-driven narratives.
-10.1.2 Image-to-Text generation
+Comics have been used also in education (Suh et al., 2021; Lima et al., 2022; Castro et al., 2022) to teach and explain in an easier way different concepts.
+An example of this is (Zhao et al., 2021) where authors proposed a system that crafts data stories from a collection of user-created charts, using a comic-style panel to imply the underlying sequence and logic of data-driven narratives.
+10.1.2. Image-to-Text generation
 Definition: Image-to-Text is a general umbrella term that covers tasks from captioning, to textbox prediction in comics, masking a specific part of the dialog.
 Motivated by supporting the Blind or Low Vision community (BLV) - also known as People with Visual Impairements (PVI) - Ramaprasad et al.
-[199] developed a two-step method to create natural language descriptions of comic strips.
+(Ramaprasad, 2023) developed a two-step method to create natural language descriptions of comic strips.
 This method first extracts information about panels, characters, and text using computer vision techniques, followed by the use of a multimodal large language model (MLLM) to generate descriptions.
-This work stands as one of the few promoting MLLM specifically for comic strip captioning, despite using “out of the box” architectures like Grounding DINO [40] to extract arbitrary elements, CLIP [14] to match character images with name-descriptions, and LLaVa [200] as MLLM.
+This work stands as one of the few promoting MLLM specifically for comic strip captioning, despite using “out of the box” architectures like Grounding DINO (Liu et al., 2023c) to extract arbitrary elements, CLIP (Radford et al., 2021) to match character images with name-descriptions, and LLaVa (Liu et al., 2023b) as MLLM.
 Vivoli et al.
-[72] proposed an advanced version of the text-cloze task with image-to-text generation, increasing the challenge level.
-They adopted the VL-T5 model [16] using two configurations with similar performances but different model sizes: with BLIP-2 visual backbone (1.3B) or a ResNet-50 fine-tuned with SimCLR on comics (0.2B).
+(Vivoli et al., 2024d) proposed an advanced version of the text-cloze task with image-to-text generation, increasing the challenge level.
+They adopted the VL-T5 model (Cho et al., 2021) using two configurations with similar performances but different model sizes: with BLIP-2 visual backbone (1.3B) or a ResNet-50 fine-tuned with SimCLR on comics (0.2B).
 They surpassed previous state of the art in the standard “text-cloze” task, and propose benchmarks in the generation version of the task.
-Authors in [77] presented a similar task but incorporated character descriptions and transcriptions, emphasizing a language model-centric approach.
+Authors in (Agrawal et al., 2023) presented a similar task but incorporated character descriptions and transcriptions, emphasizing a language model-centric approach.
 Guo et al.
-[201] introduced the multi-modal manga complement task.
+(Guo et al., 2023) introduced the multi-modal manga complement task.
 This innovative task combines visual sequences from comic pages with corresponding text dialogues, challenging the model to complete the narrative appropriately.
 They propose an effective method with CLIP as Feature Encoding, a custom cross-attention module called Fine-grained Visual Prompt Generation, and a transformer-based encoder-decoder architecture called Dialog complement.
-In a recent work [41], authors proposed a Relationformer-based architecture that generates dialog for manga pages.
+In a recent work (Sachdeva and Zisserman, 2024), authors proposed a Relationformer-based architecture that generates dialog for manga pages.
 Their approach includes detection, matching, and sorting algorithms to produce an ordered dialog that aligns with the visual narrative.
-However, all these works are limited by the dialog appearing in the comic.
-No work has explored producing a multi-page description of the scene, the actions, and the surrounding environment of every single panel.
-10.1.3 Text-to-Image generation
+However, all these works are limited by the dialog appearing in the comic, and in the context of describing a comic, not only the dialog is important, but also the scene description.
+Recent works, to fill this gap, have approached the task by exploring single panel captioning.
+In (Rigaud et al., 2024), the authors explored prompt engineering with contextual information for closed-sourced Multimodal LLMs aiming at generating an accurate text description of a single panel and subsequently of the full story.
+They successfully demonstrated to retain many of the important details in the caption, but the exploration was limited to a couple of story panels.
+In (Vivoli et al., 2024b), authors took a step further proposing a new metric for comic panel captioning called “attribute retaining metric” (ARM) which assesses whether all the objects and attributes of the panel have been identified.
+They propose a pipeline with open-source Multimodal LLMs and zero-shot grounding models for generating dense and grounded captions for more than 1M panels.
+However, in the ARM metric only the attributes are taken into consideration, which identifies the need to consider both locations and object/attribute names.
+Finally, (Hu et al., 2024) focuses on comics with contradictory narratives, where each comic consists of two panels that create a humorous contradiction, and introduces the “YesBut” benchmark, which comprises tasks of varying difficulty aimed at assessing MLLMs capabilities in recognizing and interpreting these comics.
+10.1.3. Text-to-Image generation
 Definition: The Text-to-Image task comprehends generating an image, or a sequence of images, from a text description.
-In the field of Text-to-Image generation, one of the earliest endeavors was by Jin et al.
-[202], who successfully created anime-style character faces using a GAN trained on a curated dataset of anime faces, showcasing the potential of GANs in generating stylized images.
+In the field of Text-to-Image generation, early work by Jin et al.
+(Jin et al., 2017) utilized GANs to generate anime-style character faces, demonstrating the potential of GANs for stylized image generation.
 Inoue et al.
-[68] introduced novel methods for domain transfer and pseudo-labeling in image generation, marking significant advancements in the area.
-StoryGAN [203] emerged as a groundbreaking initiative in story visualization, employing a sequential conditional GAN framework equipped with a context encoder to dynamically track story flow and a story-level discriminator.
-Enhancements to StoryGAN have included a dual learning framework to bolster semantic consistency between stories and images by refining captions [204] and a recursive architecture to handle structured text inputs via constituency parsing trees [205].
-In [206], a deep learning pipeline was proposed for generating synthetic graphic novels using GPT-2 for text and StyleGAN2 for image synthesis, trained on a large collection of manga pages.
+(Inoue et al., 2018) advanced the field with domain transfer and pseudo-labeling techniques for image generation.
+StoryGAN (Li et al., 2018a) was a groundbreaking development in story visualization, employing a sequential conditional GAN with a context encoder to track story flow and a story-level discriminator.
+Later improvements included DuCo-StoryGAN (Maharana et al., 2021), which used a dual learning framework for better semantic consistency, and VLStoryGAN (Maharana and Bansal, 2021), which incorporated recursive architectures to handle structured text inputs.
+Melistas et al.
+(Melistas et al., 2021) proposed a pipeline for generating synthetic graphic novels, using GPT-2 for text and StyleGAN2 for image synthesis, trained on a large manga collection.
 Proven-Bessel et al.
-[207] developed a text-to-image GAN for generating single-panel comics in the style of Dilbert, showcasing new possibilities for comic generation from text.
-Another approach involved FastGAN [208], enhanced with a condition vector for generating high-quality images from small datasets, such as manga faces drawn from Osamu Tezuka’s works [209], which showed improved FID scores compared to the original FastGAN.
-In [210], a multilingual text-to-image model was developed for creating webtoons, demonstrating the adaptability of GANs to multilingual contexts.
-However, despite various training experiments, the webtoon fine-tuned GAN underperformed compared to fine-tuning Korean MSCOCO, with the latter still producing abstract images that noticeably differed from authentic comics.
-Significant progress has also been observed in story visualization through diffusion models.
-Innovations like StoryDALL-E [211] employ autoregressive Transformers for extensive model tuning and adaptive adjustments.
+(Proven-Bessel et al., 2021) developed ComicGAN, a text-to-image GAN for generating single-panel comics in the style of Dilbert, while FastGAN (Liu et al., 2021b), enhanced for small datasets, was applied to manga faces (Hiruta et al., 2022) with improved FID scores.
+Yu et al.
+(Yu et al., 2023) introduced a multilingual text-to-image model for webtoon generation, showing the flexibility of GANs across languages.
+However, even fine-tuning on Korean MSCOCO resulted in abstract images that differed from authentic comics.
+Diffusion models have also made strides in story visualization.
+StoryDALL-E (Maharana et al., 2022) used autoregressive Transformers for improved model tuning.
 Everaert et al.
-[212] discovered the ability to adapt Stable Diffusion to various styles, including comics, by adjusting the initial latent tensor, showing how diffusion models can contribute to comic generation.
-Despite diffusion models, both on pixel and on latent space, being not good at generating text, the comic panels generated are in line with COMIC data distribution.
-The topic of story visualization as a subfield of text-to-image generation is very active, moving also towards integration of more powerful LLMs [213], with the main challenges being the consistency of characters and style across multiple images (panels).
-Lastly, as we saw previously for cross-modal retrieval tasks, using existing models as building blocks to automatically annotate new data is a good option, especially when the task doesn’t exist or annotations are expensive.
-In [214], authors used an LLM (ChatGPT) to generate storylines and dialogue of a “possible continuation” of the well-known manga One Piece.
-Then, they fine-tuned stable diffusion with LoRA and ControlNet to adapt to the manga layout, colorization, and character aesthetic.
-They finally generated the comic using the fine-tuned stable diffusion, rendered it in panels, and added the LLM dialogs to balloons and narration boxes.
-10.1.4 Sound Generation
-Definition: Sound generation in the comics domain stands for generating sound effects for a single panel, or in the transaction from one panel to another.
-In [215], authors proposed a system that automatically narrates the input comic page of “Clubhouse Rascals” by synthesizing realistic speeches for each character.
-The synthesized speeches carry the identity properties (gender and age) and emotional conditions (e.g., happy, sad, angry) of the characters inferred from the comic input.
-They perform a series of subsequential steps including detection of panels, speech balloons, text, character, and character’s face, as well as character re-identification, text-character association, and dialog generation.
-They classify every character’s face with gender, age, and emotional state, and generate the comic speech synthesis given all this information as input to a waveform reconstruction.
-They used the Manga109 dataset for all experiments.
-In a later study [216], authors explored whether having scene description and sound effects could improve the user experience of comics for People with Visual Impairments (PVI), discovering that both types are welcomed with the first (scene description) helping fr concentration and understanding of the situation, and the second one (sound effect) perceived as more immersive and realistic book-reading experience.
-10.2 Synthesis
-The group of Synthesis tasks, compared to the “generation” one, refers to the creation of more complex, structured, and often temporally extended outputs, such as creating a full graphic novel from a complex storyline or merging multiple elements (characters, scenes, dialogues) in a consistent way across multiple pages.
-10.2.1 3D generation from images
+(Everaert et al., 2023) adapted Stable Diffusion for comic generation by tweaking the initial latent tensor, showing how diffusion models can be tailored for comic styles.
+Despite limitations in text generation, diffusion models successfully generated comic panels aligned with COMIC data distributions.
+The consistency of characters and style across multiple images, especially in storytelling contexts, remains a key challenge (He et al., 2024).
+Recent work by Jin et al.
+(Jin and Song, 2023) leveraged large language models (LLMs) and fine-tuned Stable Diffusion to generate new manga content.
+Using ChatGPT to generate storylines and dialogues for a hypothetical continuation of the manga One Piece, they fine-tuned Stable Diffusion with LoRA and ControlNet to match the layout, style, and aesthetic of the original manga.
+The final product included LLM-generated dialogue integrated into comic panels.
+The subfield of story visualization within text-to-image generation is rapidly evolving, driven by the integration of powerful LLMs and new image synthesis techniques.
+Key challenges include maintaining character and style consistency across multiple panels and developing unified benchmarks to track advancements across the field.
+10.2. Synthesis
+Figure 15. Illustration of Synthesis tasks.
+The group of Synthesis tasks, compared to the “generation” one, refers to the creation of more complex, structured, and often temporally extended outputs, such as creating a full graphic novel from a complex storyline or merging multiple elements (characters, scenes, dialogues) in a consistent way across multiple pages (Fig.
+15).
+10.2.1. 3D generation from images
 Definition: The task of generating 3D models from 2D comic illustrations presents unique challenges.
 Unlike conventional portrait illustrations, comics often involve stylized and exaggerated features, adding complexity to this already intricate process.
-While there are advancements in using 3D character assets for tasks like pose estimation [217], re-targetting [218, 219], and character reposing [220], the field lacks scalable training resources due to the unavailability of suitable 3D character assets.
+While there are advancements in using 3D character assets for tasks like pose estimation (Kim et al., 2019), re-targetting (Khungurn, 2022; Kim et al., 2021), and character reposing (Lin et al., 2022), the field lacks scalable training resources due to the unavailability of suitable 3D character assets.
 In light of these issues, Chen et al.
-[221] addressed this gap by formalizing the stylized reconstruction task with the introduction of the AnimeRecon benchmark and the Vroid dataset.
+(Chen et al., 2023a) addressed this gap by formalizing the stylized reconstruction task with the introduction of the AnimeRecon benchmark and the Vroid dataset.
 The Vroid dataset, in particular, provides a wealth of 3D assets for scalable training in this domain.
 A significant part of their work involves solving the challenge of contour removal from illustrations, which is a crucial step in achieving accurate 3D reconstruction from stylized comic images.
-While there are numerous works on 3D reconstruction from line drawings in general [222, 223], the specific application in comics is underexplored, likely due to the absence of suitable datasets and ground truth references for comics.
+While there are numerous works on 3D reconstruction from line drawings in general (Lun et al., 2017; Sanghi et al., 2023), the specific application in comics is underexplored, likely due to the absence of suitable datasets and ground truth references for comics.
 It is worth noting that the exploration of 3D reconstruction from comic images holds great potential.
 It not only enriches the visual experience of comics but also opens up possibilities for animations, interactive media, and virtual reality applications within the comic domain.
-10.2.2 Video generation
+10.2.2. Video generation
 Definition: Video generation in comic refers to the synthesis of video content from static comic panels.
 This process represents a significant leap in merging traditional comic art with dynamic multimedia.
-Recent advancements have seen significant progress in generating videos from both images [224, 225] and text [226], although accurately defining metrics and benchmarks remains challenging [79].
-These developments are particularly notable in the broader Vision and Language community, showcasing the fast-paced evolution of the field.
-The domain of comics, however, is still in the early stages of exploring video generation capabilities.
-In one of the initial attempts, Cao et al.
-[227] introduced a framework to animate manga panels.
-This method involves classifying panels based on motion and emotion and using these classifications to animate backgrounds and characters.
-The final video creation considers both intra-panel animations and transitions between panels.
+Recent advancements in video generation from both images (Yan et al., 2023; Wang et al., 2023a) and text (Gupta et al., 2023) have made significant strides, though defining accurate metrics and benchmarks remains a challenge (Fan et al., 2024).
+These developments are rapidly evolving within the broader Vision and Language community.
+However, in the domain of comics, video generation is still in its early stages.
+One of the first attempts came from Cao et al.
+(Cao et al., 2017), who introduced a framework to animate manga panels.
+Their method classifies panels based on motion and emotion, using this information to animate characters and backgrounds.
+The resulting video includes both intra-panel animations and transitions between panels.
 Later, Gupta et al.
-[228] further advanced this concept with the Comic-to-Video Network (C2VNet).
-This framework creates panel-audio videos by cropping comic panels and adding a highlighting effect to speech balloons, simulating characters’ dialogues.
-Alongside this, they introduced the IMCDB, an annotated dataset of Indian Mythological Comics in English, facilitating research in this area.
-While these efforts mark initial steps, they still fall short of the more advanced capabilities seen in general Vision-Language video generation, such as producing fully animated video clips with sound from a single image.
-The potential for integrating comics with video and audio elements presents an exciting opportunity for enhancing storytelling and audience engagement in the comic medium.
-Future research in this area could lead to new forms of interactive and immersive comic experiences, blending traditional art with contemporary digital technologies.
-10.2.3 Narrative-based Comic Generation
+(Gupta et al., 2021) expanded on this with C2VNet, which generates panel-audio videos by cropping panels and adding a highlighting effect to speech balloons to simulate character dialogue.
+They also introduced IMCDB, a dataset of annotated Indian Mythological Comics in English, to support research in this field.
+While these efforts represent initial steps, they are still far from the capabilities seen in Vision-Language video generation, such as producing fully animated video clips with sound from a single image.
+The integration of comics with video and audio elements offers exciting potential for enhancing storytelling and audience engagement.
+Future research could pave the way for interactive and immersive comic experiences, blending traditional art with modern digital technologies.
+10.2.3. Narrative-based Comic Generation
 Definition: The task involves generating a comic (whether a panel, strip, or page) from a plain text description.
@@
 The pioneering efforts in the realm of narrative-based comic generation commenced with the introduction of StoryGAN by Li et al.
-[229], which marked a significant advancement in generating image sequences from storylines using a sequential conditional GAN framework.
+(Li et al., 2018b), which marked a significant advancement in generating image sequences from storylines using a sequential conditional GAN framework.
 This innovation incorporated a Context Encoder, which dynamically tracked the story flow, laying the groundwork for future developments in this area.
 Building on this foundation, Maharana et al.
-[230] enhanced visual quality and coherence by incorporating a dual learning framework that utilized video captioning, a copy-transform mechanism, and MART-based transformers.
+(Maharana et al., 2021) enhanced visual quality and coherence by incorporating a dual learning framework that utilized video captioning, a copy-transform mechanism, and MART-based transformers.
 This approach significantly improved the semantic alignment between the story and generated images, thus enriching the story visualization process.
 In terms of character coherence and continuity, Chen et al.
-[231] focused on maintaining consistent character portrayal throughout the narrative.
+(Chen et al., 2022) focused on maintaining consistent character portrayal throughout the narrative.
 They adapted Vector-Quantized Variational Autoencoders (VQ-VAE) with a text-to-visual-token architecture, ensuring character coherence in the visual narratives.
 Furthering this direction, Maharana et al.
-[211] introduced the concept of story continuation, where the generated visual story is conditioned on a source image, allowing narratives to incorporate new characters more fluidly.
+(Maharana et al., 2022) introduced the concept of story continuation, where the generated visual story is conditioned on a source image, allowing narratives to incorporate new characters more fluidly.
 They also introduced the DiDeMoSV dataset, which provided a platform for exploring these complex narrative structures.
 Recent advancements in story visualization are exemplified by Pan et al.
-[232], with presenting AR-LDM, a latent diffusion model that significantly raised the standards for visual quality in natural image datasets like VIST.
+(Pan et al., 2022), with presenting AR-LDM, a latent diffusion model that significantly raised the standards for visual quality in natural image datasets like VIST.
 This model was auto-regressively conditioned on history captions and generated images, capturing complex interactions between frames.
 Rahman et al.
-[233] introduced a framework with a visual memory module that implicitly captured actor and background context, generating frames that were not only of high visual quality but also consistent with the story.
+(Rahman et al., 2022) introduced a framework with a visual memory module that implicitly captured actor and background context, generating frames that were not only of high visual quality but also consistent with the story.
 Peng et al.
-[234] proposed PCSG, a diffusion-based text-to-image synthesis framework with controllable plugins for character consistency, scene layout specification, and character pose specification, further enhancing the personalized aspects of story visualization.
-These emerging trends and future directions in narrative-based comic generation showcase a notable shift towards more sophisticated, context-aware, and visually coherent storytelling techniques.
-The integration of elements like character consistency, scene dynamics, and actor-background interaction is pivotal in creating compelling visual narratives.
-Future research in this area can further refine these techniques, potentially leading to automated comic creation tools that adapt to various narrative styles and complexities, opening new horizons in the field of comics and visual storytelling.
-10.3 Future Tasks
-10.3.1 Comics to Scene graph
+(Peng et al., 2023) proposed PCSG, a diffusion-based text-to-image synthesis framework with controllable plugins for character consistency, scene layout specification, and character pose specification, further enhancing the personalized aspects of story visualization.
+These emerging trends and future directions in narrative-based comic generation showcase a notable shift towards more sophisticated, context-aware, and visually coherent storytelling techniques, potentially leading to automated comic creation tools that adapt to various narrative styles and complexities.
+10.3. Future Tasks
+10.3.1. Comics to Scene graph
 Definition: Transforming comics into scene graphs is an emerging area of research.
 This task involves dissecting comic panels to identify and represent the relationships between different elements within the panel in a graph format.
-The last work that tackles this task is [41], which employs a Relationformer architecture [145], explicitly designed for image-to-graph tasks.
+The last work that tackles this task is (Sachdeva and Zisserman, 2024), which employs a Relationformer architecture (Shit et al., 2022), explicitly designed for image-to-graph tasks.
 However, the authors convert the full page into different types of nodes (panels, text, characters) and only consider two types of arches: characters-characters (re-identification) and text-characters (speaker identification).
@@
 The task has yet to be explored to fine-grained graph creation from single panels, full pages, or a book.
-11 CONCLUSION
+11. CONCLUSION
 In reviewing the diverse and innovative efforts in the realm of comic research and technology, it becomes evident that this field is in a state of dynamic and exciting evolution.
@@
 In the future, we will continue to monitor advancements in this field, and we will update our findings in the following repository: https://github.com/emanuelevivoli/awesome-comics-understanding.
-A. Methodology details
-In this section, we will explore the process of finding the state-of-the-art in Comics and Manga.
-We utilized the software “Publish or Perish” [^13] Available at: https://harzing.com/resources/publish-or-perish, which comprehends the following Search sources: “Semantic Scholar”, “Google Scholar”, “Web of Science” and “Scopus”.
-As every source possesses its own interface and query structure, we designed a “SQL-like” query:
-  • Title: “comic*”, “manga”
-  • Keywords: (“comic*” OR “manga”) AND (“machine*” OR “learning*” OR “ml” OR “computer*” OR “science*” OR “artificial*”)
-For every source, we limited the results to the last 1000 from 2010 to 2024.
-For every paper (out of the 3500 obtained), we used the “Semantic Scholar API” [^14] API at link: https://www.semanticscholar.org/product/api to query the title and obtain the unique DOI.
-We finally filtered unique papers, automatically removing duplicated copies (DOI) copies, obtaining 2276 papers.
-This duplicate-removal process has been further implemented also in “zotero” [^15] Zotero software available at: https://forums.zotero.org, fetching data from the websites and using javascript code [^16] Zotero deduplication script.
-Finally, we define our inclusion/exclusion criteria as follows:
-Inclusion. The study regards:
-  • (I1)
-    Comics and Manga;
-  • (I2)
-    Computer science;
-  • (I3)
-    Deep Learning;
-  • (I4)
-    new Dataset and/or Task.
-Exclusion. The study is:
-  • (E1)
-    not written in English;
-  • (E2)
-    not available;
-  • (E3)
-    a duplicate or extension of an already included study;
-  • (E4)
-    a secondary or tertiary study;
-  • (E5)
-    in the form of editorials, tutorials, books, extended abstracts, etc.;
-  • (E6)
-    a non-scientific publication or grey literature;
-  • (E7)
-    cited at least 2 times (if previous 2022);
-  • (E7)
-    not published in AI conferences.
-After this process, we ended with 216 papers.
-The last phase of “snowballing” is applied at this point: whi le reading and analyzing the papers we discovered important references that were not in the collection because they were not selected by our query or were discarded by I/E criteria.
-However, this process is needed to include important papers that are highly cited from our collection.
-This step enlarged the collection to the final number of 353 papers.
-Finally, we manually check all the references and include the one that suits the best our taxonomy and scope.
-B. Taxonomy
-Task classification within the LoCU is informed by both the modalities and dimensions of task inputs and outputs.
-We define modalities as Images (ranging from no image to multiple pages), Text (from no text to complex multi-sentence structures), Audio (from no audio to complex sound mixes), and Geometry (encompassing points, lines, bounding boxes, segmentations, and 3D geometries).
-Dimensional analysis considers Space (from non-spatial to tridimensional) and Time (whether time is a factor).
-Each task is scored based on its input and output characteristics in these modalities and dimensions (see Tables IV and V), with the cumulative score determining its placement within the LoCU layers.
-TABLE IV: Modality Scoring for LoCU Tasks
-Modality Category Score
-Images No image 0
-Single panel 1
-Single page (multiple panels) 2
-Multiple pages 3
-Text No text 0
-Single word (or tag) 1
-Single sentence 2
-Multi sentences 3
-Instruct multi-sentences 4
-Complex multi-sentence 5
-Audio No audio 0
-Single sound 1
-Multi-elements sound 2
-Complex mix of sounds 4
-Geometry No 0
-Point 1
-Line 2
-Bounding box 3
-Instance segmentation 4
-Semantic segmentation (multiple masks) 5
-3D object geometry 6
-TABLE V: Dimensions Scoring for LoCU Tasks
-Dimensions Category Score
-Space No 0
-2D 1
-3D 2
-Time Not considered 0
-Considered 1
-In Table VI are reported the values of modality and dimensions, together with the Layer of Comics Understanding Table I.
-This shows the modalities and dimensions value assigned to every task’s input and output, in order to obtain an objective complexity score that could sort each of the tasks.
-The acronyms “I, T, A, G” represent “Images, Text, Audio and Geometry”, respectively.
-Similarly, the acronyms “S, T” correspond to the Dimensions “Space, and Time”, respectively.
-TABLE VI: Detailed view of Modalities and Dimensions for both inputs and outputs.
-Mods Dim Mods Dim
-Layer Category Tasks Input I T A G S T Output I T A G S T Score
-0 View Basic Image Viewing (BIV) Text Command 0 0 0 0 0 0 Image Display 0 0 0 0 0 0 0
-Image Classification (I-CLS) Image 1 0 0 0 0 0 Tag 0 1 0 0 0 0 2
-Emotion Classification Comic Panels/Images 1 0 0 0 0 0 Emotion Labels 0 1 0 0 0 0 2
-Action Detection Multiple Panels 2 0 0 0 1 1 Tag 0 1 0 0 0 0 5
-Tagging Page Stream Segmentation (PSS) Images 2 0 0 0 1 0 Tags sequence 0 1 0 0 0 0 4
-Image Super-Resolution (ISR) Image 1 0 0 0 1 0 Image 1 0 0 0 1 0 4
-Style Transfer (ST) Image 1 0 0 0 1 0 Image 1 0 0 0 1 0 4
-Vectorization Comic Panels/Images 1 0 0 0 1 0 Vector image 1 0 0 1 1 0 5
-1 Augmentation Depth Estimation Comic Panels/Images 1 0 0 0 1 0 Depth Map 1 0 0 0 1 0 4
-Object Detection Tag/s + Image 1 1 0 0 1 0 Bounding Boxes 0 0 0 2 1 0 6
-Character Re-identification (Character ID) Multiple Panels 2 1 0 0 1 0 Text [Tagging] 0 1 0 2 0 0 7
-Grounding Grounding (IG) [prompt] + Image 1 2 0 0 1 0 Bounding Boxes 0 0 0 2 1 0 7
-Character-Balloon Association (Speaker ID) Character + Balloons 2 0 0 2 1 0 Text [Tagging] 0 0 0 2 0 0 7
-Dialog transcription Image 1 0 0 0 1 1 Text 0 3 0 0 0 1 7
-Analysis Translation Image 1 3 0 0 1 0 Text 0 3 0 0 0 0 8
-2 Segmentation Instance Segmentation (IS) [prompt] + Image 1 0 0 2 1 0 Segments 0 0 0 3 1 0 8
-Image-Text Retrieval (IR) Text 1 3 0 0 1 0 Image 0 2 0 0 1 0 8
-Text-Image Retrieval (TR) Image 3 1 0 0 1 0 Text 2 0 0 0 1 0 8
-Retrieval Composed Image Retrieval (CIR) Text + Image 3 2 0 0 1 0 Image 2 0 0 0 0 0 8
-Image Inpainting (II) Text + [prompt] + Image 1 2 0 2 1 0 Image 2 0 0 0 1 0 9
-3 Modification Image Editing via Text (IET) Text + Image 1 4 0 0 1 0 Image 2 0 0 0 1 0 9
-Visual Dialog (VisDial) Image + Dialog + Text 2 4 0 0 1 1 Text 0 3 0 0 1 1 13
-Visual Reasoning (VR) Image + Text 2 3 0 0 1 1 Text 0 3 0 0 1 1 12
-4 Understanding Visual Question Answering (VQA) Image + Text 2 3 0 0 1 1 Text 0 2 0 0 1 0 10
-Image Captioning [img-2-text] (IC) Image 1 0 0 0 2 0 Text 0 3 0 0 2 0 8
-Image Generation [text-2-img] (IG) Text 0 5 0 0 1 0 Image 1 0 0 0 2 0 9
-Scene Graph Generation for Captioning Comic Panel 1 0 0 0 2 0 Scene Graph 0 3 0 3 2 0 11
-Generation Sound Generation from Single Panel Single Comic Panel 1 0 0 0 2 0 Sound/Audio 0 0 4 0 3 1 11
-3D Model Generation from Images (3DGI) Collection of Images 3 0 0 0 1 0 3D Model 0 0 0 3 3 0 10
-Video Generation from Text (VGT) Complex Long Text 0 5 0 0 1 1 Video 2 1 0 0 1 1 12
-5 Synthesis Narrative-Based Complex Scene Generation (NCSG) Detailed Narrative Text 2 5 0 0 1 1 Series of Images 3 1 0 0 1 1 15
-[Uncaptioned image] Emanuele Vivoli is pursuing its joint-PhD at the Computer Vision Center (Autonomous University of Barcelona) and at the MICC Lab (University of Florence).
-His research interests encompass multi-modal learning (vision and language integration), with a focus on comics.
-His work covers a diverse range of topics, including document and table analysis, understanding comics and manga, and creating datasets and frameworks.
-If you are interested, read his papers.
-[Uncaptioned image] Andrey Barsky is a postdoctoral researcher at the Computer Vision Center, Barcelona, Spain.
-He received his Ph.D.
-in 2015 from the University of Nottingham in the UK.
-His research focuses on computer vision and multimodal learning, as well as robustness and explainability in AI models.
-[Uncaptioned image] Mohamed Ali Souibgui is a postdoctoral researcher at Computer Vision Center, Barcelona, Spain.
-He received the Ph.D.
-degree in 2022 from the Universitat Autònoma de Barcelona (UAB), Spain.
-His current research focuses on document image analysis using computer vision and machine learning tools, with the focus on privacy and explainability.
-He is an author of several journal and peer-reviewed conference papers.
-[Uncaptioned image] Artemis Llabrés is a PhD student in the Computer Vision Center.
-She is part of the Vision, Language, and reading group where she is working on multimodal models for document understanding.
-In 2024 she was awarded best student paper by the 16th IAPR International Workshop on Document Analysis Systems, for her work titled “Image-text matching for large-scale book collections”.
-[Uncaptioned image] Marco Bertini is an Associate Professor in Computer Science at the University of Florence, Italy.
-He is the director of the Media Integration and Communication Center of the University of Florence.
-His interests are focused on multimedia and computer vision.
-On these subjects he has addressed semantic analysis, automatic content indexing, semantic retrieval and video quality improvement, applying these techniques to different domains among which cultural heritage.
-He is author of more than 30 journal papers and more than 140 peer-reviewed conference papers.
-He has been involved in 10 EU research projects as WP coordinator and researcher, among which IM3I, euTV, ORUSSI, UMETECH, AI4Media and ReInHerit.
-He has been general and program co-chair of several conferences on multimedia.
-He is co-founder of Small Pixels, an academic spin-off working on GenAI solutions to improve video quality and video compression.
-[Uncaptioned image] Dimosthenis Karatzas is a professor at the Universitat Autónoma de Barcelona and associate director of the Computer Vision Centre, leading the Vision and Language research group.
-His research focuses on computer vision and machine learning, with an emphasis on robust reading systems and document image analysis.
-He received the 2013 IAPR/ICDAR Young Investigator Award.
-He has created two spin-off companies.
-He chaired the IAPR TC11 on Reading Systems from 2016 to 2021, he is an IEEE senior member, an ELLIS fellow and co-director of the ELLIS Unit Barcelona, and a member of the Artificial Intelligence Doctoral Academy (AIDA) Research and Industry Board.
-He launched the Robust Reading Competition portal, which serves more than 50,000 researchers worldwide.